### v1

# Identify As A Human Does: A Pathfinder of Next-Generation Anti-Cheat Framework for First-Person Shooter Games

Abstract
The gaming industry has experienced substantial growth, but cheating in online games poses a significant threat to the integrity of the gaming experience. Cheating, particularly in first-person shooter (FPS) games, can lead to substantial losses for the game industry. Existing anti-cheat solutions have limitations, such as client-side hardware constraints, security risks, server-side unreliable methods, and both-sides suffer from a lack of comprehensive real-world datasets. To address these limitations, the paper proposes Hawk, a server-side FPS anti-cheat framework for the popular game CS:GO. Hawk utilizes machine learning techniques to mimic human experts’ identification process, leverages novel multi-view features, and it is equipped with a well-defined workflow. The authors evaluate Hawk with the first large and real-world datasets containing multiple cheat types and cheating sophistication, and it exhibits promising efficiency and acceptable overheads, shorter ban times compared to the in-use anti-cheat, a significant reduction in manual labor, and the ability to capture cheaters who evaded official inspections.
1 Introduction
The gaming industry has experienced substantial growth recently, with the global game market generating \$184.0 billion in 2023 and projected to reach \$218.7 billion in 2024 [47, 48]. However, cheating in online games poses a significant threat to the integrity of gaming experiences and disrupts gameplay balance. In 2023 alone, cheating led to an estimated \$29 billion in losses, with 78% of gamers deterred from playing due to cheating [32]. In particular, first-person shooter (FPS) games, which account for 20.9% of total game sales and rank second among all genres [63], are particularly targeted by cheat developers seeking monetary gains due to their competitive and multiplayer nature. The most prevalent forms of cheating involve the use of aimbots to enhance a player’s precision and wallhacks to reveal opponents’ positions, thereby compromising game integrity [4, 40, 9].
On realizing the critical importance of game security [25], both academia and industry have proposed many anti-cheats solutions, applied either on the client-side [9, 58, 49] or server-side [69, 24, 70, 2, 39, 27]. However, current solutions have the following limitations.
Client-side solutions face hardware constraints, bring security and privacy concerns, or incur unacceptable system overheads. Two state-of-the-art systems [9, 49] require specific Trusted Execution Environments (TEEs) like Intel SGX. This dependence on particular hardware, which was deprecated in 2021 [5, 31], limits compatibility and practicality for industrial use. Some solutions involve scanning hard drives and gaining root privileges, posing risks of unauthorized access and personal data leakage [45, 43, 66]. Vulnerabilities, such as RCE exploits [8] and CVEs [15, 16, 17, 18], demonstrate the potential for system instability and privacy breaches. The others rely on continuous surveillance or frequent patching to counter varying cheat signatures [22, 3, 35, 58], leading to increased client-side system overheads.
On the other hand, the existing server-side solutions use basic features and share low recall and accuracy. For instance, some solutions [69, 24, 70, 2, 39, 27] use features like winning rates or playtime, which are insufficient for detecting complex cheats such as aimbots or wallhacks. These methods also lack robust workflows to examine false positives, especially in noise-rich, real-world datasets. Additionally, they show low recall and accuracy in the real world, with extended ban cycles taking days or weeks [33, 50]. Incidents like the 2024 APEX Legends tournament hack [59] highlight the inefficacy of existing solutions under commercial anti-cheats such as Easy Anti-Cheat (EAC) [22].
Moreover, both client-side and server-side solutions suffer from a lack of comprehensive datasets, often relying on limited or synthetic data [9, 49, 24, 70, 2, 39], which may not represent real-world cheating scenarios. The simulated datasets lack varying levels of cheating sophistication (explained in Section 2.2), which could result in performance degradation or even failure in complex real-world scenarios.
To address the aforementioned limitations, we propose Hawk, a server-side approach designed to analyze data transmitted to the server, thereby avoiding the inherent limitations of client-side solutions. Hawk introduces a comprehensive set of multi-view features and a robust framework to meticulously monitor players’ points of view, statistics, and behavioral consistencies. This approach includes a corresponding workflow aimed at shortening the ban cycle and considering the occurrence of false positives. Hawk achieves high detection recall and acceptable accuracy across large, real-world datasets with varying levels of cheating sophistication.
The key innovation of Hawk lies in its ability to mimic how human experts identify cheaters by focusing on three main aspects: analyzing the player’s point of view, statistically evaluating the player’s performance, and assessing the player’s gaming sense and performance consistency. Based on these aspects, we design structured and temporal features that accurately represent player behaviors. Hawk reproduces the human identification process using Long Short-Term Memory (LSTM) and attention mechanisms, ensemble learning, and deep learning networks, respectively.
We evaluate Hawk on a mainstream FPS game, CS:GO. To the best of our knowledge, this paper is the first server-side work using a substantial volume of real-world datasets with different cheat types and levels of sophistication, making Hawk and the datasets more convincing for industrial use. Specifically, our dataset includes 2,979 aimbots and 2,971 wallhacks, totaling 56,041 players. The scale of the dataset has increased by two orders of magnitude compared to the state-of-the-art work [9]. Hawk achieves a maximum of 84% recall and 80% accuracy in detecting both aimbots and wallhacks, outperforming the official in-use anti-cheat. It successfully identifies cheaters who had evaded previous detections and those never caught by the official systems. We created a demo website [^1] Demo website. This is an anonymous website. to showcase some escapees’ illegal actions. Hawk validates its robustness against cheat evolution (introduced in Section 6.7) in real-world scenarios and showcases acceptable overheads. We particularly discuss Hawk’s generalizability in Section 7. The contributions of our work are as follows.
  • Innovative anti-cheat observations and feature designs.
  • Novel framework Hawk and a supporting workflow.
  • Real-world evaluations with strong performance.
  • Open-source Hawk and dataset [^2] Link. The repository and dataset will be released upon acceptance. Artifacts are uploaded for the reference. Ethics statement is shown in Section 10. for future research.
2 Background
This section introduces anti-cheat and cheating techniques, replay systems, and replay files in modern FPS games.
2.1 Server-side and Client-side Anti-Cheat
Client-side anti-cheats focus on data, process, and hardware protection [49, 58], or detection by either examining the existing cheats’ signatures [22, 3] or classifying through behavior [9]. However, attackers can bypass it with higher privileges, by sniffing the detection report through side-channel attack [23] and altering them in the transport layer, or by changing the program’s signature [35].
Server-side anti-cheats [69, 70, 2, 24, 39] emphasize data analysis and anomaly detection. Although the cheater’s inputs might be altered on the client, the cheaters must ensure their performance outperforms others to secure the win. Thus, it is easier to detect cheats on the server because it records all clients’ final-state actions that influence the gameplay. But the prior works have limitations mentioned in Section 1.
2.2 Cheat Types in FPS Games
Wallhack and aimbot are the two most prevalent types of cheats in FPS games [1]. Both types of cheats are mainly done by accessing the client’s memory data or image information. Wallhack enables players to visually penetrate obstructions like walls, which allows players to spot opponents across the map and effortlessly track their movements, providing a significant tactical advantage. Aimbot assists a player’s aims and shots so that the cheaters gain transcendent precision and reaction speed. Aimbot’s functionality can be customized according to player preference, allowing one to obtain different aiming performance, ranging from brutal-force aims to smooth aims like human [9]. Aimbot can be subdivided into pure aimbot (the primitive aimbot), triggerbot (auto-shoot only when the cross-hair is on the opponent), micro-settings (counteract recoil with Hotkeys), and computer-vision-based aimbot (aim with object detection models). In the following sections, aimbot refers to all aforementioned subcategories.
Notably, cheating provides different levels of sophistication [66]. The cheater might cheat cautiously, pretending to be a normal player, switching the cheat on and off, or cheating unabashedly with different configurations, various brands of cheating software, and categories. However, cheaters’ behaviors will always differ from the normal players [9]. Thus as long as cheaters want to gain unfair advantages, they must outperform the average. This is bound to create anomalies at the data level and therefore should be identifiable by models. Our evaluations in Section 6 include different levels of cheating sophistication by using large real-world datasets.
2.3 Replay System and Replay File
Replay system is a dynamic toolset for player interaction with recorded content post-game and has been extensively adopted in notable FPS games. Unlike video recordings, this system enables users to explore various perspectives, adjust camera angles, and control playback speed for analysis by reconstructing the match with the replay file. Replay file refers to a comprehensive record of a match, encapsulated in a unique data format that can be interpreted by the replay system. Demo denotes the replay file in CS:GO. Hawk innovatively utilizes demo to extract data and avoid concurrent overheads on the server. We believe the replay file can be used as a novel data source for FPS server-side anti-cheat.
3 Overview
This section provides the key observations on the human experts’ identification process and the workflow of Hawk.
3.1 Fact Observations
Although the anti-cheat tasks are still challenging for automation systems in not just the FPS field but almost any game genre, the illegal activities are identifiable for seasoned human players [11]. Therefore, by harnessing such inspiration, we discovered three observations that human experts always focus on to identify cheaters. {mdframed}[style=obs] Observation 1 (OBS.1). Review a player’s point-of-view (POV). Akin to watching a player’s game replay unfold over time, this review process involves analyzing various in-game behaviors, such as aim and shot, positioning and movement, and the in-game economy management and props utilization. These behaviors can provide temporal insights into whether the player is legitimately playing the game. {mdframed}[style=obs] Observation 2 (OBS.2). Conduct an in-depth analysis of a player’s statistical data. Cheaters often exhibit statistical anomalies that set them apart from normal players. These anomalies include abnormally high accuracy and eliminations, a low death rate, and a high number of kills made through obstacles, etc. These statistical deviations serve as strong indicators of cheating. {mdframed}[style=obs] Observation 3 (OBS.3). Examine the consistency between a player’s game sense and combat performance. A player requires substantial game experience to have better performances, which results in a positive correlation between a player’s game sense and their performance. Thus, the inconsistency between the two aforementioned factors indicates potential cheating. For instance, if a player moves like a rookie but shoots and aims incredibly accurately, it is suspicious.
3.2 Hawk’s Workflow
Figure 1 illustrates Hawk’s workflow, comprising four interlinked stages.
Stage One. Preprocessing. The demo is applied to parse in-game data from the banned databases. The multi-view features are then extracted from parsed data and then annotated for training. We discuss the feature constructions in Section 4.
Stage Two. Hawk deployment. Hawk utilizes the extracted features as inputs for its subsystems RevPov (Section 5.1), RevStats (Section 5.2), and ExSPC (Section 5.3) to mimic human identification process, respectively. The outputs are incorporated into Mvin subsystem (Section 5.4) that presents a final cheating report.
Stage Three. Game Master (GM) team verification. Given the potential personal assets tied to a player’s account, a ban should not be issued solely based on an algorithmic verdict. It is imperative to provide substantial evidence before initiating such an action. Hence, the final verdict requires GMs’ verification to avoid false bans.
Stage Four. Data updates. Upon the affirmation of cheating, the relevant replay files are added to the banned databases. This act enriches and updates the reservoir of training samples which may contain novel patterns and behaviors. Thus, four stages constitute a cyclical workflow, with each stage contributing to the subsequent one.
This ongoing cyclical workflow enhances the system’s robustness and provides a dynamic solution to multi-type cheat detection. We further discuss this in Section 6.7
Figure 1: Cyclical workflow of Hawk.
4 Feature Constructions
The feature design is one of our contributions and key to the server-side anti-cheat. This section introduces the construction of temporal features, structured features, and the classification of sense and performance features. These features are intended to be a comprehensive representation of the three aspects of human identification of cheating in Section 3.1.
4.1 Temporal Features
Temporal features are information derived from the demo in the time domain. Each temporal data tuple represents the state information (e.g., coordinates, view directions, events, etc.) associated with the present tick. Where tick denotes the smallest unit of time in the game. This section describes temporal features’ components and illustrates the distinctions in temporal features between normal players and cheaters.
4.1.1 Temporal Features Construction
According to different in-game behaviors, temporal features are segmented into three main categories and subdivided into seven specific types. This overview provides only categories and types due to the less-expressive nature of temporal features. Appendix B lists elaborate descriptions.
  • Engagement features describes in-game engagement events, which are subcategorized into five types as follows.
    • Damage features indicate damage-inflicting events, including the occurring tick, the attacker’s and victim’s locations, view directions, weapons, damage value, etc.
    • Auxiliary props features related to auxiliary props-utilization events (e.g., flashbang), including the deploy tick, victim’s affected duration, etc.
    • Offensive props features related with events involving props that aid the attack (e.g., grenade, incendiary, etc.) including the deploy and destroy ticks, the props’ type, the props’ coordinates, etc.
    • Elimination features describes kill-related events, including the occurring tick, attacker’s weapon and location, victim’s location and view direction, etc.
    • Weapon fire features denotes firing-related occurrences, including the occurring tick, the weapon information, the shooter’s location and view direction, etc.
  • Movement features represents in-game positioning, mobility, and boolean flags per tick. For instance, coordinates, view directions, velocities in three dimensions, flags that indicate if the player is ducking, blinded, reloading, etc.
  • Economy features reflects in-game financial aspects. For example, they include the equipment value and balance per round for each player.
4.1.2 Temporal Data Comparative Visualization
Figure 2 comparatively analyzes the temporal features between cheaters and normal players across one round. Where round denotes the unit determining each win-loss outcome. One match has multiple rounds, and the players who win more rounds are the winners. In this example, there are four differences between cheaters and normal players. First, since the cheaters illegally acquire their opponents’ position, the timing of cheaters’ firing (pink) and engaging (red) highly overlap (a). Second, normal players stay on guard (blue) irregularly to check for potential enemy locations, whereas cheaters are often on guard just before engagement (b). Third, normal players deploy props (orange) to gain a gunfighting or positional advantage, whereas cheaters rarely do so because they already have illegal advantages (c). Fourth, cheaters are rarely vulnerable (yellow) during the fire or engagement because they know the enemy’s location and therefore can circumvent it in advance (d).
Figure 2: Comparative behavioral visualization between honest players and cheaters with respect to time scales.
4.2 Structured Features
Structured features are single-value data calculated with well-designed algorithms to represent integrated in-game behavioral information per match. This section describes the components of structured features and visualizes the structured feature differences between normal players and cheaters.
4.2.1 Structured Features Construction
The structured features for Hawk are categorized into four main categories that comprise 28 strong-expressiveness features. For the sake of brevity and coherence, only the introduction of categories is outlined below. A meticulous breakdown can be referred to in Appendix C.
  • Aiming features represents the efficiency and efficacy of aims. We break down the aiming process into three moments (initial spot, first-time fire, and first-time hit) and two stages (reaction and adjustment). We extract different elapsed durations and angle variations with statistical representations (e.g., average or variance) to indicate the player’s different performances within a match. For example, the average of reaction duration reflects a player’s level of reaction speed, and the variance of that denotes the distribution of the reaction duration. Typically, in terms of the reaction stage, normal players obtain a higher mean value and more diverse variance compared with cheaters.
  • Firing features reflects shot proficiency. For example, the hit group distribution variance describes the dispersion of body part that suffers damage for the current player in a match. Most cheaters with aimbot prefer to lock on specific hit group (i.e., head) to achieve faster elimination.
  • Elimination features emphasizes metrics related to in-game eliminations or kill-focused activities. For example, the occluder-penetration index is a weighted index that reflects the extent of the average occluder-penetration eliminations (e.g., eliminate the opponent behind walls or smoke) conducted by the player. In terms of the cheaters exploiting wallhacks, this index is more significant than the normal players’.
  • Props utilization features captures a player’s adeptness at using in-game auxiliary and offensive props like flashbang, grenades, etc. For instance, the props utilization index denotes how frequently a player utilizes auxiliary and offensive props such as incendiaries and smoke grenades. This can help distinguish a player’s familiarity with the game and their skillfulness.
Figure 3: Box plots of comparative visualization between honest players and cheaters on top-12 structured features under Mann-Whitney U (Wilcoxon Rank-Sum) test with P-Value ascending (distinction descending) order.
4.2.2 Structured Data Comparative Visualization
Figure 3 offers a box-plot on structured features to analyze between honest players and cheaters comparatively. Each feature on the x-axis (we select top-12 features for illustration purposes) is ordered based on the outcomes of the Mann-Whitney U test, with the leftmost features, such as akpr (introduced in Appendix C), marking pronounced disparities between the two groups. This implies that features towards the left end, showcased by their lower p-values, are more distinguishable, whereas those on the right are less so. The y-axis is normalized values between 0 and 1. The box plots, differentiated by blue for honest players and orange for cheaters, provide insights into the median, interquartile range, and outlier distributions. In which, diamond shapes represent outliers, emphasizing the spread and distribution of data across the features. Despite considering dimensionality reduction techniques, such as PCA (Figure 13), empirical tests show an optimal performance when all 28 features were employed.
4.3 Sense and Performance Features
Sense and performance features are derived from the two subsections above. Appendix D includes detailed classification.
  • Sense features denote a player’s in-game cognition, reflecting strategic understanding, and tactical anticipation.
    • Temporal sense features include temporal features that reflect a player’s gaming sense such as economic considerations, movement dynamics, grenade usage, etc.
    • Structured sense features contain structured features that describe in-game understanding, e.g., flash efficiency, prop utilization, etc.
  • Performance features demonstrate in-game efficacy, indicating engagement capabilities, precision in aiming, and mastery over game mechanisms.
    • Temporal performance features include temporal features that indicate continuous in-game performances, e.g., eliminations, damage, weapon proficiency, etc.
    • Structured performance features contain the remaining structured features that statistically describe the player’s in-game performances, such as time to kill, elimination-related features, etc.
Figure 4: Hawk framework illustration. (a)-RevPov, (b)-RevStats, (c)-ExSPC are corresponding to three observations in Section 3.1, (d)-Mvin is an integration network for combining the determinations from the aforementioned three subsystems.
5 Hawk
To mimic human experts’ identification process, we propose Hawk which consists of four subsystems: Review Point-of-View (RevPov) subsystem, Review Statistics (RevStats) subsystem, Examine Sense-Performance Consistency (ExSPC) subsystem, and Multi-View Integration (Mvin) subsystem. The first three subsystems’ designs are associated with the three observations we discovered regarding the inherent nature of the cheat identification process as elaborated in Section 3.1. The fourth subsystem integrates the results derived from the first three subsystems.
5.1 RevPov
Aligned with OBS.1, RevPov thoroughly examines a player’s time-series operations within a match.
Multi-LSTM Attention Encoders: In Figure 4 (a), we feed the seven sets of temporal features (Section 4.1.1) as input to seven independent Long Short-Term Memory (LSTM) [29] networks to acquire the embeddings. For explanatory purposes, we illustrate the process using the example of the Economy features. $\mathbf{F_{eco}^{n}}$ denotes the input feature vector. A dropout layer [54] is employed from the second to last LSTM layers respectively to mitigate the risk of overfitting. $\vec{h}_{G}^{n}$ denotes the output of the encoder, consists of one set of the seven temporal features mentioned in Section 4.1.1, where $G$ represents the set and $n$ indicates the sequence number among all outputs. For each time step, we introduce the Luong-style attention [41] afterward to help the model better understand the importance of different positions in the input sequence and aggregate different parts of the input weights. In Equation 1, $a_{G}^{i}$ represents the attention weight for the time step $i$. $\mathbf{W}_{q}$ and $\mathbf{W}_{k}$ denotes the learned weight matrices of query and key in attention mechanism. $i$, $j$ and $k$ indicate different time steps. Concatenation operation is executed afterward, yielding $[\vec{a}_{G}^{n}||\vec{h}_{G}^{n}]$, where ||||| | denotes concatenation. After two dense layers, $\vec{w}_{G}^{n}$ represents the final tensor for binary classification.
$\vec{a}_{G}^{i}=\sum_{j=0}^{N}\frac{\exp{(\mathbf{W}_{q}\cdot\vec{h}_{G}^{i})\cdot(\mathbf{W}_{k}\cdot\vec{h}_{G}^{j})}}{\sum_{k=0}^{N}\exp{(\mathbf{W}_{q}\cdot\vec{h}_{G}^{i})\cdot(\mathbf{W}_{k}\cdot\vec{h}_{G}^{k})}}\cdot\vec{h}_{G}^{j}$ (1)
Embedding Generation: An average pooling is conducted to flatten the output sequence. Equation 2 indicates the average operation, where $N$ denotes the number of output sequences. $\vec{V}_{pov}$ is used as embedding for the final determination.
$\vec{w}_{G}=\frac{\sum_{n=0}^{N}\vec{w}_{G}^{n}}{N+1}$ (2)
$\vec{V}_{pov}=[\vec{w}_{eco}||\vec{w}_{mov}||\vec{w}_{dmg}||\vec{w}_{elm}||\vec{w}_{aux}||\vec{w}_{off}||\vec{w}_{wf}]$ (3)
Final Determination: In the context of cheat detection, the dataset often exhibits a class imbalance, where the majority of players are honest, inadvertently leading to biased learning towards the dominant class. Therefore, we adopt a multi-subsampling strategy to level the playing field for the minority class, creating balanced representations of both cheating and honest behaviors. Suppose $\mathcal{T}$ is the initial training set. To establish a balanced class distribution, we perform a selective sampling on $\mathcal{T}$, specifically targeting the honest player samples. We perform this operation nine times in accordance with the ratio of honest to dishonest players, generating nine distinct training sets, each denoted as $\mathcal{T}_{k}$, where $k$ indicates the iteration count. Conduct each iteration of the sampling process to select unique honest player samples, thereby adding varied examples to the learning process. Subsequently, we obtain nine separately trained Random Forest models, each corresponding to one of the diversified and balanced training sets $\mathcal{T}_{k}$, we named this strategy as multi-forest. The averaging method in Equation 4 is selected for RevPov final class prediction. Consider a random forest consisting of $N$ decision trees, each producing a prediction denoted as $h_{i}(\vec{V}_{pov}^{k})$, where $i$ indicates the $i$-th tree, $k$ denotes the $k$-th sampled set, and $\vec{V}_{lstm}^{k}$ represents the input feature vector. The prediction of each tree is a probability distribution, indicating the likelihood of the sample $\vec{V}_{pov}^{k}$ belonging to each class. $p_{i}$ represents the probability that the $i$-th tree predicts sample $\vec{V}_{pov}^{k}$ to be in class $j$. In Equation 4, $\mathbf{\overline{RF}}(\vec{V}_{pov}^{k})$ denotes the average prediction output of the random forest, which is the average of the prediction probability vectors of all decision trees. The prediction for each $\vec{D}_{k}$ can be obtained by selecting the class with the highest predicted probability shown in Equation 5, where $\left(\mathbf{\overline{RF}}(\vec{V}_{pov}^{k})\right)_{j}$ represents the probability of class $j$ in $\mathbf{\overline{RF}}(\vec{V}_{pov}^{k})$. After the multi-forest, we acquire the corresponding determination $\mathbf{D}_{k}$. Lastly, as in Equation 6, we obtain the final classification determination $\mathbf{D}_{pov}$ for RevPov.
$\mathbf{\overline{RF}}(\vec{V}_{pov}^{k})=\frac{1}{N}\sum_{i=1}^{N}h_{i}(\vec{V}_{pov}^{k})=\frac{1}{N}\sum_{i=1}^{N}\mathbf{p}_{i}$ (4)
$\mathbf{D}_{k}=\arg\max_{j}\left(\mathbf{\overline{RF}}(\vec{V}_{pov}^{k})\right)_{j}$ (5)
$\mathbf{D}_{pov}=\arg\max_{j}\sum_{k=1}^{9}\mathbf{D}_{k}^{(j)}$ (6)
5.2 RevStats
Aligned with OBS.2, RevStats subsystem is designed to perform an in-depth analysis of structured features to capture the cheaters. Cheaters often exhibit deviations in their static data compared to normal players. Even when cheaters deliberately attempt to conceal their behavior, the subtle differences tend to accumulate over time and reveal disparities in the statistical data post-game.
Feature Extraction: In Figure 4 (b), $\vec{V}_{28}$ denotes the vector representation of the structured features (Section 4.2.1).
Ensembled Committee: We empirically use seven classifiers [28, 14, 7, 13, 46, 34, 55] in an ensembled way. Each model is independently trained. Upon completion, each model independently adjudicates each test sample. The outputs are binary signifying the presence or absence of cheating.
Final Determination: Due to imbalanced dataset, a similar multi-subsampling strategy as employed in the RevPov subsystem is adopted for the RevStats subsystem to fully leverage the available data without losing valuable information. For each classification model, we generate a distinct model for each subsampled set. Each model provides an individual prediction. A majority voting scheme is utilized within each classification algorithm to aggregate the results of the nine models. The final determination is also achieved through the majority voting mechanism, which collates the determinations from all seven classification algorithms. In Equation 7, where $\mathbf{D}_{stats}$ signifies the binary final determination. Where $w_{j}$ denotes the decision from the $j$th classification algorithm, and $I(\cdot)$ is the indicator function, returning 1 if the condition within the brackets is true, and 0 otherwise. The decision, $i$, which receives the majority of votes from all seven classification algorithms, is accepted as the final determination.
$\mathbf{D}_{stats}=\arg\max_{i\in\{0,1\}}\sum_{j=1}^{7}I(w_{j}=i)$ (7)
5.3 ExSPC
Aligned with OBS.3, ExSPC subsystem is designed to inspect a comprehensive investigation of the consistency of a player’s in-game senses and overall performances.
Embedding and Feature Grouping: There are seven embeddings and two groups of vectors as input (Section 4.3) in ExSPC, in which seven embeddings are generated by using the trained models in RevPov and two vectors are the subsets of $\vec{V}_{28}$ in RevStats. Each embedding is padded to the maximum length of that in the whole dataset. $\vec{w}^{k}_{G}$ is the intermediary result from RevPov, where $k$ represents the numbering in the sequence, ${G}$ denotes the type of the temporal features. Each embeddings are represented in two-dimensional arrays. In Figure 4 (c), $P_{1}$ to $P_{k}$ indicates different players, we exemplify with one player $P_{j}$. Therefore, we flatten the embeddings into one-dimensional arrays and use a dense layer with an ELU activation function to shrink the dimensions. After the seven different similar groups of layers, we obtain $\varphi_{G}^{{P}_{j}}$. Subsequently, we divide seven different types of tensors as shown in Appendix D, and concatenate each group of tensors for each player. In which, sense vector of the temporal features is $\vec{V}_{POV}^{S}=[\varphi_{mov}^{{P}_{j}}||\varphi_{eco}^{{P}_{j}}||\varphi_{off}^{{P}_{j}}||\varphi_{aux}^{{P}_{j}}]$ and performance vector of the temporal features is $\vec{V}_{POV}^{P}=[\varphi_{wf}^{{P}_{j}}||\varphi_{elm}^{{P}_{j}}||\varphi_{dmg}^{{P}_{j}}]$. In terms of two vectors, we also divided $\vec{V}_{28}$ into two groups according to Appendix D, namely the sense vector of structured features $\vec{V}_{28}^{S}$ and the performance vector of structured features $\vec{V}_{28}^{P}$.
Dimension Reduction: $\vec{V}_{28}^{S}$ and $\vec{V}_{28}^{P}$ are normalized at first. Afterward, each group of the four vectors goes through different groups of dense layers to reduce the dimensions. Next, we concatenate four vectors and yield two vectors eventually representing sense $\vec{V}_{SENSE}=[\vec{\xi}_{SENSE}^{STATS}||\vec{\xi}_{SENSE}^{POV}]$ and performance $\vec{V}_{PREF}=[\vec{\xi}_{PREF}^{STATS}||\vec{\xi}_{PREF}^{POV}]$ respectively.
Final Determination: When acquiring $\vec{V}_{SENSE}$ and $\vec{V}_{PERF}$, we further reduce its dimension and deepen the network for better learning performances. Then, we concatenate two vectors and go through one final dense layer for reducing the dimension and another dense layer with sigmoid activation function for binary classification. Eventually, the final classification determination $\mathbf{D}_{spc}$ for ExSPC subsystem is obtained. The class weight is set as one to nine like previous subsystems. Both RevPov and ExSPC use Binary Cross-Entropy as loss function shown in Equation 8, where $N$ represents the number of samples, $y_{i}$ is the true label of the $i$-th sample, $\hat{y}_{i}$ is the predicted value for the $i$-th sample.
$\text{Loss}=-\frac{1}{N}\sum_{i=1}^{N}\left[y_{i}\log(\hat{y}_{i})+(1-y_{i})\log(1-\hat{y}_{i})\right]$ (8)
5.4 Mvin
In Figure 4 (d), Mvin is to harness the collective determinations from RevPov, RevStats, and ExSPC. By integrating the predictions from these subsystems, Mvin aims to deliver a final verdict regarding a player’s cheating disposition.
Data Integration: The subsystem integrates the outcomes from RevPov, RevStats, and ExSPC, symbolized as $\mathbf{D}_{pov}$, $\mathbf{D}_{stats}$, and $\mathbf{D}_{spc}$. Given the diverse nature of these outputs, a structured preprocessing phase becomes indispensable. This step is designed to merge the individual outputs and synchronize them for the succeeding integration process.
[…]
6 Experiments
We conduct experiments to answer six research questions (RQ.1--RQ.6):
  • RQ.1: What are the performances of Hawk and each subsystem? (Section 6.4)
  • RQ.2: Compared to the current official inspection, to what extent Hawk has improved in terms of effectiveness and efficiency? (Section 6.5)
  • RQ.3: What are the performances of Hawk under different settings of the Task-Specified Threshold Optimizer? (Section 6.6)
  • RQ.4: What is the robustness of Hawk and its subsystems against cheat evolution? (Section 6.7)
  • RQ.5: Is there any suspicious activity that still exists but has never been detected by officials? (Section 6.8)
  • RQ.6: What are the overheads of Hawk’s subsystems? (Section 6.9)
6.1 Incomparability to Prior Works
First, most prior works [69, 24, 70, 2, 39, 71] are not open-sourced. Second, data collection method is distinct and incompatible. Hawk utilizes the replay file post-game, which differs fundamentally from prior studies that modified the game engine or applied mods during gameplay. Third, the installation on commercial clients or servers is infeasible. If the installation of the prior works is implemented before the commencement of data collection, the data can then be gathered and evaluated. However, the demos we collected come directly from the game platform, which prohibits the installation of any unverified programs on clients or servers due to legal constraints. In addition to the aforementioned reasons with commonality, we also further state below the specific reasons for the detailed incomparability between Hawk and the SOTA designs. There are three aspects of incomparability with BotScreen [9]. First, CS:GO has gone through several major patch updates after BotScreen was released. The authors of BotScreen confirmed through email that BotScreen, as well as Osiris [37] (the identical cheat used in BotScreen for feature extraction and evaluation), does not apply to the in-use version of games. Second, the deprecation of Intel SGX [5, 31] prevents us from conducting comparative experiments. Third, it can only detect pure aimbot, making the comparison less significant. BlackMirror [49] and Invisibility Cloak [58] are prevention designs only for counteracting wallhacks and computer-vision-based aimbots respectively. Both studies are only functional before the cheats happen. Whereas Hawk and other prior works are detection systems that functional post-cheating.
Table 1: Dataset separation and description.
CT1 Train Validation Test
#M2 #P3 #N4 #C5 #M2 #P3 #N4 #C5 #M2 #P3 #N4 #C5
Aimbot 1,009 10,127 9,150 977 975 8,561 7,697 864 995 9,047 8,163 884
Wallhack 833 8,320 7,500 820 1,037 9,712 8,745 967 1,101 10,274 9,249 1,025
1CT: Cheat type; 2#M: Number of matches; 3#P: Number of players in total; 4#N: Number of normal (honest) players; 5#C: Number of cheaters.
Table 2: Ablation study results on each of the subsystems.
CT1 System Validation Test
TP TN FP FN Accuracy Recall NPV AUC-ROC OEI2 TP TN FP FN Accuracy Recall NPV AUC-ROC OEI2
Aimbot RevPov 458 4,796 2,901 406 0.614 0.530 0.922 0. 577 1.246 481 5,088 3,075 403 0.616 0.544 0.927 0.584 1.283
RevStats 453 7,290 407 411 0.904 0.524 0.947 0.736 4.941 566 6,797 1,366 318 0.814 0.640 0.955 0.736 2.864
ExSPC 553 6,187 1,510 311 0.787 0.640 0.952 0.779 2.529 558 6,616 1,547 326 0.793 0.631 0.953 0.786 2.586
Hawk 614 5,803 1,894 250 0.750 0.711 0.959 0.732 2.326 642 5,837 2,326 242 0.716 0.726 0.960 0.721 2.126
Wallhack RevPov 573 6,081 2,664 394 0.685 0.593 0.939 0.644 1.670 627 6,364 2,885 398 0.680 0,612 0.941 0.650 1.684
RevStats 604 8,241 504 363 0.911 0.625 0.958 0.783 4.143 855 6,988 2,261 170 0.763 0.834 0.976 0.795 1.946
ExSPC 760 7,258 1,487 207 0.826 0.786 0.972 0.885 3.303 501 8,484 765 524 0.875 0.489 0.942 0.840 3.736
Hawk 798 7,033 1,712 169 0.806 0.825 0.977 0.815 3.118 869 6,902 2,347 156 0.756 0.848 0.978 0.797 2.649
1CT: Cheat type; 2OEI: Operational Efficiency Index.
Table 3: Comparison result on the different model selection of RevStats.
CT1 Model Validation Test
TP TN FP FN Accuracy Recall NPV AUC-ROC OEI2 TP TN FP FN Accuracy Recall NPV AUC-ROC OEI2
Aimbot LogisticRegression 570 6,204 1,493 294 0.791 0.660 0.955 0.733 2.614 713 4,897 3,266 171 0.620 0.807 0.966 0.703 1.772
Naive Bayes 467 6,740 957 397 0.842 0.874 0.944 0.708 3.069 625 5,928 2,235 259 0.724 0.707 0.958 0.717 2.143
RandomForest 415 7,694 3 449 0.947 0.480 0.945 0.740 4.313 427 7,634 529 457 0.891 0.483 0.944 0.709 4.313
SVM 309 7,615 82 555 0.926 0.358 0.932 0.673 7.299 405 7,665 498 479 0.892 0.458 0.941 0.699 4.320
MLP 858 7,697 0 6 0.999 0.993 0.999 0.997 9.901 508 7,325 838 376 0.866 0.575 0.951 0.736 3.674
GaussianNB 467 6,740 957 397 0.842 0.541 0.944 0.708 3.069 625 5,928 2,235 259 0.724 0.707 0.958 0.717 2.143
QDA 350 7,181 516 514 0.880 0.405 0.933 0.669 3.737 534 6,770 1,393 350 0.807 0.604 0.951 0.717 2.697
RevStats 453 7,290 407 411 0.904 0.524 0.947 0.736 4.941 566 6,797 1,366 318 0.814 0.640 0.955 0.736 2.864
Wallhack LogisticRegression 714 7,629 1,116 253 0.859 0.738 0.968 0.805 2.861 943 5,622 3,627 82 0.639 0.920 0.986 0.764 1.420
Naive Bayes 604 7,677 1,068 363 0.853 0.625 0.955 0.751 2.741 891 6,103 3,146 134 0.681 0.869 0.979 0.765 1.545
RandomForest 967 8,745 0 0 1.000 1.000 1.000 1.000 6.713 751 8,083 1,166 274 0.860 0.733 0.967 0.803 2.872
SVM 548 8,323 422 419 0.913 0.567 0.952 0.759 4.378 817 7,227 2,022 208 0.783 0.797 0.972 0.789 2.065
MLP 472 8,489 256 495 0.923 0.488 0.945 0.729 5.186 788 7,726 1,523 237 0.829 0.769 0.970 0.802 2.471
GaussianNB 604 7,677 1,068 363 0.853 0.625 0.955 0.751 2.741 891 6,103 3,146 134 0.681 0.869 0.979 0.765 1.545
QDA 493 8,276 469 474 0.903 0.510 0.946 0.728 4.057 815 7,270 1,979 210 0.787 0.795 0.972 0.791 2.095
RevStats 604 8,241 504 363 0.911 0.625 0.958 0.783 4.143 855 6,988 2,261 170 0.763 0.834 0.976 0.795 1.946
1CT: Cheat type; 2OEI: Operational Efficiency Index.
6.2 Experimental Settings and Ground Truth
The demos are open-sourced and acquired from a well-known and recognized CS:GO platform [^3] Link. We anonymize the platform’s URL and name. The initial data extraction utilizes an open-source library awpy [67] to parse the raw data in demos into JSON files, which contain comprehensive information for the following feature constructions.
Table 1 is the dataset description. Figure 5 illustrates the experiments’ process. We evaluate Hawk with two of the most prevalent cheat types, aimbot and wallhack. For aimbot, we set one collection window (16 days in total). For wallhack, we set two collection windows (27 days in total, we explain the reason for window inconsistency in Section 6.7). All demos are real-world ranking matches within the collection windows. Therefore, the cheats used by the player are black-boxes for anti-cheating. Both datasets contain different levels of cheating sophistication.
The dates of cheaters’ label acquisition both end on the last day of data collection in each set (i.e., first labeling), this label is used for training and most of the evaluations in Figure 5. The official inspections are continuous, including multiple industry-in-use proprietary closed-source anti-cheats and manual checks. The officials update the ban list daily. However, due to the long ban cycle of the official inspections, most cheaters will not get banned on the day they cheat. Therefore, after 65 days of the first labeling, we once again acquired the labels (i.e., second labeling) of the whole datasets to screen out cheaters who had not been labeled correctly in the first labeling for further analysis on the false positives (the increased cheater numbers are shown in Table 4).
Although the two datasets contain different types of cheat, the official in-use anti-cheats can not label the specific subcategories, but can only annotate the broad categories of cheats. Cheating sophistication is subjective and difficult to quantify. Hence, even though we want to provide further information, it is infeasible to provide a more detailed breakdown of the datasets. Besides, due to the low recall and accuracy of the current in-use anti-cheats, there will be inevitable noise (cheaters who have evaded the official anti-cheats) in the negative labels within the ground truth.
Figure 5: Experiments’ process and labeling details.
6.3 Evaluation Metrics
Our chosen evaluation metrics are more skewed towards reflecting how well the cheaters can be identified. We use the following metrics in Appendix F and Operational Efficiency Index (OEI) to distinguish between cheating and normal behaviors. OEI is a new metric for evaluating the efficiency of industrial operations. Operational Efficiency Index (OEI), represented in Equation 10, where $N$ represents the total number of players. It reflects the operational efficiency gains in manual review processes. A higher OEI indicates a greater reduction in manual verification and a more effective system.
$OEI=\left(\frac{N}{TP+FP}\right)\cdot(Recall\cdot NPV)$ (10)
6.4 Ablation Study Results
To answer RQ.1, we conduct an ablation study on Hawk shown in Table 2. We set the Task-Specified Threshold Optimizer to achieve the highest accuracy with at least 70% recall for detecting aimbot, and the highest AUC-ROC for detecting wallhack, illustrating different anti-cheat scenarios. Overall, Hawk excels in identifying authentic cheaters, evidenced by the highest recall and NPV across all evaluations. For aimbot detection, RevStats enhances accuracy by effectively filtering normal players in both validation and test sets. In wallhack detection, RevStats excels in the validation set, while ExSPC leads in the test set. Although some components like RevStats and ExSPC outperform Hawk in accuracy, their lower recall (up to 64%) limits their effectiveness, affirming Hawk as the optimal choice. Notably, the Wallhack dataset yields better overall performance than the Aimbot dataset. We believe this is due to the latter’s diverse subcategories of aimbots. In Table 5 and Table 3, we present RevPov’s and RevStats’s model selection results among hundreds of tested models, and promising models are demonstrated. RevPov outperforms other models with balanced, consistent accuracy and recall. Random Forest outperformed other models under the sampled data in RevPov. Thus, we adopt multi-forest and multi-subsampling with Random Forest. RevStats outperforms individual classifiers in balancing recall and accuracy by capitalizing on the strengths of each classifier while mitigating their weaknesses. The training and validation loss for ExSPC in Figure 9 and Figure 10 (further discussed in Appendix A).
Table 4: The manual average ban duration comparison and the remediation rate among Hawk’s false positives after the first labeling.
CT1 Hawk 2 Off.3 Validation Test
#MTPs4 Rmd%5 #IncrPs6 #MTPs4 Rmd%5 #IncrPs6
Aimbot $\sim$4 min 5.88 days 93 4.91% 237 97 4.17% 233
Wallhack 4.47 days 93 5.43% 226 153 6.52% 303
1CT: Cheat type;
2
Hawk: The process time of Hawk varies with match length, Hawk’s detailed overheads are evaluated in Section 6.9;
3
Off.: The average ban duration of official manual inspections per match;
4
#MTPs: The number of missed true positives among the false positives;
5
Rmd%: The percentage of #MTPs in the number of false positives;
6
#IncrPs: The number of increased positives (cheaters) labeled by the officials after the second labeling.
Table 5: Comparison result on the different model selection of RevPov. N/A in OEI indicates the denominator is 0.
CT1 Model Validation Test
TP TN FP FN Accuracy Recall NPV AUC-ROC OEI2 TP TN FP FN Accuracy Recall NPV AUC-ROC OEI2
Aimbot REPTree [21] 19 7,660 37 845 0.897 0.022 0.901 0.509 3.028 33 8,118 45 851 0.901 0.037 0.905 0.516 3.919
RandomTree 123 6,914 783 741 0.822 0.142 0.903 0.520 1.215 125 7,399 764 759 0.832 0.141 0.907 0.524 1.305
RseslibKNN [26] 40 7,460 237 824 0.876 0.046 0.901 0.508 1.289 33 7,868 295 851 0.873 0.037 0.902 0.501 0.929
SPAARC [68] 0 7,697 0 864 0.899 0.000 0.899 0.500 N/A 0 8,163 0 884 0.902 0.000 0.902 0.500 N/A
RandomForest 16 7,665 32 848 0.897 0.019 0.900 0.507 2.974 33 8,128 35 851 0.902 0.037 0.905 0.517 4.496
Bagging [6] 18 7,666 31 846 0.898 0.021 0.901 0.508 3.278 30 8,129 34 854 0.902 0.034 0.905 0.515 4.341
VFI [20] 822 445 7,252 42 0.148 0.951 0.914 0.505 0.922 848 472 7,691 36 0.146 0.959 0.929 0.509 0.944
JRip [10] 21 7,663 34 843 0.898 0.024 0.901 0.510 3.408 42 8,116 47 842 0.902 0.048 0.906 0.521 4.376
MODLEM [57] 19 7,588 109 845 0.889 0.022 0.900 0.504 1.323 34 8,029 134 850 0.891 0.038 0.904 0.511 1.873
CHIRP [65] 0 7,697 0 864 0.899 0.000 0.899 0.500 N/A 0 8,163 0 884 0.902 0.000 0.902 0.500 N/A
CSForest [52] 50 7,292 405 814 0.858 0.058 0.900 0.503 0.980 58 7,755 408 826 0.864 0.066 0.904 0.508 1.151
RevPov 458 4,796 2,901 406 0.614 0.530 0.922 0.577 1.246 481 5,088 3,075 403 0.616 0.544 0.927 0.584 1.283
Wallhack REPTree 47 8,641 104 920 0.895 0.049 0.904 0.518 2.825 66 9,140 109 959 0.896 0.064 0.905 0.526 3.421
RandomTree 178 8,006 738 789 0.843 0.184 0.910 0.550 1.776 198 8,433 816 827 0.840 0.193 0.911 0.552 1.782
RseslibKNN 89 8,515 230 878 0.886 0.092 0.907 0.533 2.540 120 8,939 310 905 0.882 0.117 0.908 0.542 2.540
SPAARC 69 8,679 66 898 0.901 0.071 0.906 0.532 4.652 87 9,164 85 938 0.900 0.085 0.907 0.538 4.599
RandomForest 27 8,713 32 940 0.900 0.028 0.903 0.512 4.149 35 9,212 37 990 0.900 0.034 0.903 0.515 4.400
Bagging 26 8,705 40 941 0.899 0.027 0.902 0.511 3.571 37 9,201 48 988 0.899 0.036 0.903 0.515 3.940
VFI 950 563 8,182 17 0.156 0.982 0.971 0.523 1.014 1,008 689 8,560 17 0.166 0.983 0.976 0.529 1.031
JRip 36 8,670 75 931 0.896 0.037 0.903 0.514 2.941 71 9,179 70 954 0.900 0.069 0.906 0.531 4.572
MODLEM 35 8,687 58 932 0.898 0.036 0.903 0.515 3.414 56 9,171 78 969 0.898 0.055 0.904 0.523 3.789
CHIRP 6 8,738 7 961 0.900 0.006 0.901 0.503 4.176 8 9,242 7 1,017 0.900 0.008 0.901 0.504 4.816
CSForest 0 8,745 0 967 0.900 0.000 0.900 0.500 N/A 0 9,249 0 1,025 0.900 0.000 0.900 0.500 N/A
RevPov 573 6,081 2,664 394 0.685 0.593 0.939 0.644 1.670 627 6,364 2,885 398 0.680 0,612 0.941 0.650 1.684
1CT: Cheat type; 2OEI: Operational Efficiency Index.
Table 6: Hawk detection result with different Task-Specified Threshold Optimizer settings.
CT1 Optimizer2 Validation Test
TP TN FP FN Accuracy Recall AUC-ROC NPV OEI3 TP TN FP FN Accuracy Recall AUC-ROC NPV OEI3
Aimbot F1-Score 448 7,305 392 416 0.906 0.519 0.734 0.946 5.000 548 6,928 1235 336 0.826 0.620 0.734 0.954 3.000
Accuracy 269 7,528 169 595 0.911 0.311 0.645 0.927 5.640 305 7,888 275 579 0.906 0.345 0.656 0.932 5.014
AUC-ROC 648 4,603 3,094 216 0.613 0.750 0.674 0.955 1.639 691 4,364 3,799 193 0.559 0.782 0.658 0.958 1.508
$\text{Acc}_{\textit{rc}>0.75}$ 652 5,321 2,376 212 0.698 0.755 0.723 0.962 2.052 672 5,365 2,798 212 0.667 0.760 0.709 0.962 1.907
$\text{Acc}_{\textit{rc}>0.8}$ 700 4,611 3,086 164 0.620 0.810 0.705 0.966 1.769 716 4,700 3,463 168 0.599 0.810 0.693 0.965 1.693
$\text{Acc}_{\textit{rc}>0.85}$ 736 3,790 3,907 128 0.529 0.852 0.672 0.967 1.519 761 3,875 4,288 123 0.512 0.861 0.668 0.969 1.495
$\text{Acc}_{\textit{rc}>0.9}$ 778 2,859 4,838 86 0.425 0.900 0.636 0.971 1.333 787 2,955 5,208 97 0.414 0.890 0.626 0.968 1.301
Wallhack F1-Score 576 8,313 432 391 0.915 0.596 0.773 0.955 5.481 651 8,040 1,209 374 0.846 0.635 0.752 0.956 3.352
Accuracy 374 8,582 163 593 0.922 0.387 0.684 0.935 6.543 167 9,158 91 858 0.908 0.163 0.577 0.914 5.932
$\text{Acc}_{\textit{rc}>0.7}$ 677 7,898 847 290 0.883 0.700 0.802 0.965 4.304 803 7,506 1,743 222 0.809 0.783 0.797 0.971 3.071
$\text{Acc}_{\textit{rc}>0.75}$ 727 7,589 1,156 240 0.856 0.752 0.810 0.969 3.759 859 6,964 2,285 166 0.761 0.838 0.795 0.977 2.675
$\text{Acc}_{\textit{rc}>0.8}$ 776 7,186 1,559 191 0.820 0.802 0.812 0.974 3.251 583 8,280 969 442 0.863 0.569 0.732 0.949 3.574
$\text{Acc}_{\textit{rc}>0.85}$ 822 6,712 2,033 145 0.776 0.850 0.809 0.979 2.831 871 6,810 2,439 154 0.748 0.850 0.793 0.978 2.579
$\text{Acc}_{\textit{rc}>0.9}$ 871 5,755 2,990 96 0.682 0.901 0.779 0.984 2.229 855 6,885 2,364 170 0.753 0.834 0.789 0.976 2.598
1CT: Cheat type; 2Optimizer: Task-Specified Threshold Optimizer in Mvin; 3OEI: Operational Efficiency Index.
6.5 Hawk and Official Inspection Comparison
To answer RQ.2, we calculate the daily number of bans for the day and cumulative ban-sums by Hawk and the official inspections illustrated in Figure 6. The figure indicates that Hawk’s identification effectiveness (daily count in red bars) and efficiency (cumulative count in red lines) outrank those of the official manual inspections (blue bars and lines) on both cheat types. Additionally, Table 4 compares the average ban duration per match, revealing that Hawk’s seconds-scale processing time is exponentially more efficient than official inspections. The remediation rate indicates the percentage of cheaters not banned in the first labeling but caught in the second labeling. In Table 4, we list the remediation rate after the second labeling among Hawk’s false positives (which are obtained after the first labeling). It demonstrates that Hawk can capture the authentic cheaters among the normal players initially labeled by the official anti-cheats. This also suggests that while Hawk is robust to mislabeled data, using a purer dataset would further improve results.
Figure 6: Daily number of bans for the day and cumulative ban-sums by Hawk versus the official manual inspections.
6.6 Analysis on Different Mvin Settings
To address RQ.3, we conduct a comparative analysis of Hawk under different Mvin settings, as shown in Table 6. The Optimizer column indicates the priority choices of the Task-Specified Threshold Optimizer, such as F1-Score for highest F1-Score and $\text{Acc}_{\textit{rc}>0.75}$ for highest accuracy with recall above 75%. An inevitable trade-off exists between capturing more cheaters and minimizing false positives due to cheating sophistication and dataset noise (Section 6.2). As discussed in Section 5.4, GMs can dynamically adjust optimizers, allowing for quick retraining of Mvin, with new weight matrices and thresholds generated in about one minute. This adaptability aligns with the varying intensity of required anti-cheat measures, offering flexibility in balancing performance and efficiency. GMs can use different optimizers to meet specific anti-cheat needs. Interestingly, GMs do not always aim to catch all cheaters. They sometimes balance catching cheaters with permissiveness, as certain cheaters may bring short-term profits (e.g., in buyout games, where bans may be delayed until after refund periods). When complaints rise, GMs may then conduct a mass ban (a regularly organized operation to clean up the gaming environment) to restore community confidence. This also partially draws to the creation of Mvin.
6.7 Robustness to Cheat Evolution
Although anti-cheat measures are often opaque to prevent circumvention, cheat developers treat them as black boxes, adapting and camouflaging their cheats on a timely basis to evade detection. Cheat evolution refers to the diminishing effectiveness of specific anti-cheat over time. Commercially used anti-cheat methods (e.g., signature or process detection, driver scanning, etc.) are susceptible to this evolution since they target specific cheats. However, Hawk starts the analysis from the behavior, thus cheaters should appear data-level distinctions mentioned in Section 2.2. Hence, no matter how the cheats adapt, as long as the cheaters still want to gain unfair advantages, it is hard to bypass the detection of Hawk. To address RQ.4, we test Hawk’s robustness to cheat evolution using the Aimbot dataset, divided into partitions (40 matches each). Classified through date, the dataset is divided into partitions (40 matches per partition). The reason for selecting the Aimbot dataset is that during its collection window, an official mass ban on aimbots took place (at the 10th partition), and the cheats are usually updated within a few days after such an incident. Figure 7 and Figure 8 demonstrate the ablated performances of subsets under different metrics of Hawk across different numbers of partitions for training. According to the average ban duration of the official manual inspections in Table 4, approximately three sequestration ban cycles have passed during the collection window, and at least two ban cycles after the mass ban occurred. The ablated results of Hawk on both sets have been fluctuating to maintain within a small range over partitions (time) and with the cheat evolution. Besides, in Figure 6, we observe the effectiveness of Hawk has not continuously decreased over time on both datasets (we separated and extended the collection windows in Wallhack dataset for further proof). Thus, both of the results indicate Hawk is robust to cheat evolution.
Figure 7: Ablated robustness against cheat evolution over partitions on Aimbot’s validation set under different metrics.
Figure 8: Ablated robustness against cheat evolution over partitions on Aimbot’s test set under different metrics.
6.8 Detecting Uncaught Cheaters in the Wild
To answer RQ.5, given the above-known presence of cheaters in the false positives in Table 4 after the first labeling, we conduct a use case analysis to explore if there are additional illegal activities that exist after the second labeling. Given the lengthy average duration of a match, which demands significant manpower, we randomly select 30 Hawk’s false positives from the Aimbot and Wallhack validation sets for manual review, respectively (experts’ skill level is listed in Table 8). With the assistance of three experts in FPS games, we identify 27 suspects, who are labeled as normal players after the second labeling, to be highly suspicious within these 60 false positives. Of these 27 suspects, 14 originate from the Aimbot dataset (4 are considered using aimbots and 10 are considered boosting, i.e., elevate the account rank by other higher-rank players playing the account), while 13 are from the Wallhack dataset (8 are considered using wallhacks and 5 are considered boosting). We reported these suspicions to the official personnel. Upon their review, they confirm all 12 of the aforementioned aimbots and wallhacks suspects are committed to cheats, subsequently issuing bans for these cheaters on the platform. This reaffirms the poor performance of current in-use anti-cheats and the datasets used for training and evaluation are not entirely pristine. The website mentioned in Section 1 demonstrates their illegal actions. In terms of boosting suspects, there are no existing solutions for the officials to evidently verify their illegal actions. Therefore, there are no bans issued on the suspicious boosting accounts.
Table 7: Maximum overheads of Hawk’s subsystems per match. The numbers are retrieved by the Time [36] in Linux.
Max. Overheads per Match RevPov RevStats ExSPC Mvin
User time (sec)1 415.72 178.12 36.89 3.97
System time (sec)2 69.32 436.61 27.12 13.53
Elapsed time (mm:ss)3 03:40 00:14 00:45 00:01
Percent of CPU this job got4 220% 4371% 140% 2860%
Maximum resident set size (GB)5 3.56 0.27 2.57 0.10
1User time: CPU time spent in user mode;
2
System time: CPU time spent in the kernel or system mode;
3
Elapsed time: Total real-world time taken to execute the job;
4
Percent of CPU this job got: How much of the available CPU resources were allocated to the process, this value is commonly greater than 100% when the process runs on multiple CPU cores;
5
Maximum resident set size: The peak amount of physical memory that the process used.
6.9 Hawk’s Overheads
Hawk is developed as an asynchronous, load-balancing system that analyzes post-game replay files during periods of low server activity. It operates independently of active game servers and can be deployed on a separate server. However, we still conduct overhead assessments for reference. To answer RQ.6, we utilize an Ubuntu server equipped with 128 cores AMD EPYC 7713P 64-Core Processor and 256 GB RAM to mimic a real dedicated game server [19]. By leveraging the pre-trained model in the artifact, Hawk requires only CPU computing resources to function gracefully. Table 7 indicates the maximum overhead statistics per match, with each match analyzed concurrently. In addition to the basic overheads, we also estimate the Hawk-to-server overheads in terms of a real deployment scenario. Our platform partner reports that at most 150,000 matches are played daily. Each server needs to run at most 573 matches a day. According to our platform’s report and the public server statistics [56], player activity is at its lowest, typically 1-5 AM local time. During this timeframe, at most 144 matches need to be analyzed per hour. Thus, according to the elapsed time mentioned in Table 7, at most it needs to process 12 matches concurrently to finish processing all matches within the timeframe per server. We further conduct concurrent processing evaluation, it occupies at most 17.95% memory and 35.45% CPU usage for processing 12 matches concurrently. Therefore, each server possesses more than sufficient resources for deploying Hawk, even if several matches are hosted at the same time.
7 Discussions
Generalizability. Narrowed generalizability is a meta-problem for the anti-cheat field. Since the cheat is game-specified, the anti-cheat is hard to generalize for all games. Although we successfully implemented Hawk on one of the leading tactical shooter games [64] (a subgenre of FPS games like CS:GO), like other prior works [9, 69, 70, 2, 24, 39] discussed in Section 8, it requires further adaptation for other FPS subgenres with distinct gameplay. Hawk’s generalizability concerns on the core observations, methodologies, and workflow rather than specific features. The features presented are theoretically applicable to tactical shooters in FPS games like the CS series, CrossFire, and others. However, adaptation is needed for other subgenres, where certain elements like flashes and grenades may not apply. For instance, Hero Shooters (e.g., Overwatch, Valorant) where such features could be substituted with character-specific abilities, or Battle Royale (BR) games (e.g., Fortnite, PUBG) where elements like angles and velocities need substitution due to long time-to-kill. Since the replay files and labelings are significant for Hawk, also due to the dataset scarcity limitation, conducting further research on different subgenres to prove the generalizability of Hawk must require cooperation with the game vendors to obtain the data. We have initiated contact with several vendors, and one specializing in BR games has shown interest in cooperating. But, we have to leave specific implementation and evaluation for future work.
Manual involvement. Although Hawk has significantly increased the detection performance and efficiency, as mentioned in Section 3.2, Hawk-after-use manual verification is unavoidable. The presence of real false positives undeniably indicates the requirement for manual checks to avoid false bans. Nevertheless, with the mitigation of the limitations mentioned in Section 1, we still believe Hawk plays an important role in the FPS games’ server-side anti-cheat.
8 Related Works
Client-side detection. BotScreen [9], as the latest aimbot detector, utilizes TEE (Intel SGX) to store and run the detection model with aiming features. It leverages a self-setup server and hires 14 players to play in Deathmatch instead of the rank mode. Thus, BotScreen is limited to dataset scarcity, and its real-world functionality remains unclear, especially with more complex scenarios. BotScreen evaluates with a basic, open-source aimbot Osiris [37], which is game-specific and may not reflect the performance against more sophisticated cheats [44, 42]. VAC [12] can detect wallhack by using OpenGL and scanning such as DNS cache to verify if the suspect has purchased and cheated in a game for Counter-Strike [38, 12]. Warden [61] scans the system running programs, users’ browser history, and cookies to determine a user’s behavior for Blizzard’s games. EAC [22], BattlEye [3], and XIGNCODE3 [62] are commercial anti-cheats. They can achieve both cheat detection and prevention [53] for multiple platforms and different games after customization by monitoring and managing process privileges. Vanguard [30] is a kernel-level anti-cheat for VALORANT [51] that has full access and control over the system. The above client-side anti-cheats are limited to hardware constraints, security and privacy concerns, and client overheads. And reflect the limited generalizability problem for anti-cheat.
Client-side prevention. BlackMirror [49] applies SGX for anti-wallhack, providing the client GPU only visible entities based on local predictions. It only prevents without detecting and is limited to hardware constraints, client overheads, and targeting on a specific cheat type or subcategories. Invisibility Cloak [58] prevents computer-vision-based aimbots from performing object detection with the screen images but is ineffective against memory-based aimbots and wallhacks.
Server-side ML-based detection. Yeung et al. [69] initially proposed a Bayesian model to tune the probability threshold with basic statistics in 2006, but it is outdated for today’s advanced cheating techniques. Subsequent works [70, 2, 24, 39] used machine learning with basic statistics or limited features, which couldn’t monitor sophisticated behaviors or describe entire matches [9]. These earlier works are limited to naive features and workflow, unclear performance, and dataset scarcity. They require extra modification (specified to the game) on the game engine to collect and construct features in parallel with the game commencing. Therefore, it is also hard to generalize their method to other games.
9 Conclusion
We proposed an FPS games’ server-side anti-cheat framework Hawk, drawing inspiration from how humans identify cheaters. With promising experimental results, we have demonstrated Hawk’s real-world effectiveness, efficiency, and robustness in identifying cheaters in a modern FPS game. Ultimately, we hope Hawk can be a pathfinder for next-generation FPS anti-cheat, providing a valuable dataset and a reference for the follow-up research.
10 Ethics Considerations
The dataset only contains information including in-game operations and data, avatars’ IDs, and the banning information, which are all publicly accessible [^4] Link. We anonymize the platform’s URL and name.. We will anonymize all information that may disclose the player’s privacy, i.e., the player ID. The information in the published dataset will not directly or indirectly cause any form of impact or damage to the game and/or to its distributors, and/or to the players and/or the community. The data provider approves the release and reuse of the dataset for research purposes. No private information will be disclosed through our dataset. The studies were conducted in accordance with the local legislation and institutional requirements.
11 Compliance to Open Science Policy
We open-source Hawk and our datasets [^5] Link. The repository and dataset will be released upon acceptance. Artifacts are uploaded for the reference. Ethics statement is shown in Section 10. to support open science. We will make all links publicly available after paper acceptance. Artifacts will be submitted to the Artifact Evaluation Committee for further review.
Appendix A Training and Validation Loss for ExSPC Model
The training process on each of the datasets in ExSPC utilizes the validation set for optimization. Figure 9 and Figure 10 demonstrate the training and validation loss for each dataset respectively. It is noticeable that the individual results are not ideal. Thus, we use the multi-view design to complement one another of the subsystems.
Figure 9: Training and validation loss of ExSPC on Wallhack dataset.
Figure 10: Training and validation loss of ExSPC on Aimbot dataset.
Appendix B Temporal Features
B.1 Engagement features
Engagement features are extracted at each tick of the event occurrence. The introduction of the Engagement features is as follows:
  • Damage features:
    • tick: the tick number;
    • seconds: the time elapsed within a round in second;
    • attackerSteamID, victimSteamID: the ID of the player;
    • attackerSide, victimSide: the team side of the player;
[…]
    • attackerStrafe: the Boolean value indicates whether the attacker is moving sideways relative to their forward direction;
    • weapon: the ID of the damage-caused weapon;
    • weaponClass: the class ID of the damage-caused weapon;
    • hpDamage: the value of damage caused to the victim;
    • hpDamageTaken: the value of damage taken by the victim;
    • armorDamage: the value of the damage caused to the victim’s armor;
    • armorDamageTaken: the value of damage taken by the victim’s armor;
    • hitGroup: the ID of the damage-taken area on the body of the victim;
    • isFriendlyFire: the Boolean value indicates whether the damage is from the friendly;
    • distance: the distance between the attacker and the victim;
    • zoomLevel: the zoom level of the attacker’s weapon;
    • roundNum: the round number of the current tick.
  • Auxiliary props features: The tick, second, IDs, sides, location, view direction, and round number are similar to above;
    • flashDuration: the time elapsed for the blinding effect affects on the victim in second;
  • Offensive props features: The ID, side, location, view direction of the thrower and the prop, and round number are similar to above;
    • throwTick, destroyTick: the tick of the prop thrown or destroyed;
    • throwSeconds, destroySeconds: the time elapsed within a round in seconds at the moment of the prop thrown or destroyed;
  • Elimination features: The tick, second, IDs, sides, location, view direction, distance, weapon, weapon class, and round number are similar as above;
    • assisterSteamID: the ID of the assister;
    • assisterSide: the team side of the assister;
    • isSuicide: the Boolean value indicates whether the elimination is a suicide event;
    • isTeamkill: the Boolean value indicates whether the elimination is a team kill event;
    • isWallbang: the Boolean value indicates whether the elimination is conducted by occluder penetration;
    • penetratedObjects: the number of obstacles the bullet penetrated;
    • isFirstKill: the Boolean value indicates whether the elimination is the first kill within a round;
    • isHeadshot: the Boolean value indicates whether the elimination is a head-shot event;
    • victimBlinded: the Boolean value indicates whether the victim is blinded;
    • attackerBlinded: the Boolean value indicates whether the attacker is blinded;
    • flashThrowerSteamID: the ID of the flash-thrower;
    • flashThrowerSide: the team side of the flash-thrower;
    • noScope: the Boolean value indicates whether the elimination is conducted with no scoped;
    • thruSmoke: the Boolean value indicates whether the elimination is conducted through smoke;
    • isTrade: the Boolean value indicates whether the elimination is traded with the opponent;
  • Weapon fire features: The IDs, sides, location, view direction, weapon, weapon class, player strafe, zoom level of the attacker, and the tick, second, and round number are similar as above;
    • ammoInMagazine: the number of ammunition in the magazine;
    • ammoInReserve: the number of ammunition in the reserve;
B.2 Movement features
Movement features are extracted per tick. The IDs, sides, location, view direction of the player, and the tick, second, and round numbers are similar to those above. The rest of the introduction of the Movement features is as follows:
  • velocityX, velocityY, velocityZ: the velocity of the player in 3D space;
  • isAlive: the Boolean value indicates whether the player is alive;
  • isBlinded: the Boolean value indicates whether the player is blinded;
  • isAirborne: the Boolean value indicates whether the player is feet dangling;
  • isDucking: the Boolean value indicates whether the player is ducking;
  • isDuckingInProgress: the Boolean value indicates whether the player is ducking in progress;
  • isUnDuckingInProgress: the Boolean value indicates whether the player is standing up from ducking in progress;
  • isDefusing: the Boolean value indicates whether the player is defusing the bomb;
  • isPlanting: the Boolean value indicates whether the player is planting the bomb;
  • isReloading: the Boolean value indicates whether the player is reloading;
  • isInBombZone: the Boolean value indicates whether the player is in the bomb planting area;
  • isStanding: the Boolean value indicates whether the player is standing;
  • isScoped: the Boolean value indicates whether the player is scoped with the weapon;
  • isWalking: the Boolean value indicates whether the player is walking;
  • IsolationDegree: the distance to the calculation center of each teammate.
B.3 Economy features
Economy features are extracted per round. The introduction of the Economy features are as follows:
  • round: the number of round;
  • equipmentValueFreezetimeEnd: the value of the player’s equipment after the end of the freeze time (the interval from the commencement of a round to the moment players are granted the ability to move and fire. This period typically allows players to purchase equipment);
  • equipmentValueRoundStart: the value of the player’s equipment at the start of each round;
  • cash: the amount of money the player obtains at the end of each round;
  • cashSpendTotal: the amount of money the player spends within each round.
Appendix C Structured Features
In this section, the extracted data from demos are further processed to create strong-expressiveness features that can indicate the sense and performance of a player clearly. For notations used in the formulas throughout this section, please refer to Table 9.
Table 9: Notations and descriptions of structured features.
Notation Description Notation Description
$i$ the current player $\textbf{E}_{\textit{x}}^{a\rightarrow b}[k]$ the $k^{th}$ type x event conducted by player a to b
$p,p^{*}$ a particular player, any player #()#\#($\bullet$)# ( ∙ ) the number of counts of the parameter
$o,o^{*}$ a particular opponent, any opponent $wp(\bullet)$ the weapon used by a player
$a,a^{*}$ a particular ally, any ally $hg(\bullet)$ the body part hit by a player
$ts$ through smoke $\textbf{{t}}(\bullet)$ the moment of the event occurs in second
$p(\bullet)$ the number of occluders penetrated $mag(\bullet)$ the number of bullets remain in a player’s magazine
$\Delta(\bullet)$ the change in the parameter $dist(x,y)$ the distance range between entities x and y
$\textbf{ELAPSE}(\bullet)$ the duration from the start to the end of an event $\textbf{HEG}(\bullet)$ high-explosive grenade is applied by a player
$\textbf{INC}(\bullet)$ incendiary or Molotov is applied by a player $\textbf{SMK}(\bullet)$ smoke grenade is applied by a player
$\textbf{FLS}(\bullet)$ flash grenade is applied by a player $\textbf{FK}(\bullet)$ first kill is accomplished by a player
Figure 11: Illustrations of reaction and adjustment features. The black rectangle overlay in c2 hides the player’s ID.
C.1 Aiming features
Durations
Figure 11 illustrates the three distinct moments that form three different durations. In terms of the moments, $\textbf{{t}}_{0}$ denotes the moment that the current player initially spots one opponent. In Figure 11, a1 and a2 present a possible set of states at $\textbf{{t}}_{0}$. In which a1 represents a 3-dimensional world in the game map, where the opponent happens to appear at the edge of the current player’s field of view (FOV). And a2 demonstrates the 2-dimensional state correspondingly which is typically displayed through the player’s monitor, where the opponent also happens to be spotted at the edge of the current player’s POV. $\textbf{{t}}_{1}$ denotes the moment that the current player notices the hostile movement and initiates responsive action (i.e., open fire). Note that at $\textbf{{t}}_{1}$ the current firing event may or may not be equivalent to a damage event upon the opponent. $\textbf{{t}}_{2}$ denotes the moment that the current player performs a damage event upon the opponent within one period of valid engagement for the first time. In our case, the threshold for resetting the valid engagement period detection (i.e., no opponent in sight, no damage or firing event) is set as 10 seconds. In Figure 11, it demonstrates a series of generic circumstances that often happen in games where the positions of the current player and the opponent vary along from $\textbf{{t}}_{0}$ to $\textbf{{t}}_{2}$. And it is common that sometimes $\textbf{{t}}_{1}$ and $\textbf{{t}}_{2}$ are overlapped. In terms of durations, the elapsed period from $\textbf{{t}}_{0}$ to $\textbf{{t}}_{1}$ is defined as reaction time (rat). The elapsed period from $\textbf{{t}}_{1}$ to $\textbf{{t}}_{2}$ is defined as adjustment time (ajt). The elapsed period from $\textbf{{t}}_{0}$ to $\textbf{{t}}_{2}$ is defined as duration time (drt). The average and variance calculations are applied after obtaining the data. The average performance of all samples is obtained through the average, and the degree of dispersion of the samples is represented by the variance. The integration of average and variance is able to represent the overall performance and stability of the current player within a match.
rat average. Denotes the average performance regarding rat of the current player among one match. The lower this feature is, the faster the player reacts.
$\mathbf{F_{\textbf{rat-avg}}(i)}=\overline{RAT}=\frac{\sum_{j=1}^{N}RAT_{j}}{N}$ (11)
where $\text{where }N\text{ denotes }\#(RAT)\text{ within a match.}$.
rat variance. Denotes the degree of dispersion of rat distributions of the current player among one match.
$\mathbf{F_{\textbf{rat-var}}(i)}=Var(RAT)=\frac{\sum_{j=1}^{N}\left(RAT_{j}-\overline{RAT}\right)^{2}}{N}$ (12)
where $\text{where }N\text{ denotes }\#(RAT)\text{ within a match.}$.
ajt average. Denotes the average performance regarding ajt of the current player among one match. The lower this feature is, the faster the player adjusts.
$\mathbf{F_{\textbf{ajt-avg}}(i)}=\overline{AJT}=\frac{\sum_{j=1}^{N}AJT_{j}}{N}$ (13)
where $\text{where }N\text{ denotes }\#(AJT)\text{ within a match.}$.
ajt variance. Denotes the degree of dispersion of ajt distributions of the current player among one match.
$\mathbf{F_{\textbf{ajt-var}}(i)}=Var(AJT)=\frac{\sum_{j=1}^{N}\left(AJT_{j}-\overline{AJT}\right)^{2}}{N}$ (14)
where $\text{where }N\text{ denotes }\#(AJT)\text{ within a match.}$.
drt average. Denotes the average performance regarding drt of the current player among one match. The lower this feature is, the faster the player hits.
$\mathbf{F_{\textbf{drt-avg}}(i)}=\overline{DRT}=\frac{\sum_{j=1}^{N}DRT_{j}}{N}$ (15)
drt variance. Denotes the degree of dispersion of drt distributions of the current player among one match.
$\mathbf{F_{\textbf{drt-var}}(i)}=Var(DRT)=\frac{\sum_{j=1}^{N}\left(DRT_{j}-\overline{DRT}\right)^{2}}{N}$ (16)
where $\text{where }N\text{ denotes }\#(DRT)\text{ occurrences within a match.}$.
Figure 12: Complete box plots of comparative visualization between honest players and cheaters on all structured features under Mann-Whitney U (Wilcoxon Rank-Sum) test with P-Value ascending (distinction descending) order.
Velocity
velocity average. Denotes the average mouse velocity of the current player among one match between $\textbf{{t}}_{0}$ and $\textbf{{t}}_{2}$, through which way two features of distance and time period can be highly compressed into one feature. The higher the feature is, the faster the player hits the targets.
$V=\frac{dist(i,o)}{DRT_{i}}$ (17)
where $\text{where }o\text{ denotes the victim in the specified }DRT\text{.}$
$\mathbf{F_{\textbf{v-avg}}(i)}=\overline{V}=\frac{\sum_{j=1}^{N}V_{j}}{N}$ (18)
where $\text{where }N\text{ denotes }\#(DRT)\text{ within a match.}$.
velocity variance. Denotes the degree of dispersion of velocity distributions of the current player among one match.
$\mathbf{F_{\textbf{v-var}}(i)}=Var(V)=\frac{\sum_{j=1}^{N}\left(V_{j}-\overline{V}\right)^{2}}{N}$ (19)
where $\text{where }N\text{ denotes }\#(DRT)\text{ within a match.}$.
Angles
Figure 11 illustrates the three different angles, in which $\theta$ denotes divergence of aiming is extracted by Liu et al. [39] and represents the initial angular divergence between the aiming position and the opponent position. However, the time complexity for obtaining such a feature is too expensive to be accepted for industrial implementation. Therefore we extracted the following features which not only preserve strong expressiveness but also significantly reduce the time cost, as well as innovatively segment each engagement process into three stages. $\delta$ represents the reaction angle (raa), which reflects the angle of two cross-hair directions at $\textbf{{t}}_{0}$ and $\textbf{{t}}_{1}$ respectively. $\gamma$ represents the adjustment angle (aja), which, similarly reflects the angle of two cross-hair directions at $\textbf{{t}}_{1}$ and $\textbf{{t}}_{2}$ respectively. The larger $\delta$ and $\gamma$ are the longer distance the cross-hair travels, which ultimately results in performance divergence. Similarly, the average and variance operation is applied accordingly.
raa average. Denotes the average performance regarding raa of the current player among one match. The lower the feature is, the smaller the angle and the less effort the player aims for and puts into reaching their targets.
$\mathbf{F_{\textbf{raa-avg}}(i)}=\overline{RAA}=\frac{\sum_{j=1}^{N}RAA_{j}}{N}$ (20)
where $\text{where }N\text{ denotes }\#(RAA)\text{ within a match.}$.
raa variance. Denotes the degree of dispersion of raa distributions of the current player among one match.
$\mathbf{F_{\textbf{raa-var}}(i)}=Var(RAA)=\frac{\sum_{j=1}^{N}\left(RAA_{j}-\overline{RAA}\right)^{2}}{N}$ (21)
where $\text{where }N\text{ denotes }\#(RAA)\text{ within a match.}$.
aja average. Denotes the average performance regarding aja of the current player among one match. The lower the feature is, the smaller angle and the less effort the player fires for and puts into hitting their targets.
$\mathbf{F_{\textbf{aja-avg}}(i)}=\overline{AJA}=\frac{\sum_{j=1}^{N}AJA_{j}}{N}$ (22)
where $\text{where }N\text{ denotes }\#(AJA)\text{ within a match.}$.
aja variance. Denotes the degree of dispersion of aja distributions of the current player among one match.
$\mathbf{F_{\textbf{aja-var}}(i)}=Var(AJA)=\frac{\sum_{j=1}^{N}\left(AJA_{j}-\overline{AJA}\right)^{2}}{N}$ (23)
where $\text{where }N\text{ denotes }\#(AJA)\text{ within a match.}$.
C.2 Firing features
Inertial shot [39] percentage (isp)
Inertial shot denotes that after an elimination is conducted by a player, who may apply several extra firings due to the inertia. isp represents the ratio of the number of Inertial shot times to the number of eliminations. $\alpha$ is a threshold and set as 0.15 seconds in our case.
$\operatorname{\textbf{IS}}\left(\textbf{E}_{\textit{kill}}^{i\rightarrow o^{*}}[k]\right)=\left\{\begin{array}[]{ll}1,&\exists\textbf{E}_{\textit{fire}}^{i}\in\left(\textbf{{t}}\left(\textbf{E}_{\textit{kill}}^{i\rightarrow o^{*}}[k]\right),\textbf{{t}}\left(\textbf{E}_{\textit{kill}}^{i\rightarrow o^{*}}[k]\right)+\alpha\right)\\ 0,&\text{otherwise}\end{array}\right.$
$\mathbf{F_{\textbf{isp}}(i)}=\frac{\sum_{k}\operatorname{\textbf{IS}}\left(\textbf{E}_{\textit{kill}}^{i\rightarrow o^{*}}[k]\right)}{\sum_{n}\textbf{E}_{\textit{kill}}^{i\rightarrow o^{*}}[n]}$ (25)
First hit percentage (fhp)
Denote how accurate a player is while spotting an opponent and reacting to fire. Event reload represents the action of changing the magazine while the firing is performed and the bullet is consumed. Fire Round denotes the unit of the firing round. fhp refers to the ratio of the number of the first firing event equivalent to a damage event while spotting at least one opponent within a time threshold $\varepsilon$ to the number of times that the aforementioned event occurs while disregarding the equivalency and within the same time threshold $\varepsilon$. In our case, $\varepsilon$ is set as 5 seconds (i.e., 640 ticks).
$\textbf{E}_{\textit{reload}}^{i}=\Delta\left(\textit{mag}(i)\right)>0\land\textbf{{t}}(\textbf{E}_{\textit{fire}}^{i}[k])-\textbf{{t}}(\textbf{E}_{\textit{fire}}^{i}[k+1])>0$ (26)
$\textbf{FR}[i]={\{\textbf{E}_{\textit{fire}}^{i}[i]\mid\exists\textbf{E}_{\textit{reload}}^{i}\lor\textbf{{t}}(\textbf{E}_{\textit{fire}}^{i}[i][k+1])-\textbf{{t}}(\textbf{E}_{\textit{fire}}^{i}[i][k])>\varepsilon\}}$ (27)
$\mathbf{F_{\textbf{fhp}}(i)}=\frac{\#\left(\{\textbf{FR}[i][0]\mid\forall\textbf{FR}[i][0]=\textbf{E}_{\textit{damage}}^{i\rightarrow o^{*}}\}\right)}{\#\left(\textbf{FR}[i][0]\right)}$ (28)
Precision (precis)
Denote the performance of firing accuracy. precis refers to the ratio of the number of damage caused by weapons to the total number of fires applied.
$\mathbf{F_{\textbf{precis}}(i)}=\frac{\sum_{k}\textbf{E}_{\textit{damage}}^{i\rightarrow o^{*}}[k]}{\sum_{n}\textbf{E}_{\textit{fire}}^{i}[n]}$ (29)
Strafing hit percentage (shp)
Denote fire accuracy while strafing. Strafe allows a player to keep the view focused on an opponent while moving in a different direction. shp refers to the ratio of the number of damage caused while strafing to the number of damage caused.
$\mathbf{F_{\textbf{shp}}(i)}=\frac{\sum_{k}\textbf{E}_{\textit{damage}\land\textit{strafe}}^{i\rightarrow o^{*}}[k]}{\sum_{n}\textbf{E}_{\textit{damage}}^{i\rightarrow o^{*}}[n]}$ (30)
Hit-group distribution (hgd) variance
Hit group represents the body part that suffers damage, typically hit groups can be categorized as follows: chest, generic, head, neck, left arm, right arm, left leg, right leg, and stomach. hgd variance denotes the degree of dispersion of hit group distributions for the current player among one match.
$HG_{j}=\sum_{k}\textbf{E}_{\textit{damage}\land\textit{hg(j)}}^{i\rightarrow o^{*}}[k]$ (31)
$\overline{HG}=\frac{\sum_{j=chest}^{stomach}HG_{j}}{\sum_{k}\textbf{E}_{\textit{damage}}^{i\rightarrow o^{*}}[k]}$ (32)
$\mathbf{F_{\textbf{hgd}}(i)}=Var(HG)=\frac{\sum_{j=chest}^{stomach}\left(HG_{j}-\overline{HG}\right)^{2}}{\sum_{n}\textbf{E}_{\textit{damage}}^{i\rightarrow o^{*}}[n]}$ (33)
[…]
Special hit Ratio (shr)
Special hit denotes hit with sniper between 50 and 150 meters (SH50-150), head shot with sniper between 40 and 170 meters (SHS40-170), hit with normal weapon farther than 800 meters (NH800+) and headshot with normal weapon farther than 700 meters (NHS700+). In FPS games, hitting processes include using different weapons and shooting from different distances. In our game environment, the aforementioned four types of hits are considered to be potentially cheating combinations. shr denotes the ratio of the sum of the number of the aforementioned four types of special hit to the number of hits in one match.
$\mathbf{\textbf{SH}_{50-150}(i)}=\sum_{k}\textbf{E}_{\textit{damage}\land{sniper}\land{(50<=dist(i,o)}<=150)}^{i\rightarrow o^{*}}[k]$
$\mathbf{\textbf{SHS}_{40-170}(i)}=\sum_{k}\textbf{E}_{\textit{headshot}\land{sniper}\land{(40<=dist(i,o)}<=170)}^{i\rightarrow o^{*}}[k]$
$\mathbf{\textbf{NH}_{800+}(i)}=\sum_{k}\textbf{E}_{\textit{damage}\land{normal}\land{(dist(i,o)}>=800)}^{i\rightarrow o^{*}}[k]$ (36)
$\mathbf{\textbf{NHS}_{700+}(i)}=\sum_{k}\textbf{E}_{\textit{headshot}\land{normal}\land{(dist(i,o)}>=700)}^{i\rightarrow o^{*}}[k]$ (37)
$\mathbf{F_{\textbf{shr}}(i)}=\frac{\mathbf{\textbf{SH}_{50-150}(i)}+\mathbf{\textbf{SHS}_{40-170}(i)}+\mathbf{\textbf{NH}_{800+}(i)}+\mathbf{\textbf{NHS}_{700+}(i)}}{\sum_{k}\textbf{E}_{\textit{damage}}^{i\rightarrow o^{*}}[k]}$
C.3 Elimination features
Time to kill (ttk)
Denote the duration from the moment of the first damage caused to an opponent to the moment of the elimination of the opponent. In order to control the ttk to be each time one player attacks one enemy, the threshold of one statistical period of ttk is set to be 10 seconds. For instance, if the attacking period time is over 10 seconds without killing the victims, this action time will not be recorded as ttk.
$\mathbf{F_{\textbf{ttk}}(i)}=\operatorname{\textbf{ELAPSE}}\left(\textbf{E}_{\textit{damage}}^{i\rightarrow o}[0],\textbf{E}_{\textit{kill}}^{i\rightarrow o}\right)$ (39)
ttk average. Denotes the average time regarding ttk of the current player among one match. The lower this feature is, the shorter the time the player kills victims.
$\mathbf{F_{\textbf{ttk-avg}}(i)}=\overline{TTK}=\frac{\sum_{j=1}^{N}TTK_{j}}{N}$ (40)
where $\text{where }N\text{ denotes }\#(TTK)\text{ within a match.}$.
ttk variance. Denotes the degree of dispersion of ttk distributions of the current player among one match.
$\mathbf{F_{\textbf{ttk-var}}(i)}=Var(TTK)=\frac{\sum_{j=1}^{N}\left(TTK_{j}-\overline{TTK}\right)^{2}}{N}$ (41)
where $\text{where }N\text{ denotes }\#(TTK)\text{ within a match.}$.
First-blood percentage (fbp)
first-blood denotes the very first elimination in one round. fbp denotes the ratio of the number of first-blood to the number of rounds in one match.
$\mathbf{F_{\textbf{fkp}}(i)}=\frac{\#\left(\textbf{FK}(i)\right)}{\#\left(round\right)}$ (42)
Onetap percentage (otp)
Onetap denotes a head-shot fire event on an opponent that results in elimination without the use of sniper rifles, and where the inflicted damage is equal to or greater than 100 (the full health of a player). The otp represents the ratio of the number of Onetap instances to the number of times damage is inflicted by normal weapons.
$\operatorname{\textbf{ONETAP}}\left(\textbf{E}_{\textit{kill}}^{i\rightarrow o^{*}}[k]\right)=\left\{\begin{array}[]{ll}1,&\left(\textbf{E}_{\textit{fire}}^{i}=\textbf{E}_{\textit{kill}}^{i\rightarrow o^{*}}[k]\right)\land\left(\textbf{E}_{\textit{damage}}^{i\rightarrow o^{*}}\geq 100\right)\\ 0,&\text{otherwise}\end{array}\right.$
$\mathbf{F_{\textbf{otp}}(i)}=\frac{\sum_{k}\operatorname{\textbf{ONETAP}}\left(E_{\text{kill}}^{i\rightarrow o^{*}}[k]\right)}{\sum_{n}\textbf{E}_{\textit{damage}\land wp(i)\neq sniper\ rifle}^{i\rightarrow o^{*}}[n]}$ (44)
Occluder-penetration index (opi)
Occluder-penetration denotes the elimination event that happens when the victim is eliminated behind one layer or more layers of occluders by a penetrated bullet. The occluders are typically walls, players (namely wall bang), or smoke. opi refers to a weighted index that reflects the extent of the average occluder-penetration eliminations conducted by the player. The use of the number of rounds within the division aims to eliminate bias brought by that. In our case, the weights in $\alpha_{i}$ are 0.5, 1, 2, 0.5 respectively.
$\mathbf{F_{\text{opi}}(i)}=\frac{\mathbf{W_{i}}\alpha_{i}}{\#\left(\textit{round}\right)}$ (45)
$\text{where}\ \alpha_{i}=\begin{bmatrix}w_{1}\\ w_{2}\\ w_{3}\\ w_{4}\end{bmatrix},$
${\mathbf{W_{i}}}=\begin{bmatrix}\sum_{k}\textbf{E}_{\textit{kill}\land p(1)}^{i\rightarrow o^{*}}[k]&\sum_{n}\textbf{E}_{\textit{kill}\land p(2)}^{i\rightarrow o^{*}}[n]&\sum_{m}\textbf{E}_{\textit{kill}\land p(>2)}^{i\rightarrow o^{*}}[m]&\sum_{j}\textbf{E}_{\textit{kill}\land ts}^{i\rightarrow o^{*}}[j]\end{bmatrix}$
Blind kill percentage (bkp)
Denote the percentage of a player performing elimination while blinded. bkp refers to the ratio of the sum of elimination events where the attackers suffer a blinded effect to the number of eliminations.
$\mathbf{F_{\textbf{bkp}}(i)}=\frac{\sum_{k}\textbf{E}_{\textit{kill}\land\textit{blind}}^{i\rightarrow o^{*}}[k]}{\sum_{n}\textbf{E}_{\textit{kill}}^{i\rightarrow o^{*}}[n]}$ (46)
Critical hit percentage (chp)
Denote the critical hit rate. Critical hit represents a successful damage event that applies to a particular hit group (typically, head) that deals more damage than a normal blow. chp refers to the ratio of the number of critical hits to the number of damage caused.
$\mathbf{F_{\textbf{chp}}(i)}=\frac{\sum_{k}\textbf{E}_{\textit{headshot}}^{i\rightarrow o^{*}}[k]}{\sum_{n}\textbf{E}_{\textit{damage}}^{i}[n]}$ (47)
Average kill per round (akpr)
Denote the average number of eliminations per round by a player in a match.
$\mathbf{F_{\textbf{akpr}}(i)}=\frac{\sum_{k}\textbf{E}_{\textit{kill}}^{i\rightarrow o^{*}}[k]}{\#\left(round\right)}$ (48)
C.4 Props utilization features
Flash efficiency index (fei)
Denote how well a player utilizes the flash grenade. fei refers to the ratio of the total affection duration on the opponents minus that on the friendlies to the total affection duration.
$\mathbf{F_{\textbf{fei}}(i)}=\frac{\sum_{k}\operatorname{\textbf{ELAPSE}}\left(\textbf{E}_{\textit{blind}}^{i\rightarrow o^{*}}[k]\right)-\sum_{m}\operatorname{\textbf{ELAPSE}}\left(\textbf{E}_{\textit{blind}}^{i\rightarrow a^{*}}[m]\right)}{\sum_{n}\operatorname{\textbf{ELAPSE}}\left(\textbf{E}_{\textit{blind}}^{i\rightarrow p^{*}}[n]\right)}$ (49)
Props utilization index (pui)
Denote how frequently a player utilizes props such as a high-explosive grenade, Molotov or incendiary, smoke grenade, or flash grenade. pui refers to the ratio of the sum of a player and the number of times different grenades are used to a fixed constant which reflects the maximum number of grenades used within this match for all players. $\rho$ denotes the maximum number of props a player is allowed to carry per round. In our case, $\rho$ is 5.
$\mathbf{F_{\textbf{pui}}(i)}=\frac{\#\left(\textbf{HEG}(i)\right)+\#\left(\textbf{INC}(i)\right)+\#\left(\textbf{SMK}(i)\right)+\#\left(\textbf{FLS}(i)\right)}{\#\left(player\right)\times\#\left(round\right)\times\rho}$ (50)
Appendix D Sense and Performance Features Classification
All features presented in this section are recombinations of features from the first two subsystems, categorized into two main types, i.e., sense features and performance features. For a detailed description of each feature, please refer to Appendix B and Appendix C. The following merely enumerates the specific categorizations of the features and some brief descriptions.
D.1 Sense Features
Sense features indicate a player’s in-game cognition, emphasizing their strategic acumen and tactical foresight.
  • Temporal Sense Features:
    • Economy features: Financial transactions and asset management, mirroring strategic resource allocation.
    • Movement features: Spatial patterns, positioning, and navigation dynamics, indicating tactical awareness.
    • Flash features: Strategic deployment of flash grenades during in-game engagements.
    • Grenade features: Utilization of assorted grenades, highlighting understanding of map control.
  • Structured Sense Features:
    • Flash efficiency index (fei): Metrics gauging flash grenade efficiency.
    • Occluder-penetration index (opi): Achieve eliminations at specific locations, demonstrating proficiency with the map.
    • Inertial shot percentage (isp): As a normal human being’s habitual firing after destroying an opponent.
    • Blind kill percentage (bkp): Eliminations executed under visual impairment.
    • Props utilization index (pui): Metrics showcasing prop usage across the whole match.
D.2 Performance Features
Performance features encapsulate a player’s in-game efficacy, concentrating on engagement aptitude, aiming precision, and mechanism mastery.
  • Temporal Performance Features:
    • Elimination Features: Efficiency of eliminations and some special, difficult eliminations.
    • Weapon Fire Features: Proficiency metrics across weapon categories, illuminating combat adeptness.
    • Damage Features: The player’s performance with respect to inflicting effective damage in-game.
  • Structured Performance Features: The remaining structured features in Appendix C.
Figure 13: Structured Features PCA results on both datasets.
Appendix E PCA on Structured Features
In Section 4.2.1, we have demonstrated the distribution distinction between cheaters and normal players. In the following section, we will demonstrate the PCA visualization on the structured features to display their orthogonality on both Aimbot and Wallhack datasets. Figure 13 illustrates the structured features PCA results for both datasets regarding the cheaters and normal players respectively. Horizontal and vertical axes represent the first principal component (PC1) and the second principal component (PC2), respectively. These principal components capture the directions of maximum variance in the data, showing where the data varies most. The unit circle in a correlation circle plot is a circle with a radius of 1, serving as a reference. Within this circle, the vectors (represented as arrows) demonstrate the magnitude and direction of the original variables’ associations with the principal components. The arrow represents feature vectors, where the direction and length of each arrow indicate the relationship of the original data features with the principal components. The direction of the arrow shows the correlation direction of the feature with the principal components. If two arrows are pointing in similar directions, it means their corresponding features exhibit similar patterns of variation in the dataset. Length indicates the impact or correlation strength of that feature on the corresponding principal component. Longer arrows suggest that the feature contributes more significantly to the variance captured by that principal component, and vice versa. It is noticeable that within the same plot, most of the features are distinct from others by obtaining different directions and lengths of the arrow. We can also observe that the direction and length of the arrows across Aimbot and Wallhack datasets are distinctive, indicating that the data distributions of the corresponding features are statistically different. Additionally, there are differences in the direction and length of the arrows between cheaters and normal players within the same or across different datasets. Above all, the original features can not be replaced by PCA.
Appendix F Evaluation Metrics
The following are the detailed evaluation matrices that are used in Section 6.3.
True Positives (TP), shown in Equation 51, represents instances where the model accurately predicts the occurrence of cheating behavior.
$TP=\sum_{i=1}^{n}I(y_{i}=1\land\hat{y}_{i}=1)$ (51)
True Negatives (TN), represented in Equation 52, denotes instances where the model correctly identifies the absence of cheating behavior.
$TN=\sum_{i=1}^{n}I(y_{i}=0\land\hat{y}_{i}=0)$ (52)
False Positives (FP), represented in Equation 53, encompass instances where the model erroneously predicts the presence of cheating behavior.
$FP=\sum_{i=1}^{n}I(y_{i}=0\land\hat{y}_{i}=1)$ (53)
False Negatives (FN), represented in Equation 54, are instances where the model inaccurately predicts the absence of cheating behavior.
$FN=\sum_{i=1}^{n}I(y_{i}=1\land\hat{y}_{i}=0)$ (54)
In these equations, $y_{i}$ denotes the true label, $\hat{y}_{i}$ signifies the predicted label, and $I(\cdot)$ represents the indicator function.
Recall, represented in Equation 55, measures the proportion of actual positive instances that were correctly identified by the model.
$Recall=\frac{TP}{TP+FN}$ (55)
Accuracy, as shown in Equation 56, gives a holistic measure of the model’s performance across both positive and negative classes.
$Accuracy=\frac{TP+TN}{TP+TN+FP+FN}$ (56)
AUC-ROC, denoted in Equation 57, is the area under the receiver operating characteristic curve, indicating the model’s capability to differentiate between positive and negative classes. It is computed by integrating the ROC curve, as shown below, where $TPR$ (True Positive Rate) is synonymous with Recall, and $FPR$ (False Positive Rate) is defined as $\frac{FP}{FP+TN}$.
[…]
NPV, shown in Equation 58, is the proportion of actual negative instances among all instances predicted as negative.
$NPV=\frac{TN}{TN+FN}$ (58)

### v2

@@
 To address these limitations, the paper proposes Hawk, a server-side FPS anti-cheat framework for the popular game CS:GO.
-Hawk utilizes machine learning techniques to mimic human experts’ identification process, leverages novel multi-view features, and it is equipped with a well-defined workflow.
-The authors evaluate Hawk with the first large and real-world datasets containing multiple cheat types and cheating sophistication, and it exhibits promising efficiency and acceptable overheads, shorter ban times compared to the in-use anti-cheat, a significant reduction in manual labor, and the ability to capture cheaters who evaded official inspections.
-1 Introduction
-The gaming industry has experienced substantial growth recently, with the global game market generating \$184.0 billion in 2023 and projected to reach \$218.7 billion in 2024 [47, 48].
+Hawk utilizes machine learning techniques to mimic human experts’ identification process, leverages novel multi-view features, and is equipped with a well-defined workflow.
+Hawk is evaluated with the first large and real-world datasets containing multiple cheat types and cheating sophistication, and it exhibits promising efficiency and acceptable overheads, shorter ban times, higher recall and similar false positive rate compared to the in-use anti-cheat, and the ability to capture cheaters who evaded official inspections.
+I Introduction
+The gaming industry has experienced substantial growth recently, with the global games market generating \$187.7 billion in 2024 and projected to generate \$213.3 billion by 2027 [1].
 However, cheating in online games poses a significant threat to the integrity of gaming experiences and disrupts gameplay balance.
-In 2023 alone, cheating led to an estimated \$29 billion in losses, with 78% of gamers deterred from playing due to cheating [32].
-In particular, first-person shooter (FPS) games, which account for 20.9% of total game sales and rank second among all genres [63], are particularly targeted by cheat developers seeking monetary gains due to their competitive and multiplayer nature.
-The most prevalent forms of cheating involve the use of aimbots to enhance a player’s precision and wallhacks to reveal opponents’ positions, thereby compromising game integrity [4, 40, 9].
-On realizing the critical importance of game security [25], both academia and industry have proposed many anti-cheats solutions, applied either on the client-side [9, 58, 49] or server-side [69, 24, 70, 2, 39, 27].
+In 2023 alone, cheating led to an estimated \$29 billion losses, with 78% of players discouraged from playing due to cheating [2].
+In particular, first-person shooter (FPS) games, which account for 20.9% of total game sales [3], are particularly targeted by cheat developers because their stringent latency requirements make key computations on the client side, creating fertile ground for cheats to exploit.
+The most prevalent forms of cheating involve the use of aimbots to aid aiming and wallhacks to reveal opponents’ positions, thus compromising game integrity [4, 5, 6].
+Realizing the critical importance of game security [7], both academia and industry have proposed many anti-cheats solutions, applied either on the client side [6, 8, 9] or server side [10, 11, 12, 13].
 However, current solutions have the following limitations.
-Client-side solutions face hardware constraints, bring security and privacy concerns, or incur unacceptable system overheads.
-Two state-of-the-art systems [9, 49] require specific Trusted Execution Environments (TEEs) like Intel SGX.
-This dependence on particular hardware, which was deprecated in 2021 [5, 31], limits compatibility and practicality for industrial use.
-Some solutions involve scanning hard drives and gaining root privileges, posing risks of unauthorized access and personal data leakage [45, 43, 66].
-Vulnerabilities, such as RCE exploits [8] and CVEs [15, 16, 17, 18], demonstrate the potential for system instability and privacy breaches.
-The others rely on continuous surveillance or frequent patching to counter varying cheat signatures [22, 3, 35, 58], leading to increased client-side system overheads.
-On the other hand, the existing server-side solutions use basic features and share low recall and accuracy.
-For instance, some solutions [69, 24, 70, 2, 39, 27] use features like winning rates or playtime, which are insufficient for detecting complex cheats such as aimbots or wallhacks.
-These methods also lack robust workflows to examine false positives, especially in noise-rich, real-world datasets.
-Additionally, they show low recall and accuracy in the real world, with extended ban cycles taking days or weeks [33, 50].
-Incidents like the 2024 APEX Legends tournament hack [59] highlight the inefficacy of existing solutions under commercial anti-cheats such as Easy Anti-Cheat (EAC) [22].
-Moreover, both client-side and server-side solutions suffer from a lack of comprehensive datasets, often relying on limited or synthetic data [9, 49, 24, 70, 2, 39], which may not represent real-world cheating scenarios.
-The simulated datasets lack varying levels of cheating sophistication (explained in Section 2.2), which could result in performance degradation or even failure in complex real-world scenarios.
+Client-side solutions face hardware constraints, security and privacy concerns, or unacceptable system overheads.
+Two state-of-the-art systems [6, 9] require a specific Trusted Execution Environment (TEE), Intel SGX.
+This dependence on particular hardware, which was deprecated in 2021 [14, 15], limits the compatibility and practicality for industrial use.
+Some industrial solutions [16, 17, 18] involve scanning hard drives and gaining root privileges, posing risks of unauthorized access and personal data leakage [19, 20, 21].
+Vulnerabilities, e.g., RCE exploits [22] and CVEs [23, 24, 25, 26], demonstrate the potential for system instability and privacy breaches.
+And most client-side works rely on concurrent surveillance [9, 8] leading to increased client-side overheads, or frequent patching to counter varying cheat signatures [16, 17, 18, 6], which requires extra engineering efforts and is considered to be less effective [13] since it only identifies known and deprecated cheats.
+Importantly, client-side data and decision results are unreliable and easily tampered with [27, 28, 29, 13].
+Thus, the industry has shifted towards server-side anti-cheat solutions [27, 28].
+However, existing server-side solutions use basic features and have low recall and accuracy.
+For instance, previous works [10, 11, 13, 30, 31] use naive features like winning rates, playtime, or count shots on an invisible bait which are insufficient or inapplicable for detecting today’s complex cheats or specific cheat types.
+These works leverage thresholds to classify, which is not robust enough for practical use, especially in noise-rich, real-world datasets.
+Additionally, some commercial anti-cheats integrate such server-side approaches [16].
+Yet, they show low recall and accuracy, with extended ban cycles and increasing community grievances [32, 33].
+Incidents like the 2024 APEX Legends Tournament hack [34] and the study [30] highlight the ineffectiveness of existing commercial anti-cheats, e.g., EAC [16] and VAC [18].
+More broadly, prior research on anti-cheat techniques suffers from the lack of comprehensive real-world datasets, with most studies relying on limited or synthetic data [6, 9, 10, 11, 12, 13], which are hard to capture diverse cheating scenarios and varying levels of sophistication (explained in Section II-B).
+This could result in performance degradation or even failure in complex real-world scenarios.
 To address the aforementioned limitations, we propose Hawk, a server-side approach designed to analyze data transmitted to the server, thereby avoiding the inherent limitations of client-side solutions.
 Hawk introduces a comprehensive set of multi-view features and a robust framework to meticulously monitor players’ points of view, statistics, and behavioral consistencies.
-This approach includes a corresponding workflow aimed at shortening the ban cycle and considering the occurrence of false positives.
-Hawk achieves high detection recall and acceptable accuracy across large, real-world datasets with varying levels of cheating sophistication.
-The key innovation of Hawk lies in its ability to mimic how human experts identify cheaters by focusing on three main aspects: analyzing the player’s point of view, statistically evaluating the player’s performance, and assessing the player’s gaming sense and performance consistency.
-Based on these aspects, we design structured and temporal features that accurately represent player behaviors.
-Hawk reproduces the human identification process using Long Short-Term Memory (LSTM) and attention mechanisms, ensemble learning, and deep learning networks, respectively.
+This approach includes a corresponding workflow that aims to shorten the ban cycle and considers the occurrence of false positives.
+Hawk achieves high detection recall and acceptable accuracy in large real-world data sets with varying levels of sophistication of cheating.
+Table I listed the comparison between Hawk and other detection methods.
+TABLE I: Detection works comparison.
+‘Availability’: public availability of technical details (T), source code (C), and datasets (D), red and green indicate unavailable and available.
+‘Dynamic Tuning’: ability to adjust different anti-cheat demands (e.g., prioritize different metrics, introduced in Section V-D) without retraining the entire system.
+‘Data Extraction-Game Coupling Degree’: difficulty of acquiring the game’s raw data, ‘high coupling’ indicates the process is highly coupled with the game logic (e.g., modify engine, hooking process), making it challenging to generalize or reuse across different game versions.
+Side Method Work(s) Availability
+Detection Feature
+Comprehensiveness
+Data Extraction-Game
+Coupling Degree
+Real-world Functionality
+Dataset Deployment Constraint Robustness Dynamic Tuning Performance
+Client ML-based [6] T, C, D Single aspect & data type High Coupling SSL1 TEE (Intel SGX) Unknown Unknown2
+Commercial signature-based [16, 17, 18, 35, 36] T, C, D Unknown High Coupling Real-world OS and kernel drivers Unknown Unknown Low recall, acceptable FPR
+Server Statistical/Threshold-based [10, 11, 13, 30, 31] T, C, D Single aspect & data type High Coupling SSL Concurrent overheads Unknown Unknown
+ML-based [12] T, C, D Single aspect & data type High Coupling SSL Concurrent overheads Unknown Unknown
+Hawk T, C, D Multiple aspects & data types Low Coupling Real-world None CE3, N4 High recall, similar FPR
+1SSL: Simulated, small-scale, low complexity; 2Unknown: Never been tested or disclosed; 3CE: Robust to cheat evolution (an adversarial attack in Section VI-G); 4N: Robust to noise.
+The key innovation of Hawk lies in its ability to mimic how human experts identify cheaters by focusing on three dimensions: analyzing the player’s point of view, statistically evaluating the player’s performance, and assessing the player’s gaming sense and performance consistency.
+Building on these dimensions, we design structured and temporal features that precisely describe player behaviors.
+These features are selected from hundreds of candidates to ensure that they effectively represent diverse behavioral patterns.
+Hawk replicates the human identification process by leveraging machine learning techniques, including Long Short-Term Memory (LSTM) with attention mechanisms, ensemble learning, and deep neural networks, each tailored to specific aspects of the analysis.
+Hawk addresses the challenges of integrating heterogeneous data structures and dynamically adjusting anti-cheating priorities by developing a dedicated integration subsystem, Mvin.
 We evaluate Hawk on a mainstream FPS game, CS:GO.
@@
 The scale of the dataset has increased by two orders of magnitude compared to the state-of-the-art work [6].
-Hawk achieves a maximum of 84% recall and 80% accuracy in detecting both aimbots and wallhacks, outperforming the official in-use anti-cheat.
-It successfully identifies cheaters who had evaded previous detections and those never caught by the official systems.
-We created a demo website [^1] Demo website.
-This is an anonymous website.
+Hawk achieves at most 84% recall and 80% accuracy in detecting aimbots and wallhacks.
+Hawk outperforms baselines in the separate comparisons across different subsystems, given Hawk’s diverse input data structures.
+It successfully identifies cheaters who had evaded previous detections and those who were never caught by official systems.
+We created a demo website [^1] Demo website: https://hawk-anticheat.github.io/.
 to showcase some escapees’ illegal actions.
-Hawk validates its robustness against cheat evolution (introduced in Section 6.7) in real-world scenarios and showcases acceptable overheads.
-We particularly discuss Hawk’s generalizability in Section 7.
-The contributions of our work are as follows.
+Hawk validates its robustness against cheat evolution (introduced in Section VI-G) in real-world scenarios and showcases acceptable overheads.
+We particularly discuss Hawk’s generalizability in Section VII.
+The contributions of our work are summarized below.
   • Innovative anti-cheat observations and feature designs.
   • Novel framework Hawk and a supporting workflow.
-  • Real-world evaluations with strong performance.
-  • Open-source Hawk and dataset [^2] Link.
-The repository and dataset will be released upon acceptance.
-Artifacts are uploaded for the reference.
-Ethics statement is shown in Section 10.
-for future research.
-2 Background
-This section introduces anti-cheat and cheating techniques, replay systems, and replay files in modern FPS games.
-2.1 Server-side and Client-side Anti-Cheat
-Client-side anti-cheats focus on data, process, and hardware protection [49, 58], or detection by either examining the existing cheats’ signatures [22, 3] or classifying through behavior [9].
-However, attackers can bypass it with higher privileges, by sniffing the detection report through side-channel attack [23] and altering them in the transport layer, or by changing the program’s signature [35].
-Server-side anti-cheats [69, 70, 2, 24, 39] emphasize data analysis and anomaly detection.
-Although the cheater’s inputs might be altered on the client, the cheaters must ensure their performance outperforms others to secure the win.
-Thus, it is easier to detect cheats on the server because it records all clients’ final-state actions that influence the gameplay.
-But the prior works have limitations mentioned in Section 1.
-2.2 Cheat Types in FPS Games
+  • Real-world data evaluations with strong performance.
+  • Open-source Hawk and dataset [^2] Public repository and dataset: 10.6084/m9.figshare.25940818. Ethics considerations are shown in Supplemental Material A. for future research.
+II Background
+The section introduces anti-cheat, cheating techniques, the current state of cheating, and the replay design in FPS games.
+II-A Server-Side and Client-Side Anti-Cheat
+Client-side anti-cheats focus on data, process, and hardware protection [9, 8], or detection by examining the existing cheats’ signatures [16, 17] or classifying through aiming behavior [6].
+Yet, attackers can bypass it with higher privileges, by sniffing the detection report through side-channel attack [37] and altering them in the transport layer, or by changing the program’s signature [38].
+Server-side anti-cheats [10, 11, 12, 39, 13, 31] emphasize data analysis and anomaly detection.
+Although the data might be altered on the client, they must secure victory by outperforming other players.
+Thus, detecting cheats on the server is easier because it records all clients’ final-state actions influencing the gameplay.
+But the prior works have the limitations mentioned in Section I.
+II-B Cheating in FPS Games
 Wallhack and aimbot are the two most prevalent types of cheats in FPS games [40].
@@
 Aimbot can be subdivided into pure aimbot (the primitive aimbot), triggerbot (auto-shoot only when the cross-hair is on the opponent), micro-settings (counteract recoil with Hotkeys), and computer-vision-based aimbot (aim with object detection models).
-In the following sections, aimbot refers to all aforementioned subcategories.
-Notably, cheating provides different levels of sophistication [66].
+In our paper, aimbot refers to all aforementioned subcategories.
+Cheating provides different levels of sophistication [21].
 The cheater might cheat cautiously, pretending to be a normal player, switching the cheat on and off, or cheating unabashedly with different configurations, various brands of cheating software, and categories.
 However, cheaters’ behaviors will always differ from the normal players [6].
-Thus as long as cheaters want to gain unfair advantages, they must outperform the average.
+Thus, cheaters must outperform the average to gain unfair advantages.
 This is bound to create anomalies at the data level and therefore should be identifiable by models.
-Our evaluations in Section 6 include different levels of cheating sophistication by using large real-world datasets.
-2.3 Replay System and Replay File
+Our evaluations in Section VI include different levels of cheating sophistication.
+II-C Replay System and Replay File
 Replay system is a dynamic toolset for player interaction with recorded content post-game and has been extensively adopted in notable FPS games.
@@
 We believe the replay file can be used as a novel data source for FPS server-side anti-cheat.
-3 Overview
+III Overview
 This section provides the key observations on the human experts’ identification process and the workflow of Hawk.
-3.1 Fact Observations
-Although the anti-cheat tasks are still challenging for automation systems in not just the FPS field but almost any game genre, the illegal activities are identifiable for seasoned human players [11].
+III-A Fact Observations
+Although the anti-cheat tasks are still challenging for automation systems in FPS games, the illegal activities are identifiable for seasoned human players [41].
 Therefore, by harnessing such inspiration, we discovered three observations that human experts always focus on to identify cheaters.
-{mdframed}[style=obs] Observation 1 (OBS.1).
+Observation 1 (OBS.1).
 Review a player’s point-of-view (POV).
 Akin to watching a player’s game replay unfold over time, this review process involves analyzing various in-game behaviors, such as aim and shot, positioning and movement, and the in-game economy management and props utilization.
-These behaviors can provide temporal insights into whether the player is legitimately playing the game.
-{mdframed}[style=obs] Observation 2 (OBS.2).
+These behaviors can provide temporal insights into the legitimacy of a player’s gameplay.
+Observation 2 (OBS.2).
 Conduct an in-depth analysis of a player’s statistical data.
@@
 These statistical deviations serve as strong indicators of cheating.
-{mdframed}[style=obs] Observation 3 (OBS.3).
+Observation 3 (OBS.3).
 Examine the consistency between a player’s game sense and combat performance.
@@
 For instance, if a player moves like a rookie but shoots and aims incredibly accurately, it is suspicious.
-3.2 Hawk’s Workflow
-Figure 1 illustrates Hawk’s workflow, comprising four interlinked stages.
-Stage One.
-Preprocessing.
-The demo is applied to parse in-game data from the banned databases.
-The multi-view features are then extracted from parsed data and then annotated for training.
-We discuss the feature constructions in Section 4.
-Stage Two.
-Hawk deployment.
-Hawk utilizes the extracted features as inputs for its subsystems RevPov (Section 5.1), RevStats (Section 5.2), and ExSPC (Section 5.3) to mimic human identification process, respectively.
-The outputs are incorporated into Mvin subsystem (Section 5.4) that presents a final cheating report.
-Stage Three.
-Game Master (GM) team verification.
-Given the potential personal assets tied to a player’s account, a ban should not be issued solely based on an algorithmic verdict.
-It is imperative to provide substantial evidence before initiating such an action.
-Hence, the final verdict requires GMs’ verification to avoid false bans.
-Stage Four.
-Data updates.
-Upon the affirmation of cheating, the relevant replay files are added to the banned databases.
-This act enriches and updates the reservoir of training samples which may contain novel patterns and behaviors.
-Thus, four stages constitute a cyclical workflow, with each stage contributing to the subsequent one.
-This ongoing cyclical workflow enhances the system’s robustness and provides a dynamic solution to multi-type cheat detection. We further discuss this in Section 6.7
+III-B Workflow
+Figure 1 illustrates Hawk’s interlinked workflow, which enhances the system’s robustness and provides a dynamic solution to multi-type cheat detection (discussed in Section VI-G).
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
 Figure 1: Cyclical workflow of Hawk.
-4 Feature Constructions
-The feature design is one of our contributions and key to the server-side anti-cheat.
-This section introduces the construction of temporal features, structured features, and the classification of sense and performance features.
-These features are intended to be a comprehensive representation of the three aspects of human identification of cheating in Section 3.1.
-4.1 Temporal Features
-Temporal features are information derived from the demo in the time domain.
-Each temporal data tuple represents the state information (e.g., coordinates, view directions, events, etc.) associated with the present tick.
+IV Hawk’s Feature Design
+The feature design is one of our contributions.
+This section introduces the construction of structured features, temporal features, and the classification of sense and performance features.
+These features are selected among hundreds of our designed naive features and intended to be comprehensive representations of the three aspects of human identification of cheating in Section III-A to overcome the challenge of the potentially biased classifications brought by the monospecific features in the prior works mentioned in Section I.
+IV-A Structured Features
+Structured features are single-value data calculated with well-designed algorithms to represent integrated in-game behavioral information per match and ensure time complexity.
+This section describes the components of structured features and visualizes the structured feature differences between normal players and cheaters.
+IV-A1 Structured Features Construction
+The structured features for Hawk are categorized into four main categories that comprise 28 features.
+For the sake of brevity and coherence, only the introduction of categories is outlined below.
+A meticulous breakdown is shown in Supplemental Material C.
+  • Aiming features represents the efficiency and efficacy of aims.
+We break down the aiming process into three moments (initial spot, first-time fire, and first-time hit) and two stages (reaction and adjustment).
+We extract different elapsed durations and angle variations with statistical representations (e.g., average or variance) to indicate the player’s different performances within a match.
+For example, the average of reaction duration reflects a player’s level of reaction speed, and the variance of that denotes the distribution of the reaction duration.
+Typically, in terms of the reaction stage, normal players obtain a higher mean value and more diverse variance compared with cheaters.
+  • Firing features reflects shot proficiency.
+For example, the hit group distribution variance describes the dispersion of body part that suffers damage for the current player in a match.
+Most cheaters with aimbot prefer to lock on specific hit group (i.e., head) to achieve faster elimination.
+  • Elimination features emphasizes metrics related to in-game kill events.
+For example, the occluder-penetration index is a weighted index that reflects the extent of the average occluder-penetration eliminations (e.g., eliminate the opponent behind walls or smoke) conducted by the player.
+In terms of the cheaters exploiting wallhacks, this index is more significant than the normal players’.
+  • Props utilization features captures a player’s adeptness at using in-game auxiliary and offensive props like flashbang, grenades, etc.
+For instance, the props utilization index denotes how frequently a player utilizes auxiliary and offensive props such as incendiaries and smoke grenades.
+This can help distinguish a player’s familiarity with the game and their skillfulness.
+Figure 2: Box plots of comparative visualization between honest players and cheaters on top-12 structured features under the Mann-Whitney U test with distinction descending order.
+IV-A2 Structured Data Comparative Visualization
+Figure 2 offers a box-plot on structured features to analyze between honest players and cheaters comparatively.
+Each feature on the x-axis (we select top-12 features for illustration purposes) is ordered based on the outcomes of the Mann-Whitney U test, with the leftmost features, such as akpr (introduced in Supplemental Material C), marking pronounced disparities between the two groups.
+This implies that features towards the left end, showcased by their lower p-values, are more distinguishable, whereas those on the right are less so.
+The y-axis is normalized values between 0 and 1.
+The box plots, distinguished by light blue for honest players and hatched red for cheaters, provide insights into the median, interquartile range, and outlier distributions.
+In which, diamond shapes represent outliers, emphasizing the spread and distribution of data across the features.
+Despite considering dimensionality reduction techniques, such as PCA (Supplemental Material Figure 4), empirical tests show an optimal performance when all 28 features were employed.
+Therefore, cheaters often exhibit deviations in their static data compared to normal players.
+Even when cheaters deliberately attempt to conceal their behavior, the subtle differences tend to accumulate over time and reveal disparities statistically.
+IV-B Temporal Features
+Temporal features are derived in the time domain.
+Each temporal data tuple is the state information (e.g., coordinates, view directions, events, etc.) associated with the present tick.
 Where tick denotes the smallest unit of time in the game.
-This section describes temporal features’ components and illustrates the distinctions in temporal features between normal players and cheaters.
-4.1.1 Temporal Features Construction
-According to different in-game behaviors, temporal features are segmented into three main categories and subdivided into seven specific types.
-This overview provides only categories and types due to the less-expressive nature of temporal features.
-Appendix B lists elaborate descriptions.
+IV-B1 Temporal Features Construction
+According to different in-game behaviors, temporal features are segmented into three main categories and subdivided into seven specific types. Supplemental Material D lists elaborate descriptions.
   • Engagement features describes in-game engagement events, which are subcategorized into five types as follows.
@@
     • Auxiliary props features related to auxiliary props-utilization events (e.g., flashbang), including the deploy tick, victim’s affected duration, etc.
-    • Offensive props features related with events involving props that aid the attack (e.g., grenade, incendiary, etc.) including the deploy and destroy ticks, the props’ type, the props’ coordinates, etc.
+    • Offensive props features related with events involving props that aid the attack (e.g., grenade, incendiary, etc.) including the deploy and destroy ticks, the props’ type and coordinates, etc.
     • Elimination features describes kill-related events, including the occurring tick, attacker’s weapon and location, victim’s location and view direction, etc.
@@
   • Economy features reflects in-game financial aspects. For example, they include the equipment value and balance per round for each player.
-4.1.2 Temporal Data Comparative Visualization
+IV-B2 Temporal Data Comparative Visualization
 Figure 3 comparatively analyzes the temporal features between cheaters and normal players across one round.
@@
 Figure 3: Comparative behavioral visualization between honest players and cheaters with respect to time scales.
-4.2 Structured Features
-Structured features are single-value data calculated with well-designed algorithms to represent integrated in-game behavioral information per match.
-This section describes the components of structured features and visualizes the structured feature differences between normal players and cheaters.
-4.2.1 Structured Features Construction
-The structured features for Hawk are categorized into four main categories that comprise 28 strong-expressiveness features.
-For the sake of brevity and coherence, only the introduction of categories is outlined below.
-A meticulous breakdown can be referred to in Appendix C.
-  • Aiming features represents the efficiency and efficacy of aims.
-We break down the aiming process into three moments (initial spot, first-time fire, and first-time hit) and two stages (reaction and adjustment).
-We extract different elapsed durations and angle variations with statistical representations (e.g., average or variance) to indicate the player’s different performances within a match.
-For example, the average of reaction duration reflects a player’s level of reaction speed, and the variance of that denotes the distribution of the reaction duration.
-Typically, in terms of the reaction stage, normal players obtain a higher mean value and more diverse variance compared with cheaters.
-  • Firing features reflects shot proficiency.
-For example, the hit group distribution variance describes the dispersion of body part that suffers damage for the current player in a match.
-Most cheaters with aimbot prefer to lock on specific hit group (i.e., head) to achieve faster elimination.
-  • Elimination features emphasizes metrics related to in-game eliminations or kill-focused activities.
-For example, the occluder-penetration index is a weighted index that reflects the extent of the average occluder-penetration eliminations (e.g., eliminate the opponent behind walls or smoke) conducted by the player.
-In terms of the cheaters exploiting wallhacks, this index is more significant than the normal players’.
-  • Props utilization features captures a player’s adeptness at using in-game auxiliary and offensive props like flashbang, grenades, etc.
-For instance, the props utilization index denotes how frequently a player utilizes auxiliary and offensive props such as incendiaries and smoke grenades.
-This can help distinguish a player’s familiarity with the game and their skillfulness.
-Figure 3: Box plots of comparative visualization between honest players and cheaters on top-12 structured features under Mann-Whitney U (Wilcoxon Rank-Sum) test with P-Value ascending (distinction descending) order.
-4.2.2 Structured Data Comparative Visualization
-Figure 3 offers a box-plot on structured features to analyze between honest players and cheaters comparatively.
-Each feature on the x-axis (we select top-12 features for illustration purposes) is ordered based on the outcomes of the Mann-Whitney U test, with the leftmost features, such as akpr (introduced in Appendix C), marking pronounced disparities between the two groups.
-This implies that features towards the left end, showcased by their lower p-values, are more distinguishable, whereas those on the right are less so.
-The y-axis is normalized values between 0 and 1.
-The box plots, differentiated by blue for honest players and orange for cheaters, provide insights into the median, interquartile range, and outlier distributions.
-In which, diamond shapes represent outliers, emphasizing the spread and distribution of data across the features.
-Despite considering dimensionality reduction techniques, such as PCA (Figure 13), empirical tests show an optimal performance when all 28 features were employed.
-4.3 Sense and Performance Features
-Sense and performance features are derived from the two subsections above. Appendix D includes detailed classification.
+IV-C Sense and Performance Features
+Sense and performance features are derived from the previous subsections. Supplemental Material E includes detailed classification.
   • Sense features denote a player’s in-game cognition, reflecting strategic understanding, and tactical anticipation.
@@
     • Temporal performance features include temporal features that indicate continuous in-game performances, e.g., eliminations, damage, weapon proficiency, etc.
-    • Structured performance features contain the remaining structured features that statistically describe the player’s in-game performances, such as time to kill, elimination-related features, etc.
+    • Structured performance features contain the remaining structured features that statistically describe the player’s in-game performances, e.g., elimination-related features.
 Figure 4: Hawk framework illustration.
-(a)-RevPov, (b)-RevStats, (c)-ExSPC are corresponding to three observations in Section 3.1, (d)-Mvin is an integration network for combining the determinations from the aforementioned three subsystems.
-5 Hawk
-To mimic human experts’ identification process, we propose Hawk which consists of four subsystems: Review Point-of-View (RevPov) subsystem, Review Statistics (RevStats) subsystem, Examine Sense-Performance Consistency (ExSPC) subsystem, and Multi-View Integration (Mvin) subsystem.
-The first three subsystems’ designs are associated with the three observations we discovered regarding the inherent nature of the cheat identification process as elaborated in Section 3.1.
-The fourth subsystem integrates the results derived from the first three subsystems.
-5.1 RevPov
+(a)-RevPov, (b)-RevStats, (c)-ExSPC are corresponding to three observations in Section III-A, (d)-Mvin is an integration network for combining the determinations from the aforementioned three subsystems.
+V Hawk
+To mimic human experts’ identification process, we propose Hawk which consists of four subsystems: Review Point-of-View (RevPov), Review Statistics (RevStats), Examine Sense-Performance Consistency (ExSPC), and Multi-View Integration (Mvin).
+The first three subsystems are associated with the three observations we discovered regarding the inherent nature of the cheat identification process in Section III-A.
+Mvin integrates the results derived from the above three subsystems, just as humans integrate the three aspects collectively.
+V-A RevPov
 Aligned with OBS.1, RevPov thoroughly examines a player’s time-series operations within a match.
-Multi-LSTM Attention Encoders: In Figure 4 (a), we feed the seven sets of temporal features (Section 4.1.1) as input to seven independent Long Short-Term Memory (LSTM) [29] networks to acquire the embeddings.
+Multi-LSTM Attention Encoders: In Figure 4a, we feed the seven sets of temporal features (Section IV-B1) as input to seven independent Long Short-Term Memory (LSTM) [42] networks to acquire the embeddings.
 For explanatory purposes, we illustrate the process using the example of the Economy features.
 $\mathbf{F_{eco}^{n}}$ denotes the input feature vector.
-A dropout layer [54] is employed from the second to last LSTM layers respectively to mitigate the risk of overfitting.
-$\vec{h}_{G}^{n}$ denotes the output of the encoder, consists of one set of the seven temporal features mentioned in Section 4.1.1, where $G$ represents the set and $n$ indicates the sequence number among all outputs.
-For each time step, we introduce the Luong-style attention [41] afterward to help the model better understand the importance of different positions in the input sequence and aggregate different parts of the input weights.
+Dropout layers [43] are employed to mitigate the risk of overfitting.
+$\vec{h}_{G}^{n}$ denotes the output of the encoder, consists of one set of the seven temporal features mentioned in Section IV-B1, where $G$ represents the set and $n$ indicates the sequence number among all outputs.
+For each time step, we introduce the Luong-style attention [44] afterward to help the model better understand the importance of different positions in the input sequence and aggregate different parts of the input weights, overcoming the challenge of identifying varying patterns in temporal data for a large number of samples.
 In Equation 1, $a_{G}^{i}$ represents the attention weight for the time step $i$.
@@
 $i$, $j$ and $k$ indicate different time steps.
-Concatenation operation is executed afterward, yielding $[\vec{a}_{G}^{n}||\vec{h}_{G}^{n}]$, where ||||| | denotes concatenation.
+Concatenation operation is executed afterward, yielding $\vec{a}_{G}^{n}||\vec{h}_{G}^{n}]$, where $||$ denotes concatenation.
 After two dense layers, $\vec{w}_{G}^{n}$ represents the final tensor for binary classification.
 $\vec{a}_{G}^{i}=\sum_{j=0}^{N}\frac{\exp{(\mathbf{W}_{q}\cdot\vec{h}_{G}^{i})\cdot(\mathbf{W}_{k}\cdot\vec{h}_{G}^{j})}}{\sum_{k=0}^{N}\exp{(\mathbf{W}_{q}\cdot\vec{h}_{G}^{i})\cdot(\mathbf{W}_{k}\cdot\vec{h}_{G}^{k})}}\cdot\vec{h}_{G}^{j}$ (1)
-Embedding Generation: An average pooling is conducted to flatten the output sequence.
-Equation 2 indicates the average operation, where $N$ denotes the number of output sequences.
-$\vec{V}_{pov}$ is used as embedding for the final determination.
+Embedding Generation: An average pooling in Equation 2 is conducted to flatten the output sequence, where $N$ denotes the number of output sequences.
+$\vec{V}_{pov}$ is the final determination embedding.
 $\vec{w}_{G}=\frac{\sum_{n=0}^{N}\vec{w}_{G}^{n}}{N+1}$ (2)
-$\vec{V}_{pov}=[\vec{w}_{eco}||\vec{w}_{mov}||\vec{w}_{dmg}||\vec{w}_{elm}||\vec{w}_{aux}||\vec{w}_{off}||\vec{w}_{wf}]$ (3)
-Final Determination: In the context of cheat detection, the dataset often exhibits a class imbalance, where the majority of players are honest, inadvertently leading to biased learning towards the dominant class.
-Therefore, we adopt a multi-subsampling strategy to level the playing field for the minority class, creating balanced representations of both cheating and honest behaviors.
-Suppose $\mathcal{T}$ is the initial training set.
-To establish a balanced class distribution, we perform a selective sampling on $\mathcal{T}$, specifically targeting the honest player samples.
-We perform this operation nine times in accordance with the ratio of honest to dishonest players, generating nine distinct training sets, each denoted as $\mathcal{T}_{k}$, where $k$ indicates the iteration count.
-Conduct each iteration of the sampling process to select unique honest player samples, thereby adding varied examples to the learning process.
-Subsequently, we obtain nine separately trained Random Forest models, each corresponding to one of the diversified and balanced training sets $\mathcal{T}_{k}$, we named this strategy as multi-forest.
-The averaging method in Equation 4 is selected for RevPov final class prediction.
-Consider a random forest consisting of $N$ decision trees, each producing a prediction denoted as $h_{i}(\vec{V}_{pov}^{k})$, where $i$ indicates the $i$-th tree, $k$ denotes the $k$-th sampled set, and $\vec{V}_{lstm}^{k}$ represents the input feature vector.
-The prediction of each tree is a probability distribution, indicating the likelihood of the sample $\vec{V}_{pov}^{k}$ belonging to each class.
+[…]
+Final Determination: We find that only depending on the fully connected layers with different activation functions can not effectively classify.
+Therefore, we conduct the following strategies to further improve RevPov.
+In the context of cheat detection, the dataset often exhibits a class imbalance, where the majority of players are honest, inadvertently leading to biased learning towards the dominant class.
+Therefore, we create a multi-subsampling strategy to level the playing field for the minority class, creating balanced representations of both cheating and honest behaviors without dropping any samples.
+We perform a selective sampling to divide the honest players in the initial training set $\mathcal{T}$ into nine sets in accordance with the ratio of normal to cheating players, generating nine training sets containing the same set of cheaters and distinct normal players, each denoted as $\mathcal{T}_{k}$, where $k$ indicates the iteration count.
+Subsequently, after adding varied examples to the learning process, we obtain nine independent Random Forest models, each corresponding to one of the diversified and balanced training sets $\mathcal{T}_{k}$, we named this strategy as multi-forest.
+The averaging method in Equation 4 is selected for RevPov final class prediction, where random forest consisting of $N$ decision trees, each producing a prediction $h_{i}(\vec{V}_{pov}^{k})$, where $i$ indicates the $i$-th tree, $k$ denotes the $k$-th sampled set, and $\vec{V}_{pov}^{k}$ represents the input feature vector.
 $p_{i}$ represents the probability that the $i$-th tree predicts sample $\vec{V}_{pov}^{k}$ to be in class $j$.
-In Equation 4, $\mathbf{\overline{RF}}(\vec{V}_{pov}^{k})$ denotes the average prediction output of the random forest, which is the average of the prediction probability vectors of all decision trees.
-The prediction for each $\vec{D}_{k}$ can be obtained by selecting the class with the highest predicted probability shown in Equation 5, where $\left(\mathbf{\overline{RF}}(\vec{V}_{pov}^{k})\right)_{j}$ represents the probability of class $j$ in $\mathbf{\overline{RF}}(\vec{V}_{pov}^{k})$.
+$\mathbf{\overline{RF}}(\vec{V}_{pov}^{k})$ denotes the average prediction output of the random forest, which is the average of the prediction probability vectors of all decision trees.
+The prediction for each $\vec{D}_{k}$ is obtained by selecting the class with the highest predicted probability shown in Equation 5, where $(\cdot)_{j}$ represents the probability of class $j$ in $\mathbf{\overline{RF}}(\vec{V}_{pov}^{k})$.
 After the multi-forest, we acquire the corresponding determination $\mathbf{D}_{k}$.
@@
 $\mathbf{D}_{pov}=\arg\max_{j}\sum_{k=1}^{9}\mathbf{D}_{k}^{(j)}$ (6)
-5.2 RevStats
+V-B RevStats
 Aligned with OBS.2, RevStats subsystem is designed to perform an in-depth analysis of structured features to capture the cheaters.
-Cheaters often exhibit deviations in their static data compared to normal players.
-Even when cheaters deliberately attempt to conceal their behavior, the subtle differences tend to accumulate over time and reveal disparities in the statistical data post-game.
-Feature Extraction: In Figure 4 (b), $\vec{V}_{28}$ denotes the vector representation of the structured features (Section 4.2.1).
-Ensembled Committee: We empirically use seven classifiers [28, 14, 7, 13, 46, 34, 55] in an ensembled way.
-Each model is independently trained.
-Upon completion, each model independently adjudicates each test sample.
-The outputs are binary signifying the presence or absence of cheating.
+Feature Extraction: In Figure 4b, $\vec{V}_{28}$ denotes the vector representation of the structured features (Section IV-A1).
+Ensembled Committee: During preliminary testing, our selected features showed promising results across most classifiers, demonstrating their effectiveness.
+However, despite similar performance in terms of metrics alone, the testing revealed that different classifiers produced varying classifications on some individual samples.
+This suggests that relying on a single classifier, as done in prior works, can lead to unreliable outcomes.
+However, the majority of classifiers correctly classified the samples, leading to the development of the Ensembled Committee.
+Among hundreds of classifiers, we identify seven superior classifiers [45, 46, 47, 48, 49, 50, 51] after the preliminary testing and use them in an ensembled way.
 Final Determination: Due to imbalanced dataset, a similar multi-subsampling strategy as employed in the RevPov subsystem is adopted for the RevStats subsystem to fully leverage the available data without losing valuable information.
 For each classification model, we generate a distinct model for each subsampled set.
-Each model provides an individual prediction.
-A majority voting scheme is utilized within each classification algorithm to aggregate the results of the nine models.
-The final determination is also achieved through the majority voting mechanism, which collates the determinations from all seven classification algorithms.
+Each model is independently trained and provides an individual prediction.
+A majority voting scheme is utilized within each classification algorithm to aggregate the results of the seven models.
+The final determination is also achieved through majority voting, which collates the determinations from all nine subsampled sets.
 In Equation 7, where $\mathbf{D}_{stats}$ signifies the binary final determination.
@@
 $\mathbf{D}_{stats}=\arg\max_{i\in\{0,1\}}\sum_{j=1}^{7}I(w_{j}=i)$ (7)
-5.3 ExSPC
-Aligned with OBS.3, ExSPC subsystem is designed to inspect a comprehensive investigation of the consistency of a player’s in-game senses and overall performances.
-Embedding and Feature Grouping: There are seven embeddings and two groups of vectors as input (Section 4.3) in ExSPC, in which seven embeddings are generated by using the trained models in RevPov and two vectors are the subsets of $\vec{V}_{28}$ in RevStats.
-Each embedding is padded to the maximum length of that in the whole dataset.
-$\vec{w}^{k}_{G}$ is the intermediary result from RevPov, where $k$ represents the numbering in the sequence, ${G}$ denotes the type of the temporal features.
-Each embeddings are represented in two-dimensional arrays.
-In Figure 4 (c), $P_{1}$ to $P_{k}$ indicates different players, we exemplify with one player $P_{j}$.
-Therefore, we flatten the embeddings into one-dimensional arrays and use a dense layer with an ELU activation function to shrink the dimensions.
-After the seven different similar groups of layers, we obtain $\varphi_{G}^{{P}_{j}}$.
-Subsequently, we divide seven different types of tensors as shown in Appendix D, and concatenate each group of tensors for each player.
-In which, sense vector of the temporal features is $\vec{V}_{POV}^{S}=[\varphi_{mov}^{{P}_{j}}||\varphi_{eco}^{{P}_{j}}||\varphi_{off}^{{P}_{j}}||\varphi_{aux}^{{P}_{j}}]$ and performance vector of the temporal features is $\vec{V}_{POV}^{P}=[\varphi_{wf}^{{P}_{j}}||\varphi_{elm}^{{P}_{j}}||\varphi_{dmg}^{{P}_{j}}]$.
-In terms of two vectors, we also divided $\vec{V}_{28}$ into two groups according to Appendix D, namely the sense vector of structured features $\vec{V}_{28}^{S}$ and the performance vector of structured features $\vec{V}_{28}^{P}$.
-Dimension Reduction: $\vec{V}_{28}^{S}$ and $\vec{V}_{28}^{P}$ are normalized at first.
+V-C ExSPC
+Aligned with OBS.3, ExSPC is designed from scratch to inspect the consistency of a player’s in-game senses and overall performances.
+Embedding and Feature Grouping: ExSPC takes seven embeddings and two groups of vectors as input (see Section IV-C).
+The seven embeddings are generated using trained models from RevPov, while the two vectors are subsets of $\vec{V}_{28}$ from RevStats.
+Each embedding is padded to the maximum length observed in the dataset.
+The intermediary result from RevPov, denoted as $\vec{w}^{k}_{G}$, represents temporal features, where $k$ is the sequence number and G{G} denotes the type of the temporal features.
+All embeddings are structured as two-dimensional arrays.
+As illustrated in Figure 4c, where $P_{1}$ to $P_{k}$ represent different players, we exemplify with one player $P_{j}$.
+The embeddings are first flattened and dimensionally reduced.
+After processing through seven groups of similar layers, we obtain $\varphi_{G}^{{P}_{j}}$.
+These tensors are then divided into seven types and concatenated for each player.
+In which, $\vec{V}_{POV}^{S}$ denotes the temporal sense features vector and $\vec{V}_{POV}^{P}$ denotes the temporal performance features vector.
+Similarly, the structured features vector $\vec{V}_{28}$ is divided into two groups: the sense vector $\vec{V}_{28}^{S}$ and the performance vector $\vec{V}_{28}^{P}$, as detailed in Supplemental Material E.
+Dimension Reduction: $\vec{V}_{28}^{S}$ and $\vec{V}_{28}^{P}$ are normalized.
 Afterward, each group of the four vectors goes through different groups of dense layers to reduce the dimensions.
-Next, we concatenate four vectors and yield two vectors eventually representing sense $\vec{V}_{SENSE}=[\vec{\xi}_{SENSE}^{STATS}||\vec{\xi}_{SENSE}^{POV}]$ and performance $\vec{V}_{PREF}=[\vec{\xi}_{PREF}^{STATS}||\vec{\xi}_{PREF}^{POV}]$ respectively.
-Final Determination: When acquiring $\vec{V}_{SENSE}$ and $\vec{V}_{PERF}$, we further reduce its dimension and deepen the network for better learning performances.
+Next, concatenate four vectors and yield two vectors eventually representing sense $\vec{V}_{SENSE}$ and performance $\vec{V}_{PREF}$ respectively.
+Final Determination: When acquiring $\vec{V}_{SENSE}$ and $\vec{V}_{PERF}$, we further reduce their dimensions and deepen the network for better learning performances.
 Then, we concatenate two vectors and go through one final dense layer for reducing the dimension and another dense layer with sigmoid activation function for binary classification.
@@
 $\text{Loss}=-\frac{1}{N}\sum_{i=1}^{N}\left[y_{i}\log(\hat{y}_{i})+(1-y_{i})\log(1-\hat{y}_{i})\right]$ (8)
-5.4 Mvin
-In Figure 4 (d), Mvin is to harness the collective determinations from RevPov, RevStats, and ExSPC.
-By integrating the predictions from these subsystems, Mvin aims to deliver a final verdict regarding a player’s cheating disposition.
+V-D Mvin
+In Figure 4d, Mvin is to dynamically integrate determinations from RevPov, RevStats, and ExSPC.
 Data Integration: The subsystem integrates the outcomes from RevPov, RevStats, and ExSPC, symbolized as $\mathbf{D}_{pov}$, $\mathbf{D}_{stats}$, and $\mathbf{D}_{spc}$.
@@
 This step is designed to merge the individual outputs and synchronize them for the succeeding integration process.
-[…]
-[…]
-[…]
-6 Experiments
-We conduct experiments to answer six research questions (RQ.1--RQ.6):
-  • RQ.1: What are the performances of Hawk and each subsystem? (Section 6.4)
-  • RQ.2: Compared to the current official inspection, to what extent Hawk has improved in terms of effectiveness and efficiency? (Section 6.5)
-  • RQ.3: What are the performances of Hawk under different settings of the Task-Specified Threshold Optimizer? (Section 6.6)
-  • RQ.4: What is the robustness of Hawk and its subsystems against cheat evolution? (Section 6.7)
-  • RQ.5: Is there any suspicious activity that still exists but has never been detected by officials? (Section 6.8)
-  • RQ.6: What are the overheads of Hawk’s subsystems? (Section 6.9)
-6.1 Incomparability to Prior Works
-First, most prior works [69, 24, 70, 2, 39, 71] are not open-sourced.
-Second, data collection method is distinct and incompatible.
-Hawk utilizes the replay file post-game, which differs fundamentally from prior studies that modified the game engine or applied mods during gameplay.
-Third, the installation on commercial clients or servers is infeasible.
-If the installation of the prior works is implemented before the commencement of data collection, the data can then be gathered and evaluated.
-However, the demos we collected come directly from the game platform, which prohibits the installation of any unverified programs on clients or servers due to legal constraints.
-In addition to the aforementioned reasons with commonality, we also further state below the specific reasons for the detailed incomparability between Hawk and the SOTA designs.
-There are three aspects of incomparability with BotScreen [9].
-First, CS:GO has gone through several major patch updates after BotScreen was released.
-The authors of BotScreen confirmed through email that BotScreen, as well as Osiris [37] (the identical cheat used in BotScreen for feature extraction and evaluation), does not apply to the in-use version of games.
-Second, the deprecation of Intel SGX [5, 31] prevents us from conducting comparative experiments.
-Third, it can only detect pure aimbot, making the comparison less significant.
-BlackMirror [49] and Invisibility Cloak [58] are prevention designs only for counteracting wallhacks and computer-vision-based aimbots respectively.
-Both studies are only functional before the cheats happen.
-Whereas Hawk and other prior works are detection systems that functional post-cheating.
-Table 1: Dataset separation and description.
+Model Optimization: Mvin adopts a weight matrix for each subsystem’s determination, where weights are dynamically assigned based on the subsystem’s importance and reliability.
+The weighted output, $\mathbb{W}$, is calculated as Equation 9.
+Where $\lambda_{1}$, $\lambda_{2}$, and $\lambda_{3}$ represent the weights associated with the outputs of the subsystems RevPov, RevStats, and ExSPC, respectively.
+$\mathbb{W}=\lambda_{1}\cdot\mathbf{D}_{pov}+\lambda_{2}\cdot\mathbf{D}_{stats}+\lambda_{3}\cdot\mathbf{D}_{spc}$ (9)
+Final Determination: $\mathbb{W}$ undergoes a learnable thresholding via the Task-Specified Threshold Optimizer (TSTO), which provides a binary determination based on a dynamic threshold $\varepsilon$.
+The threshold optimizer can be dynamically set based on anti-cheat requirements.
+For instance, in Recall mode, the optimizer would aim to identify every potential cheater, accepting the risk of higher false positives.
+This threshold $\varepsilon$ therefore acts as a decision boundary, converting the continuous weighted value $\mathbb{W}$ into a discrete binary outcome $\mathbf{D}_{HAWK}$.
+TSTO offers Hawk a degree of selectivity, enabling flexible tuning and offering promising performance without requiring retraining any prior subsystems, overcoming the challenge of lack of system adaptability.
+By harnessing the strengths of each subsystem and compensating for their limitations, Mvin is crucial to integrate determinations from different subsystems.
+VI Experiments
+We conduct experiments to answer seven research questions (RQ.1-RQ.7):
+  • RQ.1: What are the performances of Hawk and each subsystem? (Section VI-C)
+  • RQ.2: To what extent has Hawk improved effectiveness and efficiency compared to current official inspections? (Section VI-D)
+  • RQ.3: What are the performances of Hawk under different settings of the Task-Specified Threshold Optimizer? (Section VI-F)
+  • RQ.4: What is the robustness of Hawk and its subsystems against cheat evolution? (Section VI-G)
+  • RQ.5: Is there any suspicious activity that still exists but has never been detected by officials? (Section VI-H)
+  • RQ.6: What is the overhead of Hawk? (Section VI-I)
+  • RQ.7: Is Hawk performing better than prior works? (Section VI-E)
+Figure 5: Experiments’ process and labeling details.
+TABLE II: Dataset separation and description.
 CT1 Train Validation Test
-#M2 #P3 #N4 #C5 #M2 #P3 #N4 #C5 #M2 #P3 #N4 #C5
+#M2 #P3 #N4 #C5 #M #P #N #C #M #P #N #C
 Aimbot 1,009 10,127 9,150 977 975 8,561 7,697 864 995 9,047 8,163 884
@@
 1CT: Cheat type; 2#M: Number of matches; 3#P: Number of players in total; 4#N: Number of normal (honest) players; 5#C: Number of cheaters.
-Table 2: Ablation study results on each of the subsystems.
-CT1 System Validation Test
-TP TN FP FN Accuracy Recall NPV AUC-ROC OEI2 TP TN FP FN Accuracy Recall NPV AUC-ROC OEI2
-Aimbot RevPov 458 4,796 2,901 406 0.614 0.530 0.922 0. 577 1.246 481 5,088 3,075 403 0.616 0.544 0.927 0.584 1.283
-RevStats 453 7,290 407 411 0.904 0.524 0.947 0.736 4.941 566 6,797 1,366 318 0.814 0.640 0.955 0.736 2.864
-ExSPC 553 6,187 1,510 311 0.787 0.640 0.952 0.779 2.529 558 6,616 1,547 326 0.793 0.631 0.953 0.786 2.586
-Hawk 614 5,803 1,894 250 0.750 0.711 0.959 0.732 2.326 642 5,837 2,326 242 0.716 0.726 0.960 0.721 2.126
-Wallhack RevPov 573 6,081 2,664 394 0.685 0.593 0.939 0.644 1.670 627 6,364 2,885 398 0.680 0,612 0.941 0.650 1.684
-RevStats 604 8,241 504 363 0.911 0.625 0.958 0.783 4.143 855 6,988 2,261 170 0.763 0.834 0.976 0.795 1.946
-ExSPC 760 7,258 1,487 207 0.826 0.786 0.972 0.885 3.303 501 8,484 765 524 0.875 0.489 0.942 0.840 3.736
-Hawk 798 7,033 1,712 169 0.806 0.825 0.977 0.815 3.118 869 6,902 2,347 156 0.756 0.848 0.978 0.797 2.649
-1CT: Cheat type; 2OEI: Operational Efficiency Index.
-Table 3: Comparison result on the different model selection of RevStats.
-CT1 Model Validation Test
-TP TN FP FN Accuracy Recall NPV AUC-ROC OEI2 TP TN FP FN Accuracy Recall NPV AUC-ROC OEI2
-Aimbot LogisticRegression 570 6,204 1,493 294 0.791 0.660 0.955 0.733 2.614 713 4,897 3,266 171 0.620 0.807 0.966 0.703 1.772
-Naive Bayes 467 6,740 957 397 0.842 0.874 0.944 0.708 3.069 625 5,928 2,235 259 0.724 0.707 0.958 0.717 2.143
-RandomForest 415 7,694 3 449 0.947 0.480 0.945 0.740 4.313 427 7,634 529 457 0.891 0.483 0.944 0.709 4.313
-SVM 309 7,615 82 555 0.926 0.358 0.932 0.673 7.299 405 7,665 498 479 0.892 0.458 0.941 0.699 4.320
-MLP 858 7,697 0 6 0.999 0.993 0.999 0.997 9.901 508 7,325 838 376 0.866 0.575 0.951 0.736 3.674
-GaussianNB 467 6,740 957 397 0.842 0.541 0.944 0.708 3.069 625 5,928 2,235 259 0.724 0.707 0.958 0.717 2.143
-QDA 350 7,181 516 514 0.880 0.405 0.933 0.669 3.737 534 6,770 1,393 350 0.807 0.604 0.951 0.717 2.697
-RevStats 453 7,290 407 411 0.904 0.524 0.947 0.736 4.941 566 6,797 1,366 318 0.814 0.640 0.955 0.736 2.864
-Wallhack LogisticRegression 714 7,629 1,116 253 0.859 0.738 0.968 0.805 2.861 943 5,622 3,627 82 0.639 0.920 0.986 0.764 1.420
-Naive Bayes 604 7,677 1,068 363 0.853 0.625 0.955 0.751 2.741 891 6,103 3,146 134 0.681 0.869 0.979 0.765 1.545
-RandomForest 967 8,745 0 0 1.000 1.000 1.000 1.000 6.713 751 8,083 1,166 274 0.860 0.733 0.967 0.803 2.872
-SVM 548 8,323 422 419 0.913 0.567 0.952 0.759 4.378 817 7,227 2,022 208 0.783 0.797 0.972 0.789 2.065
-MLP 472 8,489 256 495 0.923 0.488 0.945 0.729 5.186 788 7,726 1,523 237 0.829 0.769 0.970 0.802 2.471
-GaussianNB 604 7,677 1,068 363 0.853 0.625 0.955 0.751 2.741 891 6,103 3,146 134 0.681 0.869 0.979 0.765 1.545
-QDA 493 8,276 469 474 0.903 0.510 0.946 0.728 4.057 815 7,270 1,979 210 0.787 0.795 0.972 0.791 2.095
-RevStats 604 8,241 504 363 0.911 0.625 0.958 0.783 4.143 855 6,988 2,261 170 0.763 0.834 0.976 0.795 1.946
-1CT: Cheat type; 2OEI: Operational Efficiency Index.
-6.2 Experimental Settings and Ground Truth
-The demos are open-sourced and acquired from a well-known and recognized CS:GO platform [^3] Link.
-We anonymize the platform’s URL and name.
-The initial data extraction utilizes an open-source library awpy [67] to parse the raw data in demos into JSON files, which contain comprehensive information for the following feature constructions.
-Table 1 is the dataset description.
-Figure 5 illustrates the experiments’ process.
+TABLE III: Incomparability to prior works.
+Work Task Data Extraction Targeting Hardware Constraint Code Availability
+Detection Prevention Runtime Replay file Aimbot Wallhack
+BotScreen [6] ✓ (deprecated) Intel SGX (deprecated)
+BlackMirror [9] Intel SGX (deprecated)
+Invisibility Cloak [8] None
+Prior works [10, 11, 12, 13] ✓ (unavailable) None
+Hawk None
+TABLE IV: Ablation study results on each of the subsystems.
+CT System Validation Test
+TP TN FP FN Accuracy Recall NPV AUC-ROC FPR TP TN FP FN Accuracy Recall NPV AUC-ROC FPR
+Aimbot RevPov 458 4,796 2,901 406 0.614 0.530 0.922 0.577 0.378 481 5,088 3,075 403 0.616 0.544 0.927 0.584 0.377
+RevStats 453 7,290 407 411 0.904 0.524 0.947 0.736 0.053 566 6,797 1,366 318 0.814 0.640 0.955 0.736 0.167
+ExSPC 553 6,187 1,510 311 0.787 0.640 0.952 0.779 0.196 558 6,616 1,547 326 0.793 0.631 0.953 0.786 0.190
+RevPov+RevStats 648 4,603 3,094 216 0.613 0.750 0.955 0.760 0.402 691 4,364 3,799 193 0.559 0.782 0.958 0.750 0.465
+RevPov+ExSPC 611 5,733 1,964 253 0.741 0.707 0.958 0.781 0.255 608 6,082 2,081 276 0.739 0.688 0.957 0.784 0.255
+RevStats+ExSPC 605 5,828 1,869 259 0.751 0.700 0.957 0.798 0.243 639 5,877 2,286 245 0.720 0.723 0.960 0.796 0.280
+Hawk 614 5,803 1,894 250 0.750 0.711 0.959 0.732 0.246 642 5,837 2,326 242 0.716 0.726 0.960 0.721 0.285
+Wallhack RevPov 573 6,081 2,664 394 0.685 0.593 0.939 0.644 0.305 627 6,364 2,885 398 0.680 0.612 0.941 0.650 0.312
+RevStats 604 8,241 504 363 0.911 0.625 0.958 0.783 0.058 855 6,988 2,261 170 0.763 0.834 0.976 0.795 0.245
+ExSPC 760 7,258 1,487 207 0.826 0.786 0.972 0.885 0.170 501 8,484 765 524 0.875 0.489 0.942 0.840 0.083
+RevPov+RevStats 604 8,241 504 363 0.911 0.625 0.958 0.798 0.058 855 6,988 2,261 170 0.763 0.834 0.976 0.816 0.244
+RevPov+ExSPC 806 6,900 1,845 161 0.793 0.834 0.977 0.885 0.211 573 8,257 992 452 0.859 0.559 0.948 0.840 0.107
+RevStats+ExSPC 791 7,110 1,635 176 0.814 0.818 0.976 0.889 0.187 868 6,907 2,342 157 0.757 0.847 0.978 0.859 0.253
+Hawk 798 7,033 1,712 169 0.806 0.825 0.977 0.815 0.196 869 6,902 2,347 156 0.756 0.848 0.978 0.797 0.254
+TABLE V: Selected comparison result between different superior baseline models and RevStats.
+CT Model Validation Test
+TP TN FP FN Accuracy Recall NPV AUC-ROC FPR TP TN FP FN Accuracy Recall NPV AUC-ROC FPR
+Aimbot LogisticRegression [46] 570 6,204 1,493 294 0.791 0.660 0.955 0.733 0.194 713 4,897 3,266 171 0.620 0.807 0.966 0.703 0.400
+Naive Bayes [49] 467 6,740 957 397 0.842 0.874 0.944 0.708 0.124 625 5,928 2,235 259 0.724 0.707 0.958 0.717 0.274
+RandomForest [47] 415 7,694 3 449 0.947 0.480 0.945 0.740 0.000 427 7,634 529 457 0.891 0.483 0.944 0.709 0.065
+SVM [48] 309 7,615 82 555 0.926 0.358 0.932 0.673 0.011 405 7,665 498 479 0.892 0.458 0.941 0.699 0.061
+MLP [45] 858 7,697 0 6 0.999 0.993 0.999 0.997 0.000 508 7,325 838 376 0.866 0.575 0.951 0.736 0.103
+GaussianNB [50] 467 6,740 957 397 0.842 0.541 0.944 0.708 0.124 625 5,928 2,235 259 0.724 0.707 0.958 0.717 0.274
+QDA [51] 350 7,181 516 514 0.880 0.405 0.933 0.669 0.067 534 6,770 1,393 350 0.807 0.604 0.951 0.717 0.171
+RevStats 453 7,290 407 411 0.904 0.524 0.947 0.736 0.053 566 6,797 1,366 318 0.814 0.640 0.955 0.736 0.167
+Wallhack LogisticRegression 714 7,629 1,116 253 0.859 0.738 0.968 0.805 0.128 943 5,622 3,627 82 0.639 0.920 0.986 0.764 0.392
+Naive Bayes 604 7,677 1,068 363 0.853 0.625 0.955 0.751 0.122 891 6,103 3,146 134 0.681 0.869 0.979 0.765 0.340
+RandomForest 967 8,745 0 0 1.000 1.000 1.000 1.000 0.000 751 8,083 1,166 274 0.860 0.733 0.967 0.803 0.126
+SVM 548 8,323 422 419 0.913 0.567 0.952 0.759 0.048 817 7,227 2,022 208 0.783 0.797 0.972 0.789 0.219
+MLP 472 8,489 256 495 0.923 0.488 0.945 0.729 0.029 788 7,726 1,523 237 0.829 0.769 0.970 0.802 0.165
+GaussianNB 604 7,677 1,068 363 0.853 0.625 0.955 0.751 0.122 891 6,103 3,146 134 0.681 0.869 0.979 0.765 0.340
+QDA 493 8,276 469 474 0.903 0.510 0.946 0.728 0.054 815 7,270 1,979 210 0.787 0.795 0.972 0.791 0.214
+RevStats 604 8,241 504 363 0.911 0.625 0.958 0.783 0.058 855 6,988 2,261 170 0.763 0.834 0.976 0.795 0.245
+VI-A Incomparability to Prior Works
+First, most prior works are not open-sourced as listed in Table I.
+Second, the data collection method is incompatible.
+Hawk relies on post-game replay files, whereas prior studies modified the game engine or applied mods during gameplay.
+Prior works are unable to extract data from the replay files.
+Third, installation on commercial clients or servers is infeasible.
+Only if the prior designs are implemented before the data collection, the data can then be gathered and evaluated.
+However, the partner platform prohibits installing unverified programs on clients or servers due to legal restrictions.
+In addition to the aforementioned reasons with commonality, we also further state below the specific reasons for the detailed incomparability between Hawk and the SOTA designs.
+There are three aspects of incomparability with BotScreen [6].
+First, BotScreen’s data extraction is highly coupled with the game version and deprecated.
+The authors of BotScreen confirmed through email that their data extraction does not apply to the current game versions.
+Second, the deprecation of Intel SGX [14, 15] prevents us from conducting comparative experiments.
+Third, it can only detect pure aimbot, making the comparison less significant.
+Invisibility Cloak [8] and BlackMirror [9] are prevention designs only for counteracting computer-vision-based aimbots and wallhacks, respectively.
+Both studies are only functional before the cheats happen.
+Figure 6: Recall and FPR performance comparison among Hawk-Recall priority mode, Hawk-Balance mode, and our partner in-use anti-cheat [^4] Commercial anti-cheat (detailed in Section VIII), not open-source, leverages reverse engineering on the known cheats to detect cheating signatures..
+Hawk’s different modes are introduced and evaluated in Section V-D and Section VI-F.
+Whereas Hawk and other prior works are detection systems that function after cheating.
+We detailed the incomparability to prior works in Table III.
+We manage to compare Hawk with our partner’s industrial in-use anti-cheat in footnote 4 and reproduce two server-side works [11, 12] for further comparison in Table IX.
+VI-B Experimental Settings and Ground Truth
+The demos are open-sourced and acquired from a well-known CS:GO platform5555EPlay: https://www.5eplay.com/..
+The initial data extraction utilizes a public library awpy [52] to parse the demo into a JSON file, which contains the raw data for Hawk’s feature constructions.
+Table II is the dataset description.
+Figure 5 illustrates our experimental process.
 We evaluate Hawk with two of the most prevalent cheat types, aimbot and wallhack.
-For aimbot, we set one collection window (16 days in total).
-For wallhack, we set two collection windows (27 days in total, we explain the reason for window inconsistency in Section 6.7).
-All demos are real-world ranking matches within the collection windows.
-Therefore, the cheats used by the player are black-boxes for anti-cheating.
-Both datasets contain different levels of cheating sophistication.
-The dates of cheaters’ label acquisition both end on the last day of data collection in each set (i.e., first labeling), this label is used for training and most of the evaluations in Figure 5.
-The official inspections are continuous, including multiple industry-in-use proprietary closed-source anti-cheats and manual checks.
-The officials update the ban list daily.
-However, due to the long ban cycle of the official inspections, most cheaters will not get banned on the day they cheat.
-Therefore, after 65 days of the first labeling, we once again acquired the labels (i.e., second labeling) of the whole datasets to screen out cheaters who had not been labeled correctly in the first labeling for further analysis on the false positives (the increased cheater numbers are shown in Table 4).
-Although the two datasets contain different types of cheat, the official in-use anti-cheats can not label the specific subcategories, but can only annotate the broad categories of cheats.
-Cheating sophistication is subjective and difficult to quantify.
-Hence, even though we want to provide further information, it is infeasible to provide a more detailed breakdown of the datasets.
-Besides, due to the low recall and accuracy of the current in-use anti-cheats, there will be inevitable noise (cheaters who have evaded the official anti-cheats) in the negative labels within the ground truth.
-Figure 5: Experiments’ process and labeling details.
-6.3 Evaluation Metrics
-Our chosen evaluation metrics are more skewed towards reflecting how well the cheaters can be identified.
-We use the following metrics in Appendix F and Operational Efficiency Index (OEI) to distinguish between cheating and normal behaviors.
-OEI is a new metric for evaluating the efficiency of industrial operations.
-Operational Efficiency Index (OEI), represented in Equation 10, where $N$ represents the total number of players.
-It reflects the operational efficiency gains in manual review processes.
-A higher OEI indicates a greater reduction in manual verification and a more effective system.
-$OEI=\left(\frac{N}{TP+FP}\right)\cdot(Recall\cdot NPV)$ (10)
-6.4 Ablation Study Results
-To answer RQ.1, we conduct an ablation study on Hawk shown in Table 2.
-We set the Task-Specified Threshold Optimizer to achieve the highest accuracy with at least 70% recall for detecting aimbot, and the highest AUC-ROC for detecting wallhack, illustrating different anti-cheat scenarios.
+For aimbot, data is collected over a single 16-day window, while for wallhack, two collection windows totaling 27 days are used (the rationale for this inconsistency is explained in Section VI-G).
+All demos are from real-world ranked matches within these windows, ensuring that the cheats used by players remain black-box to the anti-cheat system.
+Both datasets naturally contain varying levels of cheating sophistication.
+The cheaters’ labels in the first labeling are acquired on the last day of data collection for each set.
+These labels are used for training and most evaluations, as illustrated in Figure 5.
+Official inspections, which include proprietary closed-source anti-cheat systems and manual checks, are conducted continuously, with daily updates to the ban list.
+However, due to the lengthy ban cycle, most cheaters are not banned immediately after cheating.
+To address this, a second labeling is performed 65 days after the first labeling to identify cheaters who were initially mislabeled.
+This allowed us to refine our analysis of false positives, with the updated cheater counts detailed in Table VI.
+Official inspections employ a combination of proprietary anti-cheat software, developed by 5Eplay, and manual verification.
+The anti-cheat system operates by installing a kernel-level driver to scan system memory and active processes for known cheat signatures.
+A significant drawback of this method is its inherent reliance on a blacklist, making it one step behind the cheating developers.
+Modern cheats frequently update their code and employ advanced obfuscation techniques to evade signature-based detection.
+When a potential violation is identified, the system forwards the detection results to human reviewers for final determination.
+Note that despite 5EPlay tried their best to annotate the samples as accurately as possible.
+This process inherently and inevitably introduces a small amount of false negatives (i.e., the cheaters who escape the inspections) after the second labeling due to the current limitation of anti-cheat low performance.
+This is an unsolvable dilemma for this area.
+All positive samples (cheaters) are accurately labeled.
+The increase of cheaters between the first and second labeling is indicated in Table VI’s #IncrPs.
+Based on 5EPlay’s statistics and the case study in Section VI-H, there are approximately 1.5%-2% and 3%-4% uncaught cheaters in the negative samples after the second labeling across different sets in the Aimbot and Wallhack datasets, respectively.
+Although the two datasets include different cheat types, the official in-use anti-cheat systems cannot label specific subcategories; they only annotate broad cheat categories.
+Quantifying cheating sophistication is inherently subjective and challenging, making it infeasible to provide a more detailed breakdown of the datasets.
+Additionally, due to the poor performance of current anti-cheat systems, the ground truth inevitably contains noise (cheaters who evade detection) within the negative labels.
+But to the best of our knowledge, this is the most realistic and well-annotated dataset.
+To evaluate how effectively cheaters are identified, we use the metrics, accuracy, recall, NPV, AUC-ROC, and FPR, defined in Supplemental Material G.
+TABLE VI: The manual average ban duration comparison and the remediation rate among Hawk’s false positives after the first labeling.
+CT Hawk1 Off.2 Validation Test
+#MTPs3 Rmd%4 #IncrPs5 #MTPs Rmd% #IncrPs
+Aimbot $\sim$4 min 5.88 days 93 4.91% 237 97 4.17% 233
+Wallhack 4.47 days 93 5.43% 226 153 6.52% 303
+1Hawk: The process time of Hawk varies with match length, Hawk’s detailed overheads are evaluated in Section VI-I; 2Off.: The average ban duration of official manual inspections per match; 3#MTPs: The number of missed true positives among the false positives; 4Rmd%: The percentage of #MTPs in the number of false positives; 5#IncrPs: The number of increased positives (cheaters) labeled by the officials after the second labeling.
+TABLE VII: Comparison result on the different baseline models of RevPov.
+CT Model Validation Test
+TP TN FP FN Accuracy Recall NPV AUC-ROC FPR TP TN FP FN Accuracy Recall NPV AUC-ROC FPR
+Aimbot REPTree [53] 19 7,660 37 845 0.897 0.022 0.901 0.509 0.005 33 8,118 45 851 0.901 0.037 0.905 0.516 0.006
+RandomTree 123 6,914 783 741 0.822 0.142 0.903 0.520 0.102 125 7,399 764 759 0.832 0.141 0.907 0.524 0.094
+RseslibKNN [54] 40 7,460 237 824 0.876 0.046 0.901 0.508 0.031 33 7,868 295 851 0.873 0.037 0.902 0.501 0.036
+SPAARC [55] 0 7,697 0 864 0.899 0.000 0.899 0.500 0.000 0 8,163 0 884 0.902 0.000 0.902 0.500 0.000
+RandomForest [47] 16 7,665 32 848 0.897 0.019 0.900 0.507 0.004 33 8,128 35 851 0.902 0.037 0.905 0.517 0.004
+Bagging [56] 18 7,666 31 846 0.898 0.021 0.901 0.508 0.004 30 8,129 34 854 0.902 0.034 0.905 0.515 0.004
+VFI [57] 822 445 7,252 42 0.148 0.951 0.914 0.505 0.942 848 472 7,691 36 0.146 0.959 0.929 0.509 0.942
+JRip [58] 21 7,663 34 843 0.898 0.024 0.901 0.510 0.004 42 8,116 47 842 0.902 0.048 0.906 0.521 0.006
+MODLEM [59] 19 7,588 109 845 0.889 0.022 0.900 0.504 0.014 34 8,029 134 850 0.891 0.038 0.904 0.511 0.016
+CSForest [60] 50 7,292 405 814 0.858 0.058 0.900 0.503 0.053 58 7,755 408 826 0.864 0.066 0.904 0.508 0.050
+RevPov 458 4,796 2,901 406 0.614 0.530 0.922 0.577 0.377 481 5,088 3,075 403 0.616 0.544 0.927 0.584 0.377
+Wallhack REPTree 47 8,641 104 920 0.895 0.049 0.904 0.518 0.012 66 9,140 109 959 0.896 0.064 0.905 0.526 0.012
+RandomTree 178 8,006 738 789 0.843 0.184 0.910 0.550 0.084 198 8,433 816 827 0.840 0.193 0.911 0.552 0.088
+RseslibKNN 89 8,515 230 878 0.886 0.092 0.907 0.533 0.026 120 8,939 310 905 0.882 0.117 0.908 0.542 0.034
+SPAARC 69 8,679 66 898 0.901 0.071 0.906 0.532 0.008 87 9,164 85 938 0.900 0.085 0.907 0.538 0.009
+RandomForest 27 8,713 32 940 0.900 0.028 0.903 0.512 0.004 35 9,212 37 990 0.900 0.034 0.903 0.515 0.004
+Bagging 26 8,705 40 941 0.899 0.027 0.902 0.511 0.005 37 9,201 48 988 0.899 0.036 0.903 0.515 0.005
+VFI 950 563 8,182 17 0.156 0.982 0.971 0.523 0.936 1,008 689 8,560 17 0.166 0.983 0.976 0.529 0.925
+JRip 36 8,670 75 931 0.896 0.037 0.903 0.514 0.009 71 9,179 70 954 0.900 0.069 0.906 0.531 0.008
+MODLEM 35 8,687 58 932 0.898 0.036 0.903 0.515 0.007 56 9,171 78 969 0.898 0.055 0.904 0.523 0.008
+CSForest 0 8,745 0 967 0.900 0.000 0.900 0.500 0.000 0 9,249 0 1,025 0.900 0.000 0.900 0.500 0.000
+RevPov 573 6,081 2,664 394 0.685 0.593 0.939 0.644 0.305 627 6,364 2,885 398 0.680 0,612 0.941 0.650 0.312
+TABLE VIII: Hawk detection result with different Task-Specified Threshold Optimizer (TSTO) settings.
+CT TSTO Validation Test
+TP TN FP FN Accuracy Recall AUC-ROC NPV FPR TP TN FP FN Accuracy Recall AUC-ROC NPV FPR
+Aimbot Balance1 389 7,390 307 475 0.909 0.450 0.705 0.940 0.040 457 7,444 719 427 0.873 0.517 0.714 0.946 0.088
+F1-Score 448 7,305 392 416 0.906 0.519 0.734 0.946 0.051 548 6,928 1235 336 0.826 0.620 0.734 0.954 0.151
+Accuracy 269 7,528 169 595 0.911 0.311 0.645 0.927 0.022 305 7,888 275 579 0.906 0.345 0.656 0.932 0.034
+AUC-ROC 648 4,603 3,094 216 0.613 0.750 0.674 0.955 0.402 691 4,364 3,799 193 0.559 0.782 0.658 0.958 0.465
+$\text{Acc}_{\textit{rc}>0.75}$2 652 5,321 2,376 212 0.698 0.755 0.723 0.962 0.309 672 5,365 2,798 212 0.667 0.760 0.709 0.962 0.343
+$\text{Acc}_{\textit{rc}>0.8}$ 700 4,611 3,086 164 0.620 0.810 0.705 0.966 0.401 716 4,700 3,463 168 0.599 0.810 0.693 0.965 0.424
+$\text{Acc}_{\textit{rc}>0.85}$ 736 3,790 3,907 128 0.529 0.852 0.672 0.967 0.508 761 3,875 4,288 123 0.512 0.861 0.668 0.969 0.525
+$\text{Acc}_{\textit{rc}>0.9}$ 778 2,859 4,838 86 0.425 0.900 0.636 0.971 0.628 787 2,955 5,208 97 0.414 0.890 0.626 0.968 0.638
+Wallhack Balance 592 8,256 489 375 0.911 0.612 0.778 0.957 0.056 616 8,344 905 409 0.872 0.601 0.752 0.953 0.098
+F1-Score 576 8,313 432 391 0.915 0.596 0.773 0.955 0.049 651 8,040 1,209 374 0.846 0.635 0.752 0.956 0.131
+Accuracy 374 8,582 163 593 0.922 0.387 0.684 0.935 0.019 167 9,158 91 858 0.908 0.163 0.577 0.914 0.010
+$\text{Acc}_{\textit{rc}>0.7}$ 677 7,898 847 290 0.883 0.700 0.802 0.965 0.097 803 7,506 1,743 222 0.809 0.783 0.797 0.971 0.188
+$\text{Acc}_{\textit{rc}>0.75}$ 727 7,589 1,156 240 0.856 0.752 0.810 0.969 0.132 859 6,964 2,285 166 0.761 0.838 0.795 0.977 0.247
+$\text{Acc}_{\textit{rc}>0.8}$ 776 7,186 1,559 191 0.820 0.802 0.812 0.974 0.178 583 8,280 969 442 0.863 0.569 0.732 0.949 0.105
+$\text{Acc}_{\textit{rc}>0.85}$ 822 6,712 2,033 145 0.776 0.850 0.809 0.979 0.232 871 6,810 2,439 154 0.748 0.850 0.793 0.978 0.264
+$\text{Acc}_{\textit{rc}>0.9}$ 871 5,755 2,990 96 0.682 0.901 0.779 0.984 0.342 855 6,885 2,364 170 0.753 0.834 0.789 0.976 0.256
+[…]
+VI-C Overall Performance, Ablation and Baseline Comparison
+To answer RQ.1, we conduct an ablation study, shown in Table IV.
+We set the TSTO to achieve the highest accuracy with at least 70% recall for detecting aimbot, and the highest AUC-ROC for detecting wallhack (same below unless otherwise stated), illustrating different anti-cheat scenarios (further evaluated in Section VI-F).
 Overall, Hawk excels in identifying authentic cheaters, evidenced by the highest recall and NPV across all evaluations.
@@
 In wallhack detection, RevStats excels in the validation set, while ExSPC leads in the test set.
-Although some components like RevStats and ExSPC outperform Hawk in accuracy, their lower recall (up to 64%) limits their effectiveness, affirming Hawk as the optimal choice.
-Notably, the Wallhack dataset yields better overall performance than the Aimbot dataset.
+Although some components like RevStats and ExSPC outperform Hawk in accuracy and FPR, their lower recall (up to 64%) limits their effectiveness, affirming Hawk as the optimal choice.
+Notably, the Wallhack dataset yields better overall performance than Aimbot.
 We believe this is due to the latter’s diverse subcategories of aimbots.
-In Table 5 and Table 3, we present RevPov’s and RevStats’s model selection results among hundreds of tested models, and promising models are demonstrated.
+We conduct the pair-wise ablation study to evaluate the performance among Hawk and different combinations of subsystems.
+In Table IV, we observe that RevPov and RevStats combination has significantly higher FPR on the Aimbot set and slightly lower recall on the Wallhack set.
+The other two combinations both suffer from lower recall.
+While evaluating the performance, we prioritize recall, accuracy and FPR on the test set over other metrics, because they can more directly reflect the ability to identify cheaters or the overall performance.
+Hawk’s framework is designed with both detection effectiveness and efficiency in mind.
+We select models with acceptable detection performance, and further select models with low training and prediction costs based on the overhead described in the relevant articles.
+Table VII and Table V show RevPov’s and RevStats’s promising model selection results among hundreds of tested baselines.
 RevPov outperforms other models with balanced, consistent accuracy and recall.
-Random Forest outperformed other models under the sampled data in RevPov.
+For RevStats, we found that Random Forest outperformed other models.
+Yet, it sometimes failed to overcome the overfitting problem (e.g., wallhack validation set).
 Thus, we adopt multi-forest and multi-subsampling with Random Forest.
-RevStats outperforms individual classifiers in balancing recall and accuracy by capitalizing on the strengths of each classifier while mitigating their weaknesses.
-The training and validation loss for ExSPC in Figure 9 and Figure 10 (further discussed in Appendix A).
-Table 4: The manual average ban duration comparison and the remediation rate among Hawk’s false positives after the first labeling.
-CT1 Hawk 2 Off.3 Validation Test
-#MTPs4 Rmd%5 #IncrPs6 #MTPs4 Rmd%5 #IncrPs6
-Aimbot $\sim$4 min 5.88 days 93 4.91% 237 97 4.17% 233
-Wallhack 4.47 days 93 5.43% 226 153 6.52% 303
-1CT: Cheat type;
-2
-Hawk: The process time of Hawk varies with match length, Hawk’s detailed overheads are evaluated in Section 6.9;
-3
-Off.: The average ban duration of official manual inspections per match;
-4
-#MTPs: The number of missed true positives among the false positives;
-5
-Rmd%: The percentage of #MTPs in the number of false positives;
-6
-#IncrPs: The number of increased positives (cheaters) labeled by the officials after the second labeling.
-Table 5: Comparison result on the different model selection of RevPov. N/A in OEI indicates the denominator is 0.
-CT1 Model Validation Test
-TP TN FP FN Accuracy Recall NPV AUC-ROC OEI2 TP TN FP FN Accuracy Recall NPV AUC-ROC OEI2
-Aimbot REPTree [21] 19 7,660 37 845 0.897 0.022 0.901 0.509 3.028 33 8,118 45 851 0.901 0.037 0.905 0.516 3.919
-RandomTree 123 6,914 783 741 0.822 0.142 0.903 0.520 1.215 125 7,399 764 759 0.832 0.141 0.907 0.524 1.305
-RseslibKNN [26] 40 7,460 237 824 0.876 0.046 0.901 0.508 1.289 33 7,868 295 851 0.873 0.037 0.902 0.501 0.929
-SPAARC [68] 0 7,697 0 864 0.899 0.000 0.899 0.500 N/A 0 8,163 0 884 0.902 0.000 0.902 0.500 N/A
-RandomForest 16 7,665 32 848 0.897 0.019 0.900 0.507 2.974 33 8,128 35 851 0.902 0.037 0.905 0.517 4.496
-Bagging [6] 18 7,666 31 846 0.898 0.021 0.901 0.508 3.278 30 8,129 34 854 0.902 0.034 0.905 0.515 4.341
-VFI [20] 822 445 7,252 42 0.148 0.951 0.914 0.505 0.922 848 472 7,691 36 0.146 0.959 0.929 0.509 0.944
-JRip [10] 21 7,663 34 843 0.898 0.024 0.901 0.510 3.408 42 8,116 47 842 0.902 0.048 0.906 0.521 4.376
-MODLEM [57] 19 7,588 109 845 0.889 0.022 0.900 0.504 1.323 34 8,029 134 850 0.891 0.038 0.904 0.511 1.873
-CHIRP [65] 0 7,697 0 864 0.899 0.000 0.899 0.500 N/A 0 8,163 0 884 0.902 0.000 0.902 0.500 N/A
-CSForest [52] 50 7,292 405 814 0.858 0.058 0.900 0.503 0.980 58 7,755 408 826 0.864 0.066 0.904 0.508 1.151
-RevPov 458 4,796 2,901 406 0.614 0.530 0.922 0.577 1.246 481 5,088 3,075 403 0.616 0.544 0.927 0.584 1.283
-Wallhack REPTree 47 8,641 104 920 0.895 0.049 0.904 0.518 2.825 66 9,140 109 959 0.896 0.064 0.905 0.526 3.421
-RandomTree 178 8,006 738 789 0.843 0.184 0.910 0.550 1.776 198 8,433 816 827 0.840 0.193 0.911 0.552 1.782
-RseslibKNN 89 8,515 230 878 0.886 0.092 0.907 0.533 2.540 120 8,939 310 905 0.882 0.117 0.908 0.542 2.540
-SPAARC 69 8,679 66 898 0.901 0.071 0.906 0.532 4.652 87 9,164 85 938 0.900 0.085 0.907 0.538 4.599
-RandomForest 27 8,713 32 940 0.900 0.028 0.903 0.512 4.149 35 9,212 37 990 0.900 0.034 0.903 0.515 4.400
-Bagging 26 8,705 40 941 0.899 0.027 0.902 0.511 3.571 37 9,201 48 988 0.899 0.036 0.903 0.515 3.940
-VFI 950 563 8,182 17 0.156 0.982 0.971 0.523 1.014 1,008 689 8,560 17 0.166 0.983 0.976 0.529 1.031
-JRip 36 8,670 75 931 0.896 0.037 0.903 0.514 2.941 71 9,179 70 954 0.900 0.069 0.906 0.531 4.572
-MODLEM 35 8,687 58 932 0.898 0.036 0.903 0.515 3.414 56 9,171 78 969 0.898 0.055 0.904 0.523 3.789
-CHIRP 6 8,738 7 961 0.900 0.006 0.901 0.503 4.176 8 9,242 7 1,017 0.900 0.008 0.901 0.504 4.816
-CSForest 0 8,745 0 967 0.900 0.000 0.900 0.500 N/A 0 9,249 0 1,025 0.900 0.000 0.900 0.500 N/A
-RevPov 573 6,081 2,664 394 0.685 0.593 0.939 0.644 1.670 627 6,364 2,885 398 0.680 0,612 0.941 0.650 1.684
-1CT: Cheat type; 2OEI: Operational Efficiency Index.
-Table 6: Hawk detection result with different Task-Specified Threshold Optimizer settings.
-CT1 Optimizer2 Validation Test
-TP TN FP FN Accuracy Recall AUC-ROC NPV OEI3 TP TN FP FN Accuracy Recall AUC-ROC NPV OEI3
-Aimbot F1-Score 448 7,305 392 416 0.906 0.519 0.734 0.946 5.000 548 6,928 1235 336 0.826 0.620 0.734 0.954 3.000
-Accuracy 269 7,528 169 595 0.911 0.311 0.645 0.927 5.640 305 7,888 275 579 0.906 0.345 0.656 0.932 5.014
-AUC-ROC 648 4,603 3,094 216 0.613 0.750 0.674 0.955 1.639 691 4,364 3,799 193 0.559 0.782 0.658 0.958 1.508
-$\text{Acc}_{\textit{rc}>0.75}$ 652 5,321 2,376 212 0.698 0.755 0.723 0.962 2.052 672 5,365 2,798 212 0.667 0.760 0.709 0.962 1.907
-$\text{Acc}_{\textit{rc}>0.8}$ 700 4,611 3,086 164 0.620 0.810 0.705 0.966 1.769 716 4,700 3,463 168 0.599 0.810 0.693 0.965 1.693
-$\text{Acc}_{\textit{rc}>0.85}$ 736 3,790 3,907 128 0.529 0.852 0.672 0.967 1.519 761 3,875 4,288 123 0.512 0.861 0.668 0.969 1.495
-$\text{Acc}_{\textit{rc}>0.9}$ 778 2,859 4,838 86 0.425 0.900 0.636 0.971 1.333 787 2,955 5,208 97 0.414 0.890 0.626 0.968 1.301
-Wallhack F1-Score 576 8,313 432 391 0.915 0.596 0.773 0.955 5.481 651 8,040 1,209 374 0.846 0.635 0.752 0.956 3.352
-Accuracy 374 8,582 163 593 0.922 0.387 0.684 0.935 6.543 167 9,158 91 858 0.908 0.163 0.577 0.914 5.932
-$\text{Acc}_{\textit{rc}>0.7}$ 677 7,898 847 290 0.883 0.700 0.802 0.965 4.304 803 7,506 1,743 222 0.809 0.783 0.797 0.971 3.071
-$\text{Acc}_{\textit{rc}>0.75}$ 727 7,589 1,156 240 0.856 0.752 0.810 0.969 3.759 859 6,964 2,285 166 0.761 0.838 0.795 0.977 2.675
-$\text{Acc}_{\textit{rc}>0.8}$ 776 7,186 1,559 191 0.820 0.802 0.812 0.974 3.251 583 8,280 969 442 0.863 0.569 0.732 0.949 3.574
-$\text{Acc}_{\textit{rc}>0.85}$ 822 6,712 2,033 145 0.776 0.850 0.809 0.979 2.831 871 6,810 2,439 154 0.748 0.850 0.793 0.978 2.579
-$\text{Acc}_{\textit{rc}>0.9}$ 871 5,755 2,990 96 0.682 0.901 0.779 0.984 2.229 855 6,885 2,364 170 0.753 0.834 0.789 0.976 2.598
-1CT: Cheat type; 2Optimizer: Task-Specified Threshold Optimizer in Mvin; 3OEI: Operational Efficiency Index.
-6.5 Hawk and Official Inspection Comparison
+As a result, RevStats outperforms individual classifiers in balancing recall and accuracy by capitalizing on the strengths of each classifier.
+We also prioritize metrics recall, accuracy and FPR on the test set over other metrics.
+For example, although Logistic Regression has a higher recall, its accuracy and FPR are significantly lower.
+The contribution of the recall’s increase can not make up for the decrease in accuracy and FPR.
+The current framework has optimized overall performance compared with more innovative models.
+For example, in the preliminary test of RevPov’s encoder selection, we obtain lower detection performance with extended training and testing time while using the Transformer.
+Figure 7: Daily number of bans for the day and cumulative ban-sums by Hawk versus the official manual inspections.
+VI-D Hawk and Official Inspection Comparison
 To answer RQ.2, we calculate the daily number of bans for the day and cumulative ban-sums by Hawk and the official inspections illustrated in Figure 7.
-The figure indicates that Hawk’s identification effectiveness (daily count in red bars) and efficiency (cumulative count in red lines) outrank those of the official manual inspections (blue bars and lines) on both cheat types.
-Additionally, Table 4 compares the average ban duration per match, revealing that Hawk’s seconds-scale processing time is exponentially more efficient than official inspections.
+The figure indicates that Hawk’s identification effectiveness (daily count in red bars) and efficiency (cumulative count in solid red lines) outrank those of the official manual inspections (hatched blue bars and dashed blue lines) on both cheat types.
+Additionally, Table VI compares the average ban duration per match, revealing that Hawk’s seconds-scale processing time is exponentially more efficient than official inspections.
 The remediation rate indicates the percentage of cheaters not banned in the first labeling but caught in the second labeling.
-In Table 4, we list the remediation rate after the second labeling among Hawk’s false positives (which are obtained after the first labeling).
-It demonstrates that Hawk can capture the authentic cheaters among the normal players initially labeled by the official anti-cheats.
-This also suggests that while Hawk is robust to mislabeled data, using a purer dataset would further improve results.
-Figure 6: Daily number of bans for the day and cumulative ban-sums by Hawk versus the official manual inspections.
-6.6 Analysis on Different Mvin Settings
-To address RQ.3, we conduct a comparative analysis of Hawk under different Mvin settings, as shown in Table 6.
-The Optimizer column indicates the priority choices of the Task-Specified Threshold Optimizer, such as F1-Score for highest F1-Score and $\text{Acc}_{\textit{rc}>0.75}$ for highest accuracy with recall above 75%.
-An inevitable trade-off exists between capturing more cheaters and minimizing false positives due to cheating sophistication and dataset noise (Section 6.2).
-As discussed in Section 5.4, GMs can dynamically adjust optimizers, allowing for quick retraining of Mvin, with new weight matrices and thresholds generated in about one minute.
-This adaptability aligns with the varying intensity of required anti-cheat measures, offering flexibility in balancing performance and efficiency.
-GMs can use different optimizers to meet specific anti-cheat needs.
-Interestingly, GMs do not always aim to catch all cheaters.
-They sometimes balance catching cheaters with permissiveness, as certain cheaters may bring short-term profits (e.g., in buyout games, where bans may be delayed until after refund periods).
-When complaints rise, GMs may then conduct a mass ban (a regularly organized operation to clean up the gaming environment) to restore community confidence.
-This also partially draws to the creation of Mvin.
-6.7 Robustness to Cheat Evolution
+Table VI lists the remediation rate after the second labeling among Hawk’s false positives (which are obtained after the first labeling).
+It demonstrates that Hawk can capture the authentic cheaters who escape from the official anti-cheat system.
+Moreover, this indicates that Hawk is robust against mislabeled data.
+We also compare the Hawk’s performance with the official in-use anti-cheat in the next section, as it is relevant to Mvin.
+TABLE IX: Hawk comparative study results to prior works on the Aimbot test set. The optimal results are bolded.
+Method Accuracy Recall NPV AUC-ROC FPR
+AccA[11] 0.103 0.997 0.897 0.558 0.997
+HitAcc[12] 0.675 0.545 0.931 0.658 0.311
+Hawk 0.716 0.726 0.960 0.721 0.285
+VI-E Comparative Analysis Among Hawk and Prior Works
+To answer RQ.7, we compare Hawk with two other reproduced methods using Hawk’s Aimbot dataset.
+Due to the lack of available open-source solutions, we write our own code to reproduce both aimbot detection methods based on statistical features related to aiming or elimination used in prior works.
+Both methods are related to Hawk but use naive statistical features, data, and approach.
+Since prior work only targeted aimbots, we only used the Aimbot dataset for comparison.
+The results are listed in Table IX.
+Hawk outperforms prior works across all metrics but AccA’s recall, because AccA classifies most of the samples as cheaters.
+Below is a summary of each method:
+  • AccA: Filters cheaters when the player’s field-of-view acceleration exceeds a threshold [11].
+  • HitAcc: Filters cheaters based on the weapon accuracy of the player, using SVM [12].
+VI-F Analysis on Different Mvin Settings
+To address RQ.3, we conduct a comparative analysis of Hawk under different Mvin settings, as shown in Table VIII and footnote 4.
+In Table VIII, except Mvin, other subsystems are trained and validated with recall priority.
+By only adjusting the TSTO in Mvin, Hawk achieves rapid dynamic tuning ability, by using pre-trained Mvin or retraining Mvin to generate weight matrices and thresholds in about one minute.
+For example, the TSTO settings can be adjusted to achieve a lower FPR if necessary, avoiding the potential surge of excessive manual verification workload.
+Table VIII shows that leveraging different optimizer settings, e.g., Accuracy or F1-Score optimization modes, can obtain low FPR (1%-3% or 5%-15%, respectively).
+Alternatively, the Balance mode maintains a lower FPR than current industry standards (grey bar in footnote 4) while optimizing recall, resulting in slightly lower FPR and significantly higher recall compared to industry anti-cheat system (see blue bars in footnote 4, industry anti-cheats are introduced in Section VIII).
+VI-G Robustness to Cheat Evolution
 Although anti-cheat measures are often opaque to prevent circumvention, cheat developers treat them as black boxes, adapting and camouflaging their cheats on a timely basis to evade detection.
 Cheat evolution refers to the diminishing effectiveness of specific anti-cheat over time.
-Commercially used anti-cheat methods (e.g., signature or process detection, driver scanning, etc.) are susceptible to this evolution since they target specific cheats.
-However, Hawk starts the analysis from the behavior, thus cheaters should appear data-level distinctions mentioned in Section 2.2.
+Commercially used anti-cheat methods are susceptible to this evolution (explained in Section VIII).
+However, Hawk starts the analysis from the behavior, thus cheaters must appear the data-level distinctions mentioned in Section II-B.
 Hence, no matter how the cheats adapt, as long as the cheaters still want to gain unfair advantages, it is hard to bypass the detection of Hawk.
-To address RQ.4, we test Hawk’s robustness to cheat evolution using the Aimbot dataset, divided into partitions (40 matches each).
+To address RQ.4, we test Hawk’s robustness to cheat evolution using the Aimbot dataset.
 Classified through date, the dataset is divided into partitions (40 matches per partition).
+The training and validation are performed on all data up to the current partition.
+We compared its performance on the same test set with different partitions’ trained models.
 The reason for selecting the Aimbot dataset is that during its collection window, an official mass ban on aimbots took place (at the 10th partition), and the cheats are usually updated within a few days after such an incident.
-Figure 7 and Figure 8 demonstrate the ablated performances of subsets under different metrics of Hawk across different numbers of partitions for training.
-According to the average ban duration of the official manual inspections in Table 4, approximately three sequestration ban cycles have passed during the collection window, and at least two ban cycles after the mass ban occurred.
+8(a) and 8(b) demonstrate the ablated performances of subsets under different metrics of Hawk across different numbers of partitions for training.
+According to the average ban duration of the official bans in Table VI, approximately three sequestration ban cycles have passed during the collection window, and at least two ban cycles after the mass ban occurred.
+This offers more than enough time for cheat to evolve.
 The ablated results of Hawk on both sets have been fluctuating to maintain within a small range over partitions (time) and with the cheat evolution.
+This indicates even without using the latter models of partitioned samples after the mass ban, Hawk ’s performance remains stable on the test set before and after the mass ban occurred.
 Besides, in Figure 7, we observe the effectiveness of Hawk has not continuously decreased over time on both datasets (we separated and extended the collection windows in Wallhack dataset for further proof).
-Thus, both of the results indicate Hawk is robust to cheat evolution.
-Figure 7: Ablated robustness against cheat evolution over partitions on Aimbot’s validation set under different metrics.
-Figure 8: Ablated robustness against cheat evolution over partitions on Aimbot’s test set under different metrics.
-6.8 Detecting Uncaught Cheaters in the Wild
-To answer RQ.5, given the above-known presence of cheaters in the false positives in Table 4 after the first labeling, we conduct a use case analysis to explore if there are additional illegal activities that exist after the second labeling.
-Given the lengthy average duration of a match, which demands significant manpower, we randomly select 30 Hawk’s false positives from the Aimbot and Wallhack validation sets for manual review, respectively (experts’ skill level is listed in Table 8).
+Thus, both experiments indicate Hawk is robust to cheat evolution.
+(a) Validation set
+(b) Test set
+Figure 8: Ablated robustness against cheat evolution under different metrics for validation (a) and test (b) sets.
+VI-H Detecting Uncaught Cheaters in the Wild
+To answer RQ.5, given the above-known presence of cheaters in the false positives in Table VI after the first labeling, we conduct a use case analysis to explore if there are additional illegal activities that exist after the second labeling.
+Given the lengthy average duration of a match, which demands significant manpower, we select the top 30 Hawk’s false positives based on the decreasing order of the akpr from the Aimbot and Wallhack validation sets for manual review, respectively (experts’ skill level is listed in Supplemental Material Table 1).
 With the assistance of three experts in FPS games, we identify 27 suspects, who are labeled as normal players after the second labeling, to be highly suspicious within these 60 false positives.
@@
 We reported these suspicions to the official personnel.
-Upon their review, they confirm all 12 of the aforementioned aimbots and wallhacks suspects are committed to cheats, subsequently issuing bans for these cheaters on the platform.
-This reaffirms the poor performance of current in-use anti-cheats and the datasets used for training and evaluation are not entirely pristine.
-The website mentioned in Section 1 demonstrates their illegal actions.
+Upon their review, they confirm all 12 of the aforementioned aimbots and wallhacks suspects are committed to cheats, subsequently issuing bans for these cheaters.
+This reaffirms the poor performance of current in-use anti-cheats.
+The website mentioned in Section I demonstrates their illegal actions.
 In terms of boosting suspects, there are no existing solutions for the officials to evidently verify their illegal actions.
 Therefore, there are no bans issued on the suspicious boosting accounts.
-Table 7: Maximum overheads of Hawk’s subsystems per match. The numbers are retrieved by the Time [36] in Linux.
-Max. Overheads per Match RevPov RevStats ExSPC Mvin
-User time (sec)1 415.72 178.12 36.89 3.97
-System time (sec)2 69.32 436.61 27.12 13.53
-Elapsed time (mm:ss)3 03:40 00:14 00:45 00:01
-Percent of CPU this job got4 220% 4371% 140% 2860%
-Maximum resident set size (GB)5 3.56 0.27 2.57 0.10
-1User time: CPU time spent in user mode;
-2
-System time: CPU time spent in the kernel or system mode;
-3
-Elapsed time: Total real-world time taken to execute the job;
-4
-Percent of CPU this job got: How much of the available CPU resources were allocated to the process, this value is commonly greater than 100% when the process runs on multiple CPU cores;
-5
-Maximum resident set size: The peak amount of physical memory that the process used.
-6.9 Hawk’s Overheads
-Hawk is developed as an asynchronous, load-balancing system that analyzes post-game replay files during periods of low server activity.
+This experiment further proves the limited performance of current industrial anti-cheats, while demonstrating Hawk’s superior anti-cheat capabilities and high robustness to noise.
+VI-I Overheads
+Hawk is designed as an asynchronous, load-balancing system that analyzes post-game replay files during periods of low server activity.
 It operates independently of active game servers and can be deployed on a separate server.
 However, we still conduct overhead assessments for reference.
-To answer RQ.6, we utilize an Ubuntu server equipped with 128 cores AMD EPYC 7713P 64-Core Processor and 256 GB RAM to mimic a real dedicated game server [19].
-By leveraging the pre-trained model in the artifact, Hawk requires only CPU computing resources to function gracefully.
-Table 7 indicates the maximum overhead statistics per match, with each match analyzed concurrently.
-In addition to the basic overheads, we also estimate the Hawk-to-server overheads in terms of a real deployment scenario.
-Our platform partner reports that at most 150,000 matches are played daily.
-Each server needs to run at most 573 matches a day.
-According to our platform’s report and the public server statistics [56], player activity is at its lowest, typically 1-5 AM local time.
-During this timeframe, at most 144 matches need to be analyzed per hour.
-Thus, according to the elapsed time mentioned in Table 7, at most it needs to process 12 matches concurrently to finish processing all matches within the timeframe per server.
+To answer RQ.6, we utilize an Ubuntu server equipped with 128 cores AMD EPYC 7713P 64-Core Processor and 256 GB RAM, a real dedicated game server configuration [61].
+Leveraging the pre-trained model in the artifact, Hawk requires only CPU resources to function efficiently.
+Table X shows the maximum overhead statistics per match, with each match analyzed concurrently.
+Additionally, we estimate the Hawk-to-server overheads in terms of a real deployment scenario.
+Our partner platform reports that at most 150,000 matches are played daily, with each server needing to handle up to 573 matches a day.
+According to our platform’s report and public server statistics [62], player activity is lowest between 1-5 AM local time, requiring analysis of up to 144 matches per hour.
+Based on the elapsed time in Table X, Hawk needs to process up to 12 matches concurrently to complete all analyses within the timeframe per server.
 We further conduct concurrent processing evaluation, it occupies at most 17.95% memory and 35.45% CPU usage for processing 12 matches concurrently.
-Therefore, each server possesses more than sufficient resources for deploying Hawk, even if several matches are hosted at the same time.
-7 Discussions
+Therefore, each server has more than sufficient resources for deploying Hawk, even when hosting several matches simultaneously.
+It is vital for a design to maintain stable performance in high-concurrency environments [63].
+To further validate the performance of Hawk under varying concurrency levels, we employed a game server equipped with a 24-core Intel® Xeon® Gold 5418Y CPUs and 503 GB RAM.
+During idle periods on this server, we concurrently request Hawk 10, 20, and 50 times, respectively.
+The results are listed in Table Table XI.
+The system maintains stable performance when handling different levels of concurrent requests.
+Furthermore, based on the aforementioned server load data, each server requires running Hawk concurrently 12 times per hour at most.
+Our tests reached a maximum concurrency of 50 requests, exceeding the data upper bound by over four times.
+But Hawk requires less than 7 minutes to complete all 50 requests.
+Therefore, we conclude that using Hawk under high concurrency conditions does not significantly affect the system overhead.
+TABLE X: Maximum overheads of Hawk’s subsystems per match. The numbers are retrieved by the Time [64] in Linux.
+Max. Overheads per Match RevPov RevStats ExSPC Mvin
+User time (sec) 415.72 178.12 36.89 3.97
+System time (sec) 69.32 436.61 27.12 13.53
+Elapsed time (mm:ss) 03:40 00:14 00:45 00:01
+Percent of CPU this job got 220% 4371% 140% 2860%
+Maximum resident set size (GB) 3.56 0.27 2.57 0.10
+TABLE XI: Performance under varying concurrency levels.
+# Concurrent Request 10 20 50
+Average time (sec) 384.512 365.267 377.084
+Minimum time (sec) 380.343 357.570 368.070
+Maximum time (sec) 393.472 391.536 386.522
+Standard deviation 4.502 7.793 5.282
+VII Discussions
 Generalizability.
-Narrowed generalizability is a meta-problem for the anti-cheat field.
-Since the cheat is game-specified, the anti-cheat is hard to generalize for all games.
-Although we successfully implemented Hawk on one of the leading tactical shooter games [64] (a subgenre of FPS games like CS:GO), like other prior works [9, 69, 70, 2, 24, 39] discussed in Section 8, it requires further adaptation for other FPS subgenres with distinct gameplay.
-Hawk’s generalizability concerns on the core observations, methodologies, and workflow rather than specific features.
-The features presented are theoretically applicable to tactical shooters in FPS games like the CS series, CrossFire, and others.
+Narrowed generalizability is a meta-problem in the anti-cheat field.
+Since cheats are game-specific, creating a universal anti-cheat system is challenging.
+Although we successfully implemented Hawk on a leading tactical shooter game [65] (a subgenre of FPS games like CS:GO), similar to prior works [6, 10, 11, 12, 13] discussed in Section VIII, it requires further adaptation for other FPS subgenres with distinct gameplay.
+The proposed features are applicable to tactical shooters with different engines in FPS games, e.g., CS2.
 However, adaptation is needed for other subgenres, where certain elements like flashes and grenades may not apply.
-For instance, Hero Shooters (e.g., Overwatch, Valorant) where such features could be substituted with character-specific abilities, or Battle Royale (BR) games (e.g., Fortnite, PUBG) where elements like angles and velocities need substitution due to long time-to-kill.
-Since the replay files and labelings are significant for Hawk, also due to the dataset scarcity limitation, conducting further research on different subgenres to prove the generalizability of Hawk must require cooperation with the game vendors to obtain the data.
-We have initiated contact with several vendors, and one specializing in BR games has shown interest in cooperating.
-But, we have to leave specific implementation and evaluation for future work.
+We summarize two core principles for adapting Hawk to other subgenres.
+(a) Generalized anti-cheat perspectives transform from three observations.
+The fundamental structure of transforming three key observations into three corresponding anti-cheat subsystems is a generalized framework applicable to any FPS subgenre.
+While the specific features or model architectures within each subsystem may need to be optimized for different games, the principle of using these three complementary perspectives to identify cheating behavior is universally necessary and applicable.
+(b) Adaptable feature extraction strategy.
+We propose that feature extraction for any FPS game that uses Hawk should be divided into two distinct categories.
+Core FPS features.
+These are fundamental features inherent to all FPS games and should be retained across all subgenres.
+For example, this includes data like player position, distance to target, and view direction (in RevPov); or various aiming speeds, angles, and precision/hit rates (in RevStats).
+While generalizing Hawk to other subgenres, these features are applicable without any modification.
+Subgenre-specific features.
+These features are designed around the unique mechanics that define a specific subgenre.
+The design of subgenre-specific features, as well as the classification rules in ExSPC, should be led by senior anti-cheat developers with deep expertise in the specific game.
+This ensures the system is tailored to detect cheats that exploit that subgenre’s unique mechanics.
+For instance, in a Hero Shooter (e.g., Overwatch, Valorant), the ability to use unique character skills is a key differentiator.
+To adapt Hawk to other subgenres, one would: i.
+augment RevPov with basic temporal information related to skill activation and targeting; ii.
+extend RevStats to record metrics like skill accuracy, the mean and variance of skill usage timing, and their effectiveness.
 Manual involvement.
-Although Hawk has significantly increased the detection performance and efficiency, as mentioned in Section 3.2, Hawk-after-use manual verification is unavoidable.
-The presence of real false positives undeniably indicates the requirement for manual checks to avoid false bans.
-Nevertheless, with the mitigation of the limitations mentioned in Section 1, we still believe Hawk plays an important role in the FPS games’ server-side anti-cheat.
-8 Related Works
-Client-side detection.
-BotScreen [9], as the latest aimbot detector, utilizes TEE (Intel SGX) to store and run the detection model with aiming features.
-It leverages a self-setup server and hires 14 players to play in Deathmatch instead of the rank mode.
-Thus, BotScreen is limited to dataset scarcity, and its real-world functionality remains unclear, especially with more complex scenarios.
-BotScreen evaluates with a basic, open-source aimbot Osiris [37], which is game-specific and may not reflect the performance against more sophisticated cheats [44, 42].
-VAC [12] can detect wallhack by using OpenGL and scanning such as DNS cache to verify if the suspect has purchased and cheated in a game for Counter-Strike [38, 12].
-Warden [61] scans the system running programs, users’ browser history, and cookies to determine a user’s behavior for Blizzard’s games.
-EAC [22], BattlEye [3], and XIGNCODE3 [62] are commercial anti-cheats.
-They can achieve both cheat detection and prevention [53] for multiple platforms and different games after customization by monitoring and managing process privileges.
-Vanguard [30] is a kernel-level anti-cheat for VALORANT [51] that has full access and control over the system.
-The above client-side anti-cheats are limited to hardware constraints, security and privacy concerns, and client overheads.
-And reflect the limited generalizability problem for anti-cheat.
-Client-side prevention.
-BlackMirror [49] applies SGX for anti-wallhack, providing the client GPU only visible entities based on local predictions.
-It only prevents without detecting and is limited to hardware constraints, client overheads, and targeting on a specific cheat type or subcategories.
-Invisibility Cloak [58] prevents computer-vision-based aimbots from performing object detection with the screen images but is ineffective against memory-based aimbots and wallhacks.
-Server-side ML-based detection.
+Although Hawk has significantly increased the detection performance and efficiency, as mentioned in Section III-B, Hawk-after-use manual verification is unavoidable.
+However, platform statistics indicate their anti-cheat recall (47%) is significantly lower than Hawk’s (79% on average) under recall-first mode.
+Although Hawk ’s false FPR is around 25%, the high recall outweighs the FPR increase (about 15%).
+A higher FNR indicates more cheaters undetected, leading to user attrition with greater losses than those caused by higher FPR [66].
+Hawk also offers Balance mode whose FPR is lower and recall is higher than the current in-use anti-cheat (Section VI-F and footnote 4).
+Under this mode, Hawk outperforms current anti-cheat and obtains lower manual involvement, due to lower FPR.
+Note that for any anti-cheat system, all results must undergo manual review and confirmation before any bans are issued (i.e., the GM verification stage).
+To provide an intuitive demonstration of Hawk’s performance, unless otherwise specified, all results presented for this system are shown before manual review.
+Therefore, all positive samples mentioned in the evaluation require manual review before any bans can be enforced in actual operation.
+No player will be mistakenly banned or inconvenienced before the manual revision.
+Regarding the trade-off between recall and FPR, we have demonstrated the system’s performance under different anti-cheating requirements in Table VIII.
+It can be observed that for every approximately 5% increase in recall, FPR increases by 10% and 20.00% and 20.84% increase of the manpower required on average in terms of the aimbot and the wallhack, respectively.
+Therefore, if human resources are relatively scarce, the default or balanced mode of Hawk should be used.
+Dataset imbalance and result significance.
+To statistically validate the performance stability of our models, we employed bootstrap resampling [67] with 5,000 iterations to estimate 95% confidence intervals for all evaluation metrics in Table XII.
+The bootstrap analysis confirmed the statistical significance and robustness of our models under class-imbalanced conditions.
+All performance metrics showed narrow 95% confidence intervals.
+This demonstrates that both aimbot and wallhack detection performances are not only practically meaningful but also statistically reliable, validating the effectiveness of our approach even with inherent dataset imbalances in the context of game anti-cheating.
+TABLE XII: Performance metrics with 95% confidence intervals from bootstrap analysis (5,000 resamples).
+Dataset Metric Original Bootstrap_Mean Std_Dev Std_Error CI CI_Width
+Aimbot Accuracy 0.716149 0.716142 0.00467 0.000066 [0.706972, 0.725323] 0.018351
+Recall 0.726244 0.726073 0.014762 0.000209 [0.697646, 0.755435] 0.057789
+NPV 0.960191 0.960179 0.00249 0.000035 [0.955263, 0.964983] 0.00972
+AUC-ROC 0.72065 0.720535 0.007924 0.000112 [0.704752, 0.736172] 0.03142
+FPR 0.284944 0.284831 0.004967 0.00007 [0.274992, 0.294452] 0.01946
+Wallhack Accuracy 0.756375 0.756422 0.004182 0.000059 [0.748394, 0.764651] 0.016257
+Recall 0.847805 0.847965 0.011178 0.000158 [0.826003, 0.869812] 0.043808
+NPV 0.977897 0.977882 0.001741 0.000025 [0.97437, 0.981204] 0.006834
+AUC-ROC 0.797024 0.797097 0.006064 0.000086 [0.785037, 0.80896] 0.023923
+FPR 0.253757 0.253778 0.004606 0.000065 [0.244804, 0.262858] 0.018054
+VIII Related Works
+Server-side anti-cheats.
 Yeung et al.
-[69] initially proposed a Bayesian model to tune the probability threshold with basic statistics in 2006, but it is outdated for today’s advanced cheating techniques.
-Subsequent works [70, 2, 24, 39] used machine learning with basic statistics or limited features, which couldn’t monitor sophisticated behaviors or describe entire matches [9].
-These earlier works are limited to naive features and workflow, unclear performance, and dataset scarcity.
-They require extra modification (specified to the game) on the game engine to collect and construct features in parallel with the game commencing.
-Therefore, it is also hard to generalize their method to other games.
-9 Conclusion
+[10] initially proposed a Bayesian model to tune the probability threshold with naive features (e.g., whether the player is moving, the distance between the attacker and victim) in 2006, but its performance lacks real-world validation and the features are outdated for today’s advanced cheating techniques and sophistication.
+Subsequent works [11, 12, 13, 31, 30] employed statistical tests with basic features (e.g., time-on-target) to examine data distribution difference [11], set fixed threshold with limited features (e.g., extra shots count on an invisible bait) [13, 30], or used one machine learning model (e.g., SVM) with basic non-temporal features (e.g., whether the target’s position changed) [12], which couldn’t monitor sophisticated behaviors or describe entire matches [6], or ineffective to certain type of cheat.
+Due to the scarcity of real-world datasets, these server-side studies relied on private servers and simulated data for evaluation, with the largest dataset containing only 47 cheaters [13].
+Such small-scale datasets are insufficient to reflect real-world anti-cheat performance and possibly lead to inconsistent performance between the reported results and real-world performance.
+Additionally, works like [11, 12, 13, 31] require game engine modifications for concurrent data collection, necessitating significant adaptation efforts after each update and limiting generalizability to other games.
+Importantly, prior works leverage a single model to classify and could result in biased predictions on the complex real-world dataset as discussed in Section V-B.
+CUPID [68] is a matchmaking design to optimize the user’s gameplay experience.
+In terms of anti-cheat, it may be modified to arrange cheaters within the same match so that they would not appear in the normal matches.
+However, its applicability remains unknown since it does not target identifying cheaters.
+Client-side anti-cheats.
+BotScreen [6], the latest aimbot detector, uses SGX to securely store and run its detection model with temporal aiming features.
+It constructs its dataset using a self-setup server with 14 players in Deathmatch (gameplay is different from Rank) and evaluates with the deprecated open-source aimbot Osiris [69].
+This setup may not accurately reflect performance against more sophisticated cheats or common cheating scenarios [70, 71], leaving its real-world functionality unclear.
+It is also troubled by the deprecated TEE constraint [14, 15].
+Invisibility Cloak [8] and BlackMirror [9] only prevent without detecting and are limited to hardware constraints or client overheads, and targeting on a specific cheat subcategory.
+Commercial anti-cheats like VAC [18], Vanguard [35], EAC [16] and BattlEye [17] offer both detection and prevention [72] for various games through customization.
+They monitor and manage process privileges, scan DNS caches and cookies for browsing history and cheat purchase records [73, 18], and may use kernel-level drivers [33].
+However, these solutions are limited by security and privacy concerns mentioned in Section I.
+The common approach of creating blacklists for known cheats and issuing bans based on detected malware signatures [38, 33, 6, 13] is ineffective against new or updated cheats with different signatures [6, 13], lacking robustness to cheat evolutions.
+Most cheaters avoid using legacy cheats as paid cheats often offer auto-updates to evade bans, leading to low recall in industry anti-cheats, which are always one step behind.
+These anti-cheats are integrated into games before release and must run concurrently with gameplay [33], causing client overheads and limited generalizability.
+They also combine multiple methods for cross-checking, extending the banning cycle [33].
+Importantly, these commercial anti-cheats are proprietary and closed-source [33], making it infeasible for researchers to conduct comparison studies without collaborating with game vendors.
+VESPA [74] combines computer vision techniques and human inspection to capture the user’s screen and detect the wallhack’s UI overlays and thereby identify cheaters.
+This may raise privacy concerns and can be easily bypassed by managing process privileges or disabling the visible overlays.
+Therefore, it is less practical for real-world deployment.
+FCDGame [75] leverages few-shot learning to learn highly repetitive gesture trajectory patterns on the mobile phone screen to identify cheaters.
+This pattern targets the bot identification in MMORPGs rather than FPS games.
+Because the former only requires fixed navigation and simple gesture operations to achieve repetitive tasks (e.g., collect mines), but the latter requires real-time and complex inputs to adjust aiming and perform shooting.
+The FPS trajectory patterns are more complex and irregular.
+Therefore, the screen gesture trajectory feature may be less functional in FPS games.
+IX Conclusion
 We proposed an FPS games’ server-side anti-cheat framework Hawk, drawing inspiration from how humans identify cheaters.
@@
 Ultimately, we hope Hawk can be a pathfinder for next-generation FPS anti-cheat, providing a valuable dataset and a reference for the follow-up research.
-10 Ethics Considerations
-The dataset only contains information including in-game operations and data, avatars’ IDs, and the banning information, which are all publicly accessible [^4] Link.
-We anonymize the platform’s URL and name..
-We will anonymize all information that may disclose the player’s privacy, i.e., the player ID.
-The information in the published dataset will not directly or indirectly cause any form of impact or damage to the game and/or to its distributors, and/or to the players and/or the community.
-The data provider approves the release and reuse of the dataset for research purposes.
-No private information will be disclosed through our dataset.
-The studies were conducted in accordance with the local legislation and institutional requirements.
-11 Compliance to Open Science Policy
-We open-source Hawk and our datasets [^5] Link.
-The repository and dataset will be released upon acceptance.
-Artifacts are uploaded for the reference.
-Ethics statement is shown in Section 10.
-to support open science.
-We will make all links publicly available after paper acceptance.
-Artifacts will be submitted to the Artifact Evaluation Committee for further review.
-Appendix A Training and Validation Loss for ExSPC Model
-The training process on each of the datasets in ExSPC utilizes the validation set for optimization.
-Figure 9 and Figure 10 demonstrate the training and validation loss for each dataset respectively.
-It is noticeable that the individual results are not ideal.
-Thus, we use the multi-view design to complement one another of the subsystems.
-Figure 9: Training and validation loss of ExSPC on Wallhack dataset.
-Figure 10: Training and validation loss of ExSPC on Aimbot dataset.
-Appendix B Temporal Features
-B.1 Engagement features
-Engagement features are extracted at each tick of the event occurrence. The introduction of the Engagement features is as follows:
-  • Damage features:
-    • tick: the tick number;
-    • seconds: the time elapsed within a round in second;
-    • attackerSteamID, victimSteamID: the ID of the player;
-    • attackerSide, victimSide: the team side of the player;
-[…]
-[…]
-    • attackerStrafe: the Boolean value indicates whether the attacker is moving sideways relative to their forward direction;
-    • weapon: the ID of the damage-caused weapon;
-    • weaponClass: the class ID of the damage-caused weapon;
-    • hpDamage: the value of damage caused to the victim;
-    • hpDamageTaken: the value of damage taken by the victim;
-    • armorDamage: the value of the damage caused to the victim’s armor;
-    • armorDamageTaken: the value of damage taken by the victim’s armor;
-    • hitGroup: the ID of the damage-taken area on the body of the victim;
-    • isFriendlyFire: the Boolean value indicates whether the damage is from the friendly;
-    • distance: the distance between the attacker and the victim;
-    • zoomLevel: the zoom level of the attacker’s weapon;
-    • roundNum: the round number of the current tick.
-  • Auxiliary props features: The tick, second, IDs, sides, location, view direction, and round number are similar to above;
-    • flashDuration: the time elapsed for the blinding effect affects on the victim in second;
-  • Offensive props features: The ID, side, location, view direction of the thrower and the prop, and round number are similar to above;
-    • throwTick, destroyTick: the tick of the prop thrown or destroyed;
-    • throwSeconds, destroySeconds: the time elapsed within a round in seconds at the moment of the prop thrown or destroyed;
-  • Elimination features: The tick, second, IDs, sides, location, view direction, distance, weapon, weapon class, and round number are similar as above;
-    • assisterSteamID: the ID of the assister;
-    • assisterSide: the team side of the assister;
-    • isSuicide: the Boolean value indicates whether the elimination is a suicide event;
-    • isTeamkill: the Boolean value indicates whether the elimination is a team kill event;
-    • isWallbang: the Boolean value indicates whether the elimination is conducted by occluder penetration;
-    • penetratedObjects: the number of obstacles the bullet penetrated;
-    • isFirstKill: the Boolean value indicates whether the elimination is the first kill within a round;
-    • isHeadshot: the Boolean value indicates whether the elimination is a head-shot event;
-    • victimBlinded: the Boolean value indicates whether the victim is blinded;
-    • attackerBlinded: the Boolean value indicates whether the attacker is blinded;
-    • flashThrowerSteamID: the ID of the flash-thrower;
-    • flashThrowerSide: the team side of the flash-thrower;
-    • noScope: the Boolean value indicates whether the elimination is conducted with no scoped;
-    • thruSmoke: the Boolean value indicates whether the elimination is conducted through smoke;
-    • isTrade: the Boolean value indicates whether the elimination is traded with the opponent;
-  • Weapon fire features: The IDs, sides, location, view direction, weapon, weapon class, player strafe, zoom level of the attacker, and the tick, second, and round number are similar as above;
-    • ammoInMagazine: the number of ammunition in the magazine;
-    • ammoInReserve: the number of ammunition in the reserve;
-B.2 Movement features
-Movement features are extracted per tick.
-The IDs, sides, location, view direction of the player, and the tick, second, and round numbers are similar to those above.
-The rest of the introduction of the Movement features is as follows:
-  • velocityX, velocityY, velocityZ: the velocity of the player in 3D space;
-  • isAlive: the Boolean value indicates whether the player is alive;
-  • isBlinded: the Boolean value indicates whether the player is blinded;
-  • isAirborne: the Boolean value indicates whether the player is feet dangling;
-  • isDucking: the Boolean value indicates whether the player is ducking;
-  • isDuckingInProgress: the Boolean value indicates whether the player is ducking in progress;
-  • isUnDuckingInProgress: the Boolean value indicates whether the player is standing up from ducking in progress;
-  • isDefusing: the Boolean value indicates whether the player is defusing the bomb;
-  • isPlanting: the Boolean value indicates whether the player is planting the bomb;
-  • isReloading: the Boolean value indicates whether the player is reloading;
-  • isInBombZone: the Boolean value indicates whether the player is in the bomb planting area;
-  • isStanding: the Boolean value indicates whether the player is standing;
-  • isScoped: the Boolean value indicates whether the player is scoped with the weapon;
-  • isWalking: the Boolean value indicates whether the player is walking;
-  • IsolationDegree: the distance to the calculation center of each teammate.
-B.3 Economy features
-Economy features are extracted per round. The introduction of the Economy features are as follows:
-  • round: the number of round;
-  • equipmentValueFreezetimeEnd: the value of the player’s equipment after the end of the freeze time (the interval from the commencement of a round to the moment players are granted the ability to move and fire.
-This period typically allows players to purchase equipment);
-  • equipmentValueRoundStart: the value of the player’s equipment at the start of each round;
-  • cash: the amount of money the player obtains at the end of each round;
-  • cashSpendTotal: the amount of money the player spends within each round.
-Appendix C Structured Features
-In this section, the extracted data from demos are further processed to create strong-expressiveness features that can indicate the sense and performance of a player clearly.
-For notations used in the formulas throughout this section, please refer to Table 9.
-Table 9: Notations and descriptions of structured features.
-Notation Description Notation Description
-$i$ the current player $\textbf{E}_{\textit{x}}^{a\rightarrow b}[k]$ the $k^{th}$ type x event conducted by player a to b
-$p,p^{*}$ a particular player, any player #()#\#($\bullet$)# ( ∙ ) the number of counts of the parameter
-$o,o^{*}$ a particular opponent, any opponent $wp(\bullet)$ the weapon used by a player
-$a,a^{*}$ a particular ally, any ally $hg(\bullet)$ the body part hit by a player
-$ts$ through smoke $\textbf{{t}}(\bullet)$ the moment of the event occurs in second
-$p(\bullet)$ the number of occluders penetrated $mag(\bullet)$ the number of bullets remain in a player’s magazine
-$\Delta(\bullet)$ the change in the parameter $dist(x,y)$ the distance range between entities x and y
-$\textbf{ELAPSE}(\bullet)$ the duration from the start to the end of an event $\textbf{HEG}(\bullet)$ high-explosive grenade is applied by a player
-$\textbf{INC}(\bullet)$ incendiary or Molotov is applied by a player $\textbf{SMK}(\bullet)$ smoke grenade is applied by a player
-$\textbf{FLS}(\bullet)$ flash grenade is applied by a player $\textbf{FK}(\bullet)$ first kill is accomplished by a player
-Figure 11: Illustrations of reaction and adjustment features. The black rectangle overlay in c2 hides the player’s ID.
-C.1 Aiming features
-Durations
-Figure 11 illustrates the three distinct moments that form three different durations.
-In terms of the moments, $\textbf{{t}}_{0}$ denotes the moment that the current player initially spots one opponent.
-In Figure 11, a1 and a2 present a possible set of states at $\textbf{{t}}_{0}$.
-In which a1 represents a 3-dimensional world in the game map, where the opponent happens to appear at the edge of the current player’s field of view (FOV).
-And a2 demonstrates the 2-dimensional state correspondingly which is typically displayed through the player’s monitor, where the opponent also happens to be spotted at the edge of the current player’s POV.
-$\textbf{{t}}_{1}$ denotes the moment that the current player notices the hostile movement and initiates responsive action (i.e., open fire).
-Note that at $\textbf{{t}}_{1}$ the current firing event may or may not be equivalent to a damage event upon the opponent.
-$\textbf{{t}}_{2}$ denotes the moment that the current player performs a damage event upon the opponent within one period of valid engagement for the first time.
-In our case, the threshold for resetting the valid engagement period detection (i.e., no opponent in sight, no damage or firing event) is set as 10 seconds.
-In Figure 11, it demonstrates a series of generic circumstances that often happen in games where the positions of the current player and the opponent vary along from $\textbf{{t}}_{0}$ to $\textbf{{t}}_{2}$.
-And it is common that sometimes $\textbf{{t}}_{1}$ and $\textbf{{t}}_{2}$ are overlapped.
-In terms of durations, the elapsed period from $\textbf{{t}}_{0}$ to $\textbf{{t}}_{1}$ is defined as reaction time (rat).
-The elapsed period from $\textbf{{t}}_{1}$ to $\textbf{{t}}_{2}$ is defined as adjustment time (ajt).
-The elapsed period from $\textbf{{t}}_{0}$ to $\textbf{{t}}_{2}$ is defined as duration time (drt).
-The average and variance calculations are applied after obtaining the data.
-The average performance of all samples is obtained through the average, and the degree of dispersion of the samples is represented by the variance.
-The integration of average and variance is able to represent the overall performance and stability of the current player within a match.
-rat average. Denotes the average performance regarding rat of the current player among one match. The lower this feature is, the faster the player reacts.
-$\mathbf{F_{\textbf{rat-avg}}(i)}=\overline{RAT}=\frac{\sum_{j=1}^{N}RAT_{j}}{N}$ (11)
-where $\text{where }N\text{ denotes }\#(RAT)\text{ within a match.}$.
-rat variance. Denotes the degree of dispersion of rat distributions of the current player among one match.
-$\mathbf{F_{\textbf{rat-var}}(i)}=Var(RAT)=\frac{\sum_{j=1}^{N}\left(RAT_{j}-\overline{RAT}\right)^{2}}{N}$ (12)
-where $\text{where }N\text{ denotes }\#(RAT)\text{ within a match.}$.
-ajt average. Denotes the average performance regarding ajt of the current player among one match. The lower this feature is, the faster the player adjusts.
-$\mathbf{F_{\textbf{ajt-avg}}(i)}=\overline{AJT}=\frac{\sum_{j=1}^{N}AJT_{j}}{N}$ (13)
-where $\text{where }N\text{ denotes }\#(AJT)\text{ within a match.}$.
-ajt variance. Denotes the degree of dispersion of ajt distributions of the current player among one match.
-$\mathbf{F_{\textbf{ajt-var}}(i)}=Var(AJT)=\frac{\sum_{j=1}^{N}\left(AJT_{j}-\overline{AJT}\right)^{2}}{N}$ (14)
-where $\text{where }N\text{ denotes }\#(AJT)\text{ within a match.}$.
-drt average. Denotes the average performance regarding drt of the current player among one match. The lower this feature is, the faster the player hits.
-$\mathbf{F_{\textbf{drt-avg}}(i)}=\overline{DRT}=\frac{\sum_{j=1}^{N}DRT_{j}}{N}$ (15)
-drt variance. Denotes the degree of dispersion of drt distributions of the current player among one match.
-$\mathbf{F_{\textbf{drt-var}}(i)}=Var(DRT)=\frac{\sum_{j=1}^{N}\left(DRT_{j}-\overline{DRT}\right)^{2}}{N}$ (16)
-where $\text{where }N\text{ denotes }\#(DRT)\text{ occurrences within a match.}$.
-Figure 12: Complete box plots of comparative visualization between honest players and cheaters on all structured features under Mann-Whitney U (Wilcoxon Rank-Sum) test with P-Value ascending (distinction descending) order.
-Velocity
-velocity average.
-Denotes the average mouse velocity of the current player among one match between $\textbf{{t}}_{0}$ and $\textbf{{t}}_{2}$, through which way two features of distance and time period can be highly compressed into one feature.
-The higher the feature is, the faster the player hits the targets.
-$V=\frac{dist(i,o)}{DRT_{i}}$ (17)
-where $\text{where }o\text{ denotes the victim in the specified }DRT\text{.}$
-$\mathbf{F_{\textbf{v-avg}}(i)}=\overline{V}=\frac{\sum_{j=1}^{N}V_{j}}{N}$ (18)
-where $\text{where }N\text{ denotes }\#(DRT)\text{ within a match.}$.
-velocity variance. Denotes the degree of dispersion of velocity distributions of the current player among one match.
-$\mathbf{F_{\textbf{v-var}}(i)}=Var(V)=\frac{\sum_{j=1}^{N}\left(V_{j}-\overline{V}\right)^{2}}{N}$ (19)
-where $\text{where }N\text{ denotes }\#(DRT)\text{ within a match.}$.
-Angles
-Figure 11 illustrates the three different angles, in which $\theta$ denotes divergence of aiming is extracted by Liu et al.
-[39] and represents the initial angular divergence between the aiming position and the opponent position.
-However, the time complexity for obtaining such a feature is too expensive to be accepted for industrial implementation.
-Therefore we extracted the following features which not only preserve strong expressiveness but also significantly reduce the time cost, as well as innovatively segment each engagement process into three stages.
-$\delta$ represents the reaction angle (raa), which reflects the angle of two cross-hair directions at $\textbf{{t}}_{0}$ and $\textbf{{t}}_{1}$ respectively.
-$\gamma$ represents the adjustment angle (aja), which, similarly reflects the angle of two cross-hair directions at $\textbf{{t}}_{1}$ and $\textbf{{t}}_{2}$ respectively.
-The larger $\delta$ and $\gamma$ are the longer distance the cross-hair travels, which ultimately results in performance divergence.
-Similarly, the average and variance operation is applied accordingly.
-raa average.
-Denotes the average performance regarding raa of the current player among one match.
-The lower the feature is, the smaller the angle and the less effort the player aims for and puts into reaching their targets.
-$\mathbf{F_{\textbf{raa-avg}}(i)}=\overline{RAA}=\frac{\sum_{j=1}^{N}RAA_{j}}{N}$ (20)
-where $\text{where }N\text{ denotes }\#(RAA)\text{ within a match.}$.
-raa variance. Denotes the degree of dispersion of raa distributions of the current player among one match.
-$\mathbf{F_{\textbf{raa-var}}(i)}=Var(RAA)=\frac{\sum_{j=1}^{N}\left(RAA_{j}-\overline{RAA}\right)^{2}}{N}$ (21)
-where $\text{where }N\text{ denotes }\#(RAA)\text{ within a match.}$.
-aja average.
-Denotes the average performance regarding aja of the current player among one match.
-The lower the feature is, the smaller angle and the less effort the player fires for and puts into hitting their targets.
-$\mathbf{F_{\textbf{aja-avg}}(i)}=\overline{AJA}=\frac{\sum_{j=1}^{N}AJA_{j}}{N}$ (22)
-where $\text{where }N\text{ denotes }\#(AJA)\text{ within a match.}$.
-aja variance. Denotes the degree of dispersion of aja distributions of the current player among one match.
-$\mathbf{F_{\textbf{aja-var}}(i)}=Var(AJA)=\frac{\sum_{j=1}^{N}\left(AJA_{j}-\overline{AJA}\right)^{2}}{N}$ (23)
-where $\text{where }N\text{ denotes }\#(AJA)\text{ within a match.}$.
-C.2 Firing features
-Inertial shot [39] percentage (isp)
-Inertial shot denotes that after an elimination is conducted by a player, who may apply several extra firings due to the inertia.
-isp represents the ratio of the number of Inertial shot times to the number of eliminations.
-$\alpha$ is a threshold and set as 0.15 seconds in our case.
-$\operatorname{\textbf{IS}}\left(\textbf{E}_{\textit{kill}}^{i\rightarrow o^{*}}[k]\right)=\left\{\begin{array}[]{ll}1,&\exists\textbf{E}_{\textit{fire}}^{i}\in\left(\textbf{{t}}\left(\textbf{E}_{\textit{kill}}^{i\rightarrow o^{*}}[k]\right),\textbf{{t}}\left(\textbf{E}_{\textit{kill}}^{i\rightarrow o^{*}}[k]\right)+\alpha\right)\\ 0,&\text{otherwise}\end{array}\right.$
-$\mathbf{F_{\textbf{isp}}(i)}=\frac{\sum_{k}\operatorname{\textbf{IS}}\left(\textbf{E}_{\textit{kill}}^{i\rightarrow o^{*}}[k]\right)}{\sum_{n}\textbf{E}_{\textit{kill}}^{i\rightarrow o^{*}}[n]}$ (25)
-First hit percentage (fhp)
-Denote how accurate a player is while spotting an opponent and reacting to fire.
-Event reload represents the action of changing the magazine while the firing is performed and the bullet is consumed.
-Fire Round denotes the unit of the firing round.
-fhp refers to the ratio of the number of the first firing event equivalent to a damage event while spotting at least one opponent within a time threshold $\varepsilon$ to the number of times that the aforementioned event occurs while disregarding the equivalency and within the same time threshold $\varepsilon$.
-In our case, $\varepsilon$ is set as 5 seconds (i.e., 640 ticks).
-$\textbf{E}_{\textit{reload}}^{i}=\Delta\left(\textit{mag}(i)\right)>0\land\textbf{{t}}(\textbf{E}_{\textit{fire}}^{i}[k])-\textbf{{t}}(\textbf{E}_{\textit{fire}}^{i}[k+1])>0$ (26)
-$\textbf{FR}[i]={\{\textbf{E}_{\textit{fire}}^{i}[i]\mid\exists\textbf{E}_{\textit{reload}}^{i}\lor\textbf{{t}}(\textbf{E}_{\textit{fire}}^{i}[i][k+1])-\textbf{{t}}(\textbf{E}_{\textit{fire}}^{i}[i][k])>\varepsilon\}}$ (27)
-$\mathbf{F_{\textbf{fhp}}(i)}=\frac{\#\left(\{\textbf{FR}[i][0]\mid\forall\textbf{FR}[i][0]=\textbf{E}_{\textit{damage}}^{i\rightarrow o^{*}}\}\right)}{\#\left(\textbf{FR}[i][0]\right)}$ (28)
-Precision (precis)
-Denote the performance of firing accuracy. precis refers to the ratio of the number of damage caused by weapons to the total number of fires applied.
-$\mathbf{F_{\textbf{precis}}(i)}=\frac{\sum_{k}\textbf{E}_{\textit{damage}}^{i\rightarrow o^{*}}[k]}{\sum_{n}\textbf{E}_{\textit{fire}}^{i}[n]}$ (29)
-Strafing hit percentage (shp)
-Denote fire accuracy while strafing.
-Strafe allows a player to keep the view focused on an opponent while moving in a different direction.
-shp refers to the ratio of the number of damage caused while strafing to the number of damage caused.
-$\mathbf{F_{\textbf{shp}}(i)}=\frac{\sum_{k}\textbf{E}_{\textit{damage}\land\textit{strafe}}^{i\rightarrow o^{*}}[k]}{\sum_{n}\textbf{E}_{\textit{damage}}^{i\rightarrow o^{*}}[n]}$ (30)
-Hit-group distribution (hgd) variance
-Hit group represents the body part that suffers damage, typically hit groups can be categorized as follows: chest, generic, head, neck, left arm, right arm, left leg, right leg, and stomach.
-hgd variance denotes the degree of dispersion of hit group distributions for the current player among one match.
-$HG_{j}=\sum_{k}\textbf{E}_{\textit{damage}\land\textit{hg(j)}}^{i\rightarrow o^{*}}[k]$ (31)
-$\overline{HG}=\frac{\sum_{j=chest}^{stomach}HG_{j}}{\sum_{k}\textbf{E}_{\textit{damage}}^{i\rightarrow o^{*}}[k]}$ (32)
-$\mathbf{F_{\textbf{hgd}}(i)}=Var(HG)=\frac{\sum_{j=chest}^{stomach}\left(HG_{j}-\overline{HG}\right)^{2}}{\sum_{n}\textbf{E}_{\textit{damage}}^{i\rightarrow o^{*}}[n]}$ (33)
-[…]
-[…]
-Special hit Ratio (shr)
-Special hit denotes hit with sniper between 50 and 150 meters (SH50-150), head shot with sniper between 40 and 170 meters (SHS40-170), hit with normal weapon farther than 800 meters (NH800+) and headshot with normal weapon farther than 700 meters (NHS700+).
-In FPS games, hitting processes include using different weapons and shooting from different distances.
-In our game environment, the aforementioned four types of hits are considered to be potentially cheating combinations.
-shr denotes the ratio of the sum of the number of the aforementioned four types of special hit to the number of hits in one match.
-$\mathbf{\textbf{SH}_{50-150}(i)}=\sum_{k}\textbf{E}_{\textit{damage}\land{sniper}\land{(50<=dist(i,o)}<=150)}^{i\rightarrow o^{*}}[k]$
-$\mathbf{\textbf{SHS}_{40-170}(i)}=\sum_{k}\textbf{E}_{\textit{headshot}\land{sniper}\land{(40<=dist(i,o)}<=170)}^{i\rightarrow o^{*}}[k]$
-$\mathbf{\textbf{NH}_{800+}(i)}=\sum_{k}\textbf{E}_{\textit{damage}\land{normal}\land{(dist(i,o)}>=800)}^{i\rightarrow o^{*}}[k]$ (36)
-$\mathbf{\textbf{NHS}_{700+}(i)}=\sum_{k}\textbf{E}_{\textit{headshot}\land{normal}\land{(dist(i,o)}>=700)}^{i\rightarrow o^{*}}[k]$ (37)
-$\mathbf{F_{\textbf{shr}}(i)}=\frac{\mathbf{\textbf{SH}_{50-150}(i)}+\mathbf{\textbf{SHS}_{40-170}(i)}+\mathbf{\textbf{NH}_{800+}(i)}+\mathbf{\textbf{NHS}_{700+}(i)}}{\sum_{k}\textbf{E}_{\textit{damage}}^{i\rightarrow o^{*}}[k]}$
-C.3 Elimination features
-Time to kill (ttk)
-Denote the duration from the moment of the first damage caused to an opponent to the moment of the elimination of the opponent.
-In order to control the ttk to be each time one player attacks one enemy, the threshold of one statistical period of ttk is set to be 10 seconds.
-For instance, if the attacking period time is over 10 seconds without killing the victims, this action time will not be recorded as ttk.
-$\mathbf{F_{\textbf{ttk}}(i)}=\operatorname{\textbf{ELAPSE}}\left(\textbf{E}_{\textit{damage}}^{i\rightarrow o}[0],\textbf{E}_{\textit{kill}}^{i\rightarrow o}\right)$ (39)
-ttk average. Denotes the average time regarding ttk of the current player among one match. The lower this feature is, the shorter the time the player kills victims.
-$\mathbf{F_{\textbf{ttk-avg}}(i)}=\overline{TTK}=\frac{\sum_{j=1}^{N}TTK_{j}}{N}$ (40)
-where $\text{where }N\text{ denotes }\#(TTK)\text{ within a match.}$.
-ttk variance. Denotes the degree of dispersion of ttk distributions of the current player among one match.
-$\mathbf{F_{\textbf{ttk-var}}(i)}=Var(TTK)=\frac{\sum_{j=1}^{N}\left(TTK_{j}-\overline{TTK}\right)^{2}}{N}$ (41)
-where $\text{where }N\text{ denotes }\#(TTK)\text{ within a match.}$.
-First-blood percentage (fbp)
-first-blood denotes the very first elimination in one round. fbp denotes the ratio of the number of first-blood to the number of rounds in one match.
-$\mathbf{F_{\textbf{fkp}}(i)}=\frac{\#\left(\textbf{FK}(i)\right)}{\#\left(round\right)}$ (42)
-Onetap percentage (otp)
-Onetap denotes a head-shot fire event on an opponent that results in elimination without the use of sniper rifles, and where the inflicted damage is equal to or greater than 100 (the full health of a player).
-The otp represents the ratio of the number of Onetap instances to the number of times damage is inflicted by normal weapons.
-$\operatorname{\textbf{ONETAP}}\left(\textbf{E}_{\textit{kill}}^{i\rightarrow o^{*}}[k]\right)=\left\{\begin{array}[]{ll}1,&\left(\textbf{E}_{\textit{fire}}^{i}=\textbf{E}_{\textit{kill}}^{i\rightarrow o^{*}}[k]\right)\land\left(\textbf{E}_{\textit{damage}}^{i\rightarrow o^{*}}\geq 100\right)\\ 0,&\text{otherwise}\end{array}\right.$
-$\mathbf{F_{\textbf{otp}}(i)}=\frac{\sum_{k}\operatorname{\textbf{ONETAP}}\left(E_{\text{kill}}^{i\rightarrow o^{*}}[k]\right)}{\sum_{n}\textbf{E}_{\textit{damage}\land wp(i)\neq sniper\ rifle}^{i\rightarrow o^{*}}[n]}$ (44)
-Occluder-penetration index (opi)
-Occluder-penetration denotes the elimination event that happens when the victim is eliminated behind one layer or more layers of occluders by a penetrated bullet.
-The occluders are typically walls, players (namely wall bang), or smoke.
-opi refers to a weighted index that reflects the extent of the average occluder-penetration eliminations conducted by the player.
-The use of the number of rounds within the division aims to eliminate bias brought by that.
-In our case, the weights in $\alpha_{i}$ are 0.5, 1, 2, 0.5 respectively.
-$\mathbf{F_{\text{opi}}(i)}=\frac{\mathbf{W_{i}}\alpha_{i}}{\#\left(\textit{round}\right)}$ (45)
-$\text{where}\ \alpha_{i}=\begin{bmatrix}w_{1}\\ w_{2}\\ w_{3}\\ w_{4}\end{bmatrix},$
-${\mathbf{W_{i}}}=\begin{bmatrix}\sum_{k}\textbf{E}_{\textit{kill}\land p(1)}^{i\rightarrow o^{*}}[k]&\sum_{n}\textbf{E}_{\textit{kill}\land p(2)}^{i\rightarrow o^{*}}[n]&\sum_{m}\textbf{E}_{\textit{kill}\land p(>2)}^{i\rightarrow o^{*}}[m]&\sum_{j}\textbf{E}_{\textit{kill}\land ts}^{i\rightarrow o^{*}}[j]\end{bmatrix}$
-Blind kill percentage (bkp)
-Denote the percentage of a player performing elimination while blinded.
-bkp refers to the ratio of the sum of elimination events where the attackers suffer a blinded effect to the number of eliminations.
-$\mathbf{F_{\textbf{bkp}}(i)}=\frac{\sum_{k}\textbf{E}_{\textit{kill}\land\textit{blind}}^{i\rightarrow o^{*}}[k]}{\sum_{n}\textbf{E}_{\textit{kill}}^{i\rightarrow o^{*}}[n]}$ (46)
-Critical hit percentage (chp)
-Denote the critical hit rate.
-Critical hit represents a successful damage event that applies to a particular hit group (typically, head) that deals more damage than a normal blow.
-chp refers to the ratio of the number of critical hits to the number of damage caused.
-$\mathbf{F_{\textbf{chp}}(i)}=\frac{\sum_{k}\textbf{E}_{\textit{headshot}}^{i\rightarrow o^{*}}[k]}{\sum_{n}\textbf{E}_{\textit{damage}}^{i}[n]}$ (47)
-Average kill per round (akpr)
-Denote the average number of eliminations per round by a player in a match.
-$\mathbf{F_{\textbf{akpr}}(i)}=\frac{\sum_{k}\textbf{E}_{\textit{kill}}^{i\rightarrow o^{*}}[k]}{\#\left(round\right)}$ (48)
-C.4 Props utilization features
-Flash efficiency index (fei)
-Denote how well a player utilizes the flash grenade. fei refers to the ratio of the total affection duration on the opponents minus that on the friendlies to the total affection duration.
-$\mathbf{F_{\textbf{fei}}(i)}=\frac{\sum_{k}\operatorname{\textbf{ELAPSE}}\left(\textbf{E}_{\textit{blind}}^{i\rightarrow o^{*}}[k]\right)-\sum_{m}\operatorname{\textbf{ELAPSE}}\left(\textbf{E}_{\textit{blind}}^{i\rightarrow a^{*}}[m]\right)}{\sum_{n}\operatorname{\textbf{ELAPSE}}\left(\textbf{E}_{\textit{blind}}^{i\rightarrow p^{*}}[n]\right)}$ (49)
-Props utilization index (pui)
-Denote how frequently a player utilizes props such as a high-explosive grenade, Molotov or incendiary, smoke grenade, or flash grenade.
-pui refers to the ratio of the sum of a player and the number of times different grenades are used to a fixed constant which reflects the maximum number of grenades used within this match for all players.
-$\rho$ denotes the maximum number of props a player is allowed to carry per round.
-In our case, $\rho$ is 5.
-$\mathbf{F_{\textbf{pui}}(i)}=\frac{\#\left(\textbf{HEG}(i)\right)+\#\left(\textbf{INC}(i)\right)+\#\left(\textbf{SMK}(i)\right)+\#\left(\textbf{FLS}(i)\right)}{\#\left(player\right)\times\#\left(round\right)\times\rho}$ (50)
-Appendix D Sense and Performance Features Classification
-All features presented in this section are recombinations of features from the first two subsystems, categorized into two main types, i.e., sense features and performance features.
-For a detailed description of each feature, please refer to Appendix B and Appendix C.
-The following merely enumerates the specific categorizations of the features and some brief descriptions.
-D.1 Sense Features
-Sense features indicate a player’s in-game cognition, emphasizing their strategic acumen and tactical foresight.
-  • Temporal Sense Features:
-    • Economy features: Financial transactions and asset management, mirroring strategic resource allocation.
-    • Movement features: Spatial patterns, positioning, and navigation dynamics, indicating tactical awareness.
-    • Flash features: Strategic deployment of flash grenades during in-game engagements.
-    • Grenade features: Utilization of assorted grenades, highlighting understanding of map control.
-  • Structured Sense Features:
-    • Flash efficiency index (fei): Metrics gauging flash grenade efficiency.
-    • Occluder-penetration index (opi): Achieve eliminations at specific locations, demonstrating proficiency with the map.
-    • Inertial shot percentage (isp): As a normal human being’s habitual firing after destroying an opponent.
-    • Blind kill percentage (bkp): Eliminations executed under visual impairment.
-    • Props utilization index (pui): Metrics showcasing prop usage across the whole match.
-D.2 Performance Features
-Performance features encapsulate a player’s in-game efficacy, concentrating on engagement aptitude, aiming precision, and mechanism mastery.
-  • Temporal Performance Features:
-    • Elimination Features: Efficiency of eliminations and some special, difficult eliminations.
-    • Weapon Fire Features: Proficiency metrics across weapon categories, illuminating combat adeptness.
-    • Damage Features: The player’s performance with respect to inflicting effective damage in-game.
-  • Structured Performance Features: The remaining structured features in Appendix C.
-Figure 13: Structured Features PCA results on both datasets.
-Appendix E PCA on Structured Features
-In Section 4.2.1, we have demonstrated the distribution distinction between cheaters and normal players.
-In the following section, we will demonstrate the PCA visualization on the structured features to display their orthogonality on both Aimbot and Wallhack datasets.
-Figure 13 illustrates the structured features PCA results for both datasets regarding the cheaters and normal players respectively.
-Horizontal and vertical axes represent the first principal component (PC1) and the second principal component (PC2), respectively.
-These principal components capture the directions of maximum variance in the data, showing where the data varies most.
-The unit circle in a correlation circle plot is a circle with a radius of 1, serving as a reference.
-Within this circle, the vectors (represented as arrows) demonstrate the magnitude and direction of the original variables’ associations with the principal components.
-The arrow represents feature vectors, where the direction and length of each arrow indicate the relationship of the original data features with the principal components.
-The direction of the arrow shows the correlation direction of the feature with the principal components.
-If two arrows are pointing in similar directions, it means their corresponding features exhibit similar patterns of variation in the dataset.
-Length indicates the impact or correlation strength of that feature on the corresponding principal component.
-Longer arrows suggest that the feature contributes more significantly to the variance captured by that principal component, and vice versa.
-It is noticeable that within the same plot, most of the features are distinct from others by obtaining different directions and lengths of the arrow.
-We can also observe that the direction and length of the arrows across Aimbot and Wallhack datasets are distinctive, indicating that the data distributions of the corresponding features are statistically different.
-Additionally, there are differences in the direction and length of the arrows between cheaters and normal players within the same or across different datasets.
-Above all, the original features can not be replaced by PCA.
-Appendix F Evaluation Metrics
-The following are the detailed evaluation matrices that are used in Section 6.3.
-True Positives (TP), shown in Equation 51, represents instances where the model accurately predicts the occurrence of cheating behavior.
-$TP=\sum_{i=1}^{n}I(y_{i}=1\land\hat{y}_{i}=1)$ (51)
-True Negatives (TN), represented in Equation 52, denotes instances where the model correctly identifies the absence of cheating behavior.
-$TN=\sum_{i=1}^{n}I(y_{i}=0\land\hat{y}_{i}=0)$ (52)
-False Positives (FP), represented in Equation 53, encompass instances where the model erroneously predicts the presence of cheating behavior.
-$FP=\sum_{i=1}^{n}I(y_{i}=0\land\hat{y}_{i}=1)$ (53)
-False Negatives (FN), represented in Equation 54, are instances where the model inaccurately predicts the absence of cheating behavior.
-$FN=\sum_{i=1}^{n}I(y_{i}=1\land\hat{y}_{i}=0)$ (54)
-In these equations, $y_{i}$ denotes the true label, $\hat{y}_{i}$ signifies the predicted label, and $I(\cdot)$ represents the indicator function.
-Recall, represented in Equation 55, measures the proportion of actual positive instances that were correctly identified by the model.
-$Recall=\frac{TP}{TP+FN}$ (55)
-Accuracy, as shown in Equation 56, gives a holistic measure of the model’s performance across both positive and negative classes.
-$Accuracy=\frac{TP+TN}{TP+TN+FP+FN}$ (56)
-AUC-ROC, denoted in Equation 57, is the area under the receiver operating characteristic curve, indicating the model’s capability to differentiate between positive and negative classes.
-It is computed by integrating the ROC curve, as shown below, where $TPR$ (True Positive Rate) is synonymous with Recall, and $FPR$ (False Positive Rate) is defined as $\frac{FP}{FP+TN}$.
-[…]
-NPV, shown in Equation 58, is the proportion of actual negative instances among all instances predicted as negative.
-$NPV=\frac{TN}{TN+FN}$ (58)
+Acknowledgments
+We thank 5EPlay for sharing and labeling the dataset.
+We appreciate Associate Editor Prof.
+Abderrahim Benslimane for arranging the review process, and three anonymous reviewers for their professional reviews.
+We also acknowledge the anonymous reviewers from IEEE S&P, ACM CCS, USENIX Security and NDSS for their constructive reviews on the past versions of this research paper.
+This work is partially supported by the NSFC for Young Scientists of China (No.62202400) and the RGC for Early Career Scheme (No.27210024).
+Any opinions, findings, or conclusions expressed in this material are those of the authors and do not necessarily reflect the views of NSFC and RGC.